### v6

# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

Abstract
I analyze the risk-coping behaviors among factory worker households in early 20th-century Tokyo. I digitize a unique daily longitudinal survey dataset on household budgets to examine the extent to which consumption is affected by idiosyncratic shocks. I find that while the households were so vulnerable that the shocks impacted their consumption levels, the income elasticity for food consumption is relatively low in the short-run perspective. The result from mechanism analysis suggests that credit purchases with local retailers played a role in smoothing short-run food consumption. The event-study analysis using the adverse health shock as the idiosyncratic income shock confirms the robustness of the results. I also find evidence that the misassignment of payday in data aggregation results in a systematic attenuation bias due to measurement error in the standard consumption smoothing regressions.
Keywords: consumption smoothing; credit purchases; daily budget data; local retailers; measurement error; risk-coping strategy; risk-sharing
JEL Codes: E21; N35
1 Introduction
Economic theory illustrates the mechanism behind the smoothing of consumption in complete markets (Arrow and Debreu 1954). A series of empirical studies have tested the smoothness of consumption among households in various economies under systematic empirical specification. Evidence shows that markets in contemporary developing countries are often imperfect.111 The early representative literature includes, for example, Townsend (1994), Townsend (1995), Ravallion and Chaudhuri (1997), Grimard (1997), and Fafchamps and Lund (2003). Morduch (1999), Townsend (1995), and Dercon (2004) provide a review of the earlier studies. Attanasio and Pistaferri (2016) and Meyer and Sullivan (2023) provide recent discussions. The households are vulnerable to risk, and the idiosyncratic shocks impact the levels of their consumption because risks are not fully insured in the markets. Therefore, consumption smoothing is an essential agenda for mitigating poverty, especially in developing economies. [^2] Smoothness of consumption is also an important factor in human capital formation. For example, Foster (1995) shows that children’s growth patterns are influenced by the inefficiency of the credit markets using the floods of 1988 in Bangladesh as a shock. Rose (1999) investigates the relationship between consumption smoothing and excess losses of girls, revealing that the favorable rainfall shocks in early life increased the survival probabilities of the girls in rural India. Gertler and Gruber (2002) also offers evidence that the smoothing of consumption against health shocks is fairly imperfect in Indonesia.
Historical economies offer experimental environments to investigate the means and degrees of consumption smoothing under less developed societies. Lessons from historical institutions in developed economies shall provide meaningful insights for today’s developing economies. Despite this advantage, little has been known about historical consumption smoothing behaviors, partially because of the lack of data on longitudinal household budgets. This study is to digitize a unique daily longitudinal budget survey on working-class households conducted in Tokyo around 1919 and employed systematic econometric methods to explore the consumption behaviors among factory worker households in a representative manufacturing area. The result shows that the full insurance hypothesis is strongly rejected for overall consumption expenditure. The estimated income elasticity is close to that for the rural households in southern India around 1980. This means that the factory worker households in historical metropolis suffered certain consumption losses when they faced idiosyncratic adverse income shocks. However, I also found evidence that while expenses on dispensable categories are likely to be cut in response to the idiosyncratic shocks, expenses on indispensable consumption categories are partially smoothed in the short-run. Mechanism analyses provide suggestive evidence that credit purchases in the local economy played an important role in mitigating adverse idiosyncratic shocks.
The first contribution of this study is to offer the first case study of a metropolis in a developing economy. Urbanization is now a general phenomenon in the world, and the shares of urban population in many developing countries, especially in East Asia, are reaching a similar level to the developed countries (Henderson and Turner 2020). This fact raises the need for a deeper understanding of consumption behavior among urban working households in the process of economic development. Despite this, there is still little evidence for the urban worker households in the literature. While Townsend (1995) offers the results for Bangkok in Thailand between 1975 and 1990, a large part of the existing studies focused on the rural economies (Townsend 1994; Ravallion and Chaudhuri 1997; Gertler and Gruber 2002). [^3] Recent field experimental studies also focus on the village economies (Comola and Prina 2023; Macours et al. 2022). A few exceptional studies provide case studies on the middle-sized contemporary economies (Skoufias and Quisumbing 2005; Aguila et al. 2017). In historical context, Ogasawara (2024) analyzes the factory workers’ households in Osaka City around 1920, which was roughly half the size of Tokyo City (Online Appendix Table A.1). This study considers a set of homogeneous households in a representative manufacturing area in Tokyo, which provides a valuable empirical setting for testing the risk-sharing behaviors in the target urban economy. [^4] Focusing on the sample sharing the similar occupational and regional characteristics systematically mitigates bias in estimating income elasticity under the standard risk-sharing regressions due to the heterogeneity in the risk preferences (Schulhofer-Wohl 2011; Mazzocco and Saini 2012). I found suggestive evidence of consumption smoothing using credit transactions with local retailers among working-class households in a historical metropolis.
Macroeconomics literature has investigated the theoretical mechanisms behind the false rejections of the null of the risk-sharing hypothesis that are sometimes observed in developed economies. An essential factor disturbing the valid statistical inference is measurement error in aggregate data (Cochrane 1991; Nelson 1994; Dynarski et al. 1997; Gervais and Klein 2010). This study reveals a specific generating process of systematic measurement errors in the aggregated income and expenditure series in the survey dataset. In my empirical setting, the factory workers had two paydays in a month: one was set in the middle and the other at the end of the month. Hence, the panel dataset based on the calendar months leads to a substantial mismatch between the aggregate income and expenditure series. [^5] If payday is set soon after the initial date in a certain time cell, it does not cause a mismatch in income and expenditure. However, when payday is set just before the initial date in the cell, it leads to measurement errors because the consumption schedule in that cell shall based on the earnings in the previous cell. In other words, this way of matching assumes that the households in a certain cell make their consumption schedule based on future earnings, which is unrealistic. I found that the aggregation ignoring this mechanism causes false rejections in most consumption subcategories due to the attenuation bias. This evidence shows the importance of reflecting the timing of paydays in the aggregation and invokes careful reviews of the existing results from the aggregate budget dataset, a typical format widely used previously.
The second contribution is to add a new analytical perspective to economic history studies. How working-class households overcame their economic hardships during the late 19th to the early 20th century Britain has attracted wider attention. Weights of evidence from these studies have revealed the roles of historical financial institutions and labor supply adjustments in compensating income losses in working-class households. [^6] Horrell and Oxley (2000) found the important roles of the labor supply adjustments among family members in coping with the income losses of the head in working-class households in late 19th-century Britain. Scott and Walker (2012) also found that interwar British working-class households had access to various institutions such as clothing clubs, hire-purchase, contractural savings, and insurance. The households with less earnings were likely to use credit-based institutions like clubs and hire-purchase, whereas those with higher incomes could use the savings institutions. Kiesling (1996) and Boyer (1997) focused on the risk-coping behaviors among the cotton workers households in Lancashire during the recession due to Civil War. Another strand of literature in financial history provided supply-side (lender-side) evidence on the functions of the micro-financial institutions for the workers (O’Connell and Reid 2005). However, there is little evidence of how those risk-coping strategies contributed to smooth household consumption. In addition, the previous studies have utilized the cross-sectional variations in income to estimate the elasticities for the temporary income sources. [^7] For example, Horrell and Oxley (2000) and James and Suto (2011) consider the deviations of actual earnings from the predicted earnings to sort out the shocks. While Scott and Walker (2012) considers the smoothness of expenditure rather than income based on the deviation from the mean expenditure, their analyses employed the cross-sectional variations in the pooled cross-sectional sample. An exceptional study is Saaritsa (2011), which shows the estimated elasticity to temporary income from a fixed effect one-way error component model using a longitudinal household budget survey in 1928 Helsinki. Hence, the estimated elasticity tells the mean tendency that the households with lower (higher) income were more (less) likely to use a specific risk-coping strategy. [^8] In other words, the elasticity is derived from the observations across different households with different income levels. However, since risk-coping is a concept about the dynamic behavior of a specific household, comparing the different time cells is more natural to evaluate the roles of the risk-coping strategies. By using the panel dataset, this study implements the stylized econometric method to systematically investigate the households’ smoothing mechanisms to offer comparable estimates.
Another historical contribution is adding micro-level evidence on consumer behaviors during industrialization. Francks (2009) provides a comprehensive view of the long-run consumption history in Japan from the Edo era to the twentieth century. [^9] There are several business history studies analyzing the relationship between consumption and production. A representative work is Gordon (2013), who investigated the expansion of the sewing machine from many aspects, including the production, distribution, and gender social norms in Japan. However, these studies have little interest in the households’ risk-coping behaviors. A few studies investigated on the relationships between household chores and consumption behaviors in interwar households (Tanimoto 2011, 2012). This study employs a different analytical view by focusing on short-run consumption behaviors.Looking into the extent and how the local society could ensure the households’ consumption would provide valuable insights into the consumption history literature.
The rest of this paper is organized as follows. Section 2 overviews the Japanese economy around the First World War (WWI) and reviews Tsukisima, the target area in my analysis. I also summarize the financial institutions that were accessible to working-class households. Section 3 introduces the documents used to construct my panel dataset and discusses the sample characteristics. Section 4 summarizes the empirical strategies and results. Section 5 overviews the credit purchases in the local economy. Section 6 concludes.
2 Historical Background
2.1 Japanese Economy around the First World War
Japan had the foundations of an industrial nation by WWI. Between 1885 and 1915, real GNP grew annually by 2.6%, while population increased at 1.1%. [^10] Given that economic growth is faster than population growth, this implies that modern economic growth had started in this period as defined by Simon Kuznets (Miyamoto 2008, pp. 60; 64). In 1915, the percentage of the employed population in the primary, secondary, and tertiary industries were 63%, 20%, and 18%, respectively. Based on domestic genuine production, these industries accounted for just under 30%, over 30%, and approximately 40%, respectively.
Soon after the outbreak of WWI, exports from European countries declined. Japan increased exports to the Asian market on behalf of those European countries and exported munitions and foodstuffs to the European Allies, leading to a current account surplus. Given the declines in the export capacity of the European countries, import substitution had progressed. This expands the domestic market and leads to a rise in the machinery, metal, and chemical sectors. A worldwide shortage of ships also stimulated the development of shipbuilding, which proliferated as a leading industry. The development of the shipbuilding industry stimulated demand for the machinery and steel industries. Through this process of heavy and chemical industrialization, the total number of enterprises increased. [^11] Descriptions about the process of heavy chemical industrialization are based on Nakamura (1971, p. 128), Nakamura and Odaka (1989, p. 10), and Sawai and Tanimoto (2016, p. 250). Year-on-year growth in the number of incorporated firms was 17.8% in 1918, 22.3% in 1919, and 13.5% in 1920. These are the top three years of the 34 years from after the Russo-Japanese War to 1940 (Nakamura and Odaka 1989, p. 22). Figure 1 shows that the GNP growth from 1915-19 was 11.5% in nominal terms (4.4% in real terms), the highest growth throughout the 1910s and 1920s.
This rapid industrialization impacted the Japanese economy in many ways. Due to export-led economic growth, large inflows of gold increased domestic currency balances, causing prices to soar. Although wages began to rise around 1916 as a response to rising prices and labor demand due to the increased number of firms, wage increases failed to keep pace with soaring prices. By the end of the war, the number of labor disputes grew sharply, and the labor movement flourished. Furthermore, the rise of business accelerated urbanization as the rural population moved to cities in search of job opportunities. The share of large cities (with people of 50,000 or more) as a percentage of the total population was less than 10% in 1889 but increased to nearly 16% in 1920 (Online Appendix Figure A.1a). The growth rate was exceptionally high during the WWI period, indicating that structural changes through the war accelerated the rise of the urban population. Urbanization has increased public investment in large cities, especially in construction investment in social capital such as roads and water supply systems. [^12] Descriptions on the export-led growths, urbanization, and investments are from Nakamura and Odaka (1989, pp. 21–23) and Nakamura (1971, pp. 142; 147).
Exports in the 1920s were lower than before because of the appreciation of the yen against the real exchange rate due to the accumulation of domestic and foreign species during WWI. Since the current account deficit was settled through the disbursement of foreign species owned by the government and the Bank of Japan, there was no reduction in the domestic money supply despite the sizeable current account deficit. This resulted in the domestic prices, especially prices of nontradable goods, being kept high. The first half of the 1920s saw a temporary decline in industrial investment as the capital investment boom of the WWI period came to a halt under declining demand. [^13] On the other hand, construction investment by the private sector, including electric power and private railways, increased (Nakamura 1971, p. 145). Descriptions about the post-war economy in this paragraph are from Sawai and Tanimoto (2016, pp. 253–254).
Figure 1: Gross National Product per Capita and Personal Consumption Expenditure (in yen)
Note: This figure shows the gross national product per capita (solid line) and personal consumption expenditure (dashed line) between 1910 and 1926. The prices are deflated using the CPI in 1934-1935 price. Sources: Created by the author using Ohkawa et. al. (1974, p. 237).
Figure 1: Gross National Product per Capita and Personal Consumption Expenditure (in yen)
As described, there were short-term business cycles, such as the boom caused by WWI and the subsequent recession. However, taken as a whole, Japan achieved a high economic growth rate relative to other countries in the early 20th century and before WWII. As shown in Figure 1, the growth rates of personal consumption expenditure were particularly high during and immediately after WWI, with average growth rates reaching 112% in 1915-19 and 118% in 1920-24.141414Total consumption as a percentage of the GNP remained above 80%. Although this level is not apparent among other countries (Kuznets 1962), the growth rate of real personal consumption per capita was relatively high: from 1875 to 1939, the average growth rate of personal consumption was 1.36%, and the rates never turned negative except during the Russo-Japanese War period. This was a high level even by international standards. Given the rapid urbanization, therefore, it is plausible to focus on the consumption behavior of urban households when considering the Japanese economy in this period. However, progress has yet to be made. [^15] Descriptions on the personal consumpsion expenditure are from Nakamura (1971, pp. 2–5; 27; 30–31).
2.2 Tsukishima: A Representative Manufacturing Area
The population in Tokyo City has been growing since the end of the 19th century, and industrialization due to WWI spurred its growth. Tokyo City’s population share increased from 2.5% in 1889 to approximately 4% by 1920.161616In 1920, there were more than 10,000 municipalities in Japan, suggesting how huge Tokyo was. There was a difference of about 1 million between Tokyo and Osaka, the second-largest city by population. Online Appendix Figure A.1a summarizes the trends in the number of people in the top five cities between 1889 and 1920. Throughout the 1910s, the percentage of industrial and commercial workers increased in Tokyo City, building the city as a gigantic commercial and industrial place (Tanimoto 2002, p. 8). Column 1 of Panel A in Table 1 shows that nearly half of the male workers in Tokyo worked in the industrial sector as of 1920.
Figure 2: Administrative Wards in Tokyo City
Notes: The name in each lattice indicates the administrative ward in Tokyo. The gray ward indicates the Kyobashi Ward and an island in the Kyobashi Ward indicates Tsukishima. The border surrounded by Asakusa, Honjyo, Nihonbashi, and Fukagawa wards includes the Sumida River. The borders of each ward are based on the administrative district in 1920. Geographical coordinate system is based on JGD2000/(B, L). Source: Created by the author using the official shapefile (Ministry of Land, Infrastructure, Transport and Tourism, database).
Figure 2: Administrative Wards in Tokyo City
Figure 3: Factories in Tsukishima
Notes: Blocks in green and light green indicate the shipyards. Blocks in red and pink indicate the factories in the machinery sector. Blocks in blue and light blue indicate the factories in the other manufacturing sectors. Blocks in brown are the warehouses. Small circles colored in black and white show the workplaces in the machinery sector and the other manufacturing sectors, respectively. Source: Department of Health, Ministry of the Interior 1923c, second map. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
Figure 3: Factories in Tsukishima
The subject of this paper is Tsukishima, an industrial area located southeast of Kyobashi Ward in Tokyo (Figure 2). [^17] The base of the island was reclaimed during the Edo and Meiji periods from a delta at the mouth of the Sumida River. Online Appendix A.3 summarizes the development history of Tsukishima in detail. Column (2) of Panel A in Table 1 shows that Kyobashi had an industrial composition similar to that of Tokyo City in 1920. Column (3) confirms that Tsukishima is a representative industrial area in the city; approximately 68% of male workers were employed in the manufacturing sector. In the same year, there were 214 factories on this island, and approximately 80% of them belonged to the machinery and related sectors (Department of Health, Ministry of the Interior 1923a, pp 385–389). [^18] According to the census of 1920, the number of workers in the metal and machinery equipment manufacturing industry ($27.1$%) was the highest compared to other industries in the entire city of Tokyo. The same tendency was observed in Kyobashi Ward ($35.5$%). Although similar statistics are not available for Tsukishima, this supports the evidence that Kyobashi was not an area with a unique industrial pattern (Statistics Bureau of the Cabinet 1929c, pp. 85–86; 108–109)). Figure 3 shows the spatial distribution of factories in Tsukishima as of June 30, 1918. Tsukishima had many machine factories and shipyards, and the workers’ residences were clustered there. [^19] For example, a representative shipbuilding plant was located on Ishikawajima, an island in the north of Tsukishima (Figure 3). Ships were strongly influenced by the import substitution that was underway at the beginning of the 20th century, and domestic production of ships was progressing (Section 2.1). The production of a hydraulic 80-ton hardened crane by Ishikawajima Shipyard in 1917 is an example of the domestic output of a production good that had previously relied on imports (Nakamura and Odaka 1989, pp. 10; 32). The focus of this study is the consumption behavior of these worker households in the industrial sector.
Table 1: Industrial and Family Structures in Tsukishima
Panel A: Industrial structure
Name of survey (1) 1920 Population census (2) 1920 Population census (3) 1920 Population census
Survey area Tokyo City Kyobashi Ward Tsukishima
Survey subject Complete survey Complete survey Complete survey
Survey month and year October 1920 October 1920 October 1920
Agriculture 1.0 0.6 0.4
Fisheries 0.1 0.3 0.8
Mining 0.4 0.4 0.8
Manufacturing 44.5 45.2 67.7
Commerce 32.4 34.0 16.9
Transport 7.6 10.2 6.1
Public service and professions 11.3 10.2 4.6
Housework 0.3 0.1 0.2
Other industry 2.4 2.0 2.5
Panel B: Demographic structure
Name of survey (1) 1920 Population census (2) 1920 Population census (3) 1920 Population census
Survey area Tokyo City Kyobashi Ward Tsukishima
Survey subject Complete survey Complete survey Complete survey
Survey month and year October 1920 October 1920 October 1920
Average household size (in person) 4.6 4.7 4.3
Share of married males (%) 38 36 40
Sex ratio (males/females) 1.2 1.2 1.3
    0–13 years 1.1 1.1 1.1
    14–19 years 1.5 1.7 2.2
    20–29 years 1.3 1.4 1.6
    30–39 years 1.2 1.2 1.3
    40–49 years 1.2 1.2 1.3
    50–59 years 1.1 1.1 1.2
    60+ 0.8 0.8 0.9
Average age of males 25.3 25.3 25.1
Notes: Panel A: This panel summarizes the industrial structures based on the occupations of male workers measured in the 1920 population census. Each share is calculated as the number of males worked as regular workers in each sector divided by the total number of male workers (in percentage). All the figures do not include the number of unemployed (mugyō sha) males. Panel B: The average household size is the number of people divided by the number of households. The share of married males is the number of married males living with/without spouses divided by the total number of males (in percentage). The average age of males is calculated using the tables for population by age group reported in the census. Since the tables are open-ended, the class for over 60 years old is rounded as $64.5$ following the range of the former category (i.e., 50–59). All the figures in this panel do not include a small number of quasi-households (jyun setai), which include the person, such as the students in dormitories and the patients in hospitals. The figures for Tsukishima listed in column 3 of both panels are based on the total sum of statistics of all the blocks (chōme) in Tsukishima (i.e., from Tsukudajima to Tsukishima dōri jyūni chōme). Sources: Columns 1 and 2 of Panel A use the Statistics Bureau of the Cabinet (1929c, pp. 18–19). Columns 1 and 2 of Panel B use the Statistics Bureau of the Cabinet (1929c, pp. 38–43); Tokyo City Office (1922b, pp. 2–3; 42–46). Column 3 of Panel A uses the Tokyo City Office (1922c, pp. 42–47). Column 3 of Panel B uses the Tokyo City Office (1922a, pp. 26–31); Tokyo City Office (1922b, pp. 2–3; 42–46).
To look more closely at the characteristics of working-class households in Tsukishima, Panel B in Table 1 summarizes the population and occupational statistics from the census. First, the average household size in Tsukishima is $4.3$, similar but slightly smaller than the values in Tokyo City ($4.6$) and Kyobashi Ward ($4.7$). The percentage of the married male population is $40$%, which is in the similar range of the figures of Tokyo ($38$%) and Kyobashi ($36$%). The average age of men in Tsukishima is $25.1$, slightly lower than the figures for Tokyo and Kyobashi. While the sex ratio by age group in Tsukishima shows a similar trend to that of Tokyo and Kyobashi, the ratios for the 1414141419191919 and 2020202029292929 age bins are higher than those. Overall, these statistics reflect the fact that Tsukishima is a typical industrial area and many young male workers were employed there. The slightly smaller household size in Tsukishima can be attributed to the relatively large number of young factory worker households. The percentage of married men in Tsukishima suggests that a certain proportion of the households in their early 20s might have been couples before they started raising children.
Finally, I briefly overview the lives of the factory workers’ households in Tsukishima. Factories and workers’ residences were scattered throughout the island. [^20] Tsukishima is composed of Ishikawajima island at the northern end, the adjacent Tsukudajima and Shin-tsukudajima islands, and the rectangular Lots 1–3 in the south (Online Appendix Figure A.3). Specifically, factories and worker residences were scattered throughout Shin-tsukudajima and Lots 1 and 2, whereas Lot 3, a relatively new reclaimed site completed in 1913, was undeveloped. The factory workers had lived in similar tenements and used the commercial stores and restaurants in the central lot. [^21] Online Appendix A.4 summarizes the housing in Tsukishima in detail using a housing survey. Regarding the stores, there were 209 stores selling daily necessities, 109 selling clothing, 74 selling tools, and 120 selling food, drink, and entertainment. As for medical personnel, there were eight doctors, four dentists, eight obstetricians, and 14 drug stores. There were also ten bathhouses (Department of Health, Ministry of the Interior 1923a, pp. 41–47). There were two postal offices, 13 pawnshops, and five moneylenders, and those were scattered throughout the island (Online Appendix Figure A.2). The main routes to the mainland were the wooden Aioi Bridge (Aioi bashi) and three ferries (watashi bune). [^22] See Online Appendix Figure A.3 for these routes. The average daily total number of people using the ferries in 1918 was reported to be $53,318$, accounting for more than twice as many people as the Tsukishima population. [^23] See Department of Health, Ministry of the Interior (1923a, p. 30). The number of people and households of Tsukishima at the end of December 1918 is reported to be $24,399$ and $5,562$, respectively (National Police Agency 1920a, pp. 50–51). These descriptions illustrate Tsukishima as a busy machine industry area during the industrialization.
2.3 Risk-Coping Institutions
Working-class households had been exposed to several risks that led to losses of income at that time. Generally, social insurance, such as health, accident, and unemployment insurance, can help mitigate the idiosyncratic income shocks on the workers. [^24] Gertler and Gruber (2002) shows the impacts of illness on consumption smoothing behavior in early 1990s Indonesia and argues the importance of public support for the illness. Kantor and Fishback (1996) show that the installation of workers’ compensation reduced the precautionary savings in the interwar U.S.. Gruber (1997) provides evidence that unemployment insurance smooths individual consumption as the case of the U.S. from 1968–1987. However, comprehensive social insurance systems did not exist in the early 1920s in Japan. The former health insurance was enacted in 1927 and unemployment insurance had not been established throughout the prewar period. Therefore, the workers had basically no public support for their illness, injuries, and joblessness. Instead, they might have used private financial institutions to deal with the idiosyncratic risks.
Several financial institutions were indeed available for working-class households. A representative savings institution for households was the postal savings (yūbin chokin). In Tokyo City, roughly $1.4$ million people had postal savings accounts, covering approximately $67$% of the entire citizens in 1920 (Online Appendix Table A.2). The average savings amount per capita in Kyobashi Ward was $44.5$ yen (Tokyo City Office 1921, pp. 892–893), which is less than one month’s earnings of the skilled-factory workers (Section 3.2). Savings bank (chochiku ginkō) is another famous savings institution among the working class in the 1910s (Ito and Saito 2019, pp. 79–81). The number of people with savings accounts was roughly $1.4$ million, accounting for approximately $58$% of total people in Tokyo (Online Appendix Table A.2). However, $84.8$% of all the depositors were the commerce and miscellaneous workers, and the manufacturing sector shared only $11.5$% in Kyobashi Ward (Online Appendix Table A.3). This reflects the fact that savings bank was an institution designed for lower-income workers than factory workers (Tokyo Institute for Municipal Research 1925b, p. 81). This is consistent with the fact that Tsukishima did not have any savings bank branches, while there were two postal offices (Section 2.2). Although another savings institution was the mutual loan association (mujin), it had been used among small business owners and merchants for running their businesses (Online Appendix A.2). [^25] Generally, the ordinary bank in prewar Japan was the institution for providing large loans to businesses. There was no bank in Tsukishima, as workers did not borrow money from banks to make ends meet. The official report states that “a commercial bank is not necessary in a laboring area such as this Tsukishima, and even for large entrepreneurs such as factory owners, there is no particular need to establish a bank for their business transactions, and transactions with banks in other areas are sufficient” (Department of Health, Ministry of the Interior 1923a, p. 48).
There also be a few lending institutions. Credit purchases (kakegai) in retail shops were the most popular institution. Although the systematic statistics on credit purchases are unavailable, the THBS households often used credit purchases in various retail shops for rice, fish, vegetables, firewood, and charcoal. [^26] Cooperative societies were also available for worker households to buy daily commodities, albeit they were not lending institutions. There were 26 cooperative societies in Tokyo prefecture in 1924, of which 23 were organized by the citizens and workers (Central Federation of Industrial Associations 1925, p. 57–59). However, while the two largest societies had roughly $4,000$ members, the others were generally small societies with less than 100 members. This figure accounts for only a few percentage points in the workers in the city, given that the number of male workers was more than 768 thousand. Most of the THBS households had not used the cooperatives. I will describe this institution in detail in Section 5. The pawnshop was another lending institution. The accessibilities were higher than the other lending institutions: [^27] The number of pawnshops in Tokyo were $1,334$ in 1918 and $1,261$ in 1919, respectively. The total number of cases was $8,226,883$ and $7,573,406$ each year. These figures are more than three times greater than the entire population in Tokyo City. there were $13$ pawnshops in Tsukishima, more than twice the large number of moneylenders (Section 2.2). However, the core users of the pawnshops were the low-income working-class households rather than the factory workers (Tokyo City Social Affairs Bureau 1921, p. 9). According to the statistics on the pawnshop users in Tokyo prefecture in 1923, the most representative users were the day laborers and workers classified as “miscellaneous”, accounting for $63.4$% of all users. In contrast, those for agriculture, commerce, and manufacturing sectors were only $1.7$, $19.5$, and $15.6$%, respectively (Tokyo Institute for Municipal Research 1926, p. 25). The inexpensive article means that pawnshops were used to deal with very short-run necessities only. Among the pawnshops of Kyobashi, for example, the average amount per case was $7.8$ yen which accounts for $12$% of the monthly disposable income of the THBS households (Panel A of Table 3). In this light, the pawnshop is a small lending institution for low-income working-class households who might have faced credit constraints. Lenders did not need to screen the borrower’s credit, and the borrowers did not worry about incurring heavy debt as the main articles were inexpensive clothes (Shibuya and Shojiro 1982). In addition, the interest rates were regulated by the Pawnbroker Regulation Act of 1895, and the average redemption rate was substantially high under the lower interest rates.  [^28] This means that the average redemption rate was substantially high: it was approximately $94$% in 1920 among $88$ pawnshops in Kyobashi ($401,222$ cases/$427,265$ cases) (Tokyo City Office 1922b, pp. 888–889).
Other lending institutions are money lenders (kinsen kashitsuke gyō) and credit unions (shinyō kumiai). Generally, moneylenders were used by business owners with very high-interest rates (Shibuya 2000, pp. 184; 248). Similarly, the coverage of the credit unions was substantially low, say a few percentage points of the male workers in the manufacturing sector (Online Appendix A.2). Informal gifts may be another type of risk-coping institution, although systematic statistics are not available. A household survey on the 185 factory-worker households in Tokyo suggests that the share of average monthly gifts in total income was approximately $3$% in November 1922 (Social Affairs Division 1925, pp. 58-59). While this means that the gifts comprised a few percentage points of the monthly income, it is unclear whether these informal income transfers had functioned as mutual aid.
To summarize, while there were several institutions, the available institutions for the factory workers were few in the practical sense. For savings institutions, the postal savings might have been the best available devices for Tsukishima workers. Regarding lending institutions, credit purchase is the plausible device among factory workers. While pawnshops were accessible financial institutions, they might not provide sufficient money as the articles were usually inexpensive. The factory workers in Tsukishima might have combined these devices to deal with the idiosyncratic income shocks.
3 Data
This section is designed to understand the characteristics of the data used in the study. I give an overview of the Tsukishima Survey in Section 3.1. I then utilize the statistics from the complete surveys in Section 3.2 to assess the representativeness of the sample households. In Section 3.3, I explain my data aggregation method to overcome the potential measurement error issue. The variables used, the trend in income and expenditure, and variations in idiosyncratic shocks are described in Section 3.4.
3.1 Tsukishima Household Budget Survey
The Tsukishima Survey was the first urban social survey conducted in Japan by the Ministry of Home Affairs around 1919. The purpose was to investigate the status of the lives of urban worker households amid the rapid progression of industrialization throughout WWI. It was designed as a field survey consisting of multiple survey items, including the budget survey (hereafter THBS). [^29] Online Appendix B.1 summarizes detailed information about the Tsukishima Survey. The THBS was the first household survey in Japan to use the budget book method and is considered the prototype for subsequent household surveys in Japan (Sekiya 1970, p. 43). The Ministry of Interior published an official report on the Tsukishima Survey in 1921, which contains some aggregated information on the target households. However, household-level budget information is required to analyze consumption smoothing behaviors.
Fortunately, the household budget book (kinsendeiri hikaechō) gathered in the THBS is preserved in the archives of the Ohara Institute of Social Research (hereafter, OISR). [^30] The author is permitted to access the materials related to the Tsukishima Survey by the OISR: Archives of the Tsukishima Survey (THBS, unreleased). While most of the sources are unreleased, the official reports of the survey published by the Ministry of Home Affairs are included in the archives of Gonda Yasunosuke at the OISR and are available to the public (The OISR, Archives of Gonda Yasunosuke (7-2; 7-3; 7-4)). These budget books were transferred to the OISR by Iwasaburo Takano, who was responsible for the Tsukishima Survey as a member of the 7th Section of the Health and Hygiene Research Committee of the Ministry of Home Affairs and became the first director of the OISR. [^31] Iwasaburo Takano was a professor at Tokyo Imperial University and a specialist in social statistics and social policy. He emphasized gathering statistics about workers through social surveys and is known as a researcher who significantly contributed to the formation of social survey methodology in Japan. His contribution to the Tsukishima Survey is summarized in detail in the Online Appendix B.1.
I corrected and digitized all the 40 THBS budget books for the factory worker households that were included in the official report. [^32] The archives of the OISR also reserve several other household budget books of factory worker households. However, these budget books are generally in poor record keeping. Therefore, I do not use these incomplete books in this study. The characteristics of the remaining household budget books are summarized in detail in the Online Appendix B.5. During the study period, each household recorded their income and expenses daily in a budget book. Notably, a survey office was set up in Tsukishima, where a few surveyors lived and were closely connected with the survey households. Furthermore, weekly administrative meetings were held to discuss the progress of the survey (Miyoshi 1980, p. 38). This confirms that the information in these budget books is reliable and that the THBS budget dataset ensures a quality that allows quantitative analysis. [^33] Online Appendix B.1 provides finder details on how the Tsukishima Survey had been conducted carefully as a social survey project of Takano.
3.2 Sample Characteristics
Analytical Sample
The estimation strategy used for testing consumption smoothing and risk-coping behaviors requires the panel structures for aggregate income and expenditure (Section 4). Among the 40 THBS households, there are five cross-sectional units in the semi-monthly and monthly data. [^34] Section 3.3 provides details on the definitions of aggregate income and expenditure. I excluded these five cross-sectional units, leaving 35 THBS households. Next, a household with incomplete income information, which is necessary for my analysis, is removed. Finally, I exclude a household for which no information on payments to the credit purchases was recorded. [^35] Online Appendix B.3 summarizes the finer details of the original data structure and trimming. Consequently, my THBS sample includes 33 households with panel structures from 12th January to 11th November 1919.
Before assessing the representatives of the THBS sample, I test the potential selection bias due to attrition of the units. The panel units with shorter (longer) time-series observations might have different preferences on the consumption than those with longer (shorter) time-series observations. In other words, the households with shorter panels should have similar family size characteristics to those with more extended panels. To test this potential issue, I regressed an indicator variable for the shorter panel units on the family size variables. Online Appendix Table B.2 summarizes the results. All the estimated coefficients on the covariates are close to zero and statistically insignificant, and the Wald statistics support the null results. This result is robust to using different thresholds such as first quantile, median, and third quantile. This supports the evidence that the household’s preference is less likely to be correlated with the unbalancedness.
Representativeness
Table 2: Assessing Representativeness: Tsukishima Household Budget Survey
Panel A: Manufacturing sector
Name of survey (1) 1919 Statistics of (2) 1920 Tsukishima Factory (3) The THBS
National Police Agency Survey
Survey area Tsukishima Tsukishima Tsukishima
Survey month and year December 1919 November 1920 1919
Survey subject All factories with All factories with # of 33 households
# of workers $\geq 15$ workers $\geq 15$ workers $<15$
Textile 0.2 2.9 0.0 0
Machinery 87.4 85.3 91.8 93.9
Chemical 1.1 5.9 5.5 0
Food 1.2 1.5 0.0 3.0
Miscellaneous 10.2 4.4 2.7 3.0
Observations $7,647$ workers 68 factories 146 factories 33 heads
Panel B: Family structure
Name of survey (1) 1920 Population (2) 1919 Statistics of (3) The THBS
Census National Police Agency
Survey area Tsukishima Tsukishima Tsukishima
Survey month and year October 1920 December 1919 1919
Survey subject Complete survey Complete survey 33 households
Average household size 4.3 4.3 4.3
Sex ratio 1.3 1.2 1.2
Household size (share)
     1 6.0 0
     2 16.7 (18.7) 15.2
     3–5 50.8 (57.1) 57.6
     6–8 21.5 (24.2) 27.3
     9+ 4.9 0
Pearson $\chi^{2}$ statistic $p$-value 0.948
Panel C: Monthly earnings
Name of survey (1) The THBS (2) Manufacturing Census
Survey area Tsukishima Tokyo City
Survey/equivalent year 1919 1919
Survey subject THBS heads Male factory workers
(Ave. age: 33.2) in machinery factories
(Age range: 30-40)
Average monthly earnings 52.7 55.0
95%CI [48.0, 57.4]
Notes: Panel A: Column 1 shows the share of factory workers to the total factory workers in each sector (in percentage). The number of factory workers includes female and male factory workers in the 68 factories, with 15 workers and more. Column 2 indicates the share of factories to the total number of factories in each sector by factory size (in percentage). Column 3 summarizes the manufacturing sectors for the heads of the THBS. The definition of classification in the Statistics of the National Police Agency (column 1) is used in columns 2 and 3. Following this definition, several “woodworking” factories measured in the Tsukishima Factory Survey are included in the machinery sector in column 2, as it consists of shipbuilding and wooden pattern factories. The report also supports that the wooden patterns are used in the casting process and shall be regarded as one of the machinery sector (Department of Health, Ministry of the Interior 1923a, pp. 388–389). Panel B: The first and second rows in columns 1 and 2 show the average household size and sex ratio measured in the 1920 Population Census and Statistics of the National Police Agency of 1919, respectively. The third to eighth rows in columns 1 and 3 show household size distributions in the census and the THBS sample, respectively. The figures in parentheses in Column (1) are the percentage share for family size bins between 2 and 8 people. The ninth row in column 3 shows the $p$-value from the Pearson $\chi^{2}$ test for the equality of the household size distributions between 2 and 8 people. The share in each household size bin measured in the census is used as the theoretical probability in calculating the test statistic (Online Appendix B.5). Panel C: Column 1 lists the average monthly earnings of the THBS heads calculated using the adjusted monthly panel dataset. Gaussian-based 95% confidence interval (CI) using bootstrap standard error is reported in the brackets. Column 2 shows the average monthly wage of the male factory workers in the skilled age range (30-40 years old) in machinery factories in Tokyo City. This figure is calculated using the average monthly wage and average monthly ancillary wage measured in the manufacturing censuses. All the wage statistics are deflated using the CPI for cities provided by Ohkawa et al. (1967, p. 255). Online Appendix B.6 summarizes the finer details of the process of this wage estimation. Sources: Panel A: Column 1 uses the National Police Agency (1920b, p. 236). Column 2 uses the Department of Health, Ministry of the Interior (1923a, p. 385–386). Column 3 uses the Department of Health, Ministry of the Interior (1923a, p. 154). Panel B: Column 1 uses the Tokyo City Office (1922a, pp. 26–31; 262–283); Tokyo City Office (1922b, pp. 2–3; 42–46). Column 2 uses the National Police Agency (1920b, p. 53). The figures in column 3 are calculated by the author using the THBS dataset. Panel C: Column 1 uses the THBS dataset. Column 2 uses the Tokyo City Statistics Division (1926a, pp. 16; 244–245); Tokyo City Office (1921, pp. 726–757).
Next, I assess the representativeness of the THBS sample in Tsukishima using the manufacturing and population censuses,
Panel A of Table 2 summarizes the share of male workers in the manufacturing industry by sector. Statistics of the National Police Agency (hereafter SNPA) focusing on factories with more than or equal to 15 workers show that roughly 90% of male workers had been engaged in the machinery sector in 1919 (Column 1). Column 2 summarizes the statistics in the manufacturing census conducted in the Tsukishima Factory Survey. The proportion of machine factories is $85$% in larger factories and $92$% in smaller factories, indicating that most of the male workers worked in the machinery sector.363636 I have confirmed that the paternal occupational statistics for the primary school students in Tsukishima provide materially similar results (Online Appendix A.5). In Column 3, I confirm that the THBS heads’ occupations show a similar proportion: $94$% worked in the machinery sector. In fact, the report states that “It is not too much to say that most of the industry in Tsukishima is machinery. Therefore, to describe the labor situation in Tsukishima, it is sufficient to describe its machinery industry” (Department of Health, Ministry of the Interior 1923a, p. 389). Figure 3 illustrating a number of machinery factories supports this statement. As for the other sectors, the THBS includes a few heads who worked in food and miscellaneous sectors, and the proportions are similar to those listed in column 2.
Panel B of Table 2 summarizes the family structure. The average household sizes are identical in the Population Census (Column 1), SNPA (Column 2), and THBS (Column 3). The sex ratio is also similar across the surveys (1.21.21.21.21.31.31.31.3). The number of people in the THBS households ranges from 2 to 8, meaning the THBS sample does not include single or very large households. [^37] Those family households with 2–8 people cover most of the households in Tsukishima because the share of single and vary large (9+ people) households was roughly $10$% (Column 1 of Panel B). In other words, the THBS mainly focuses on the consumption behavior in couples and households with a few children. For this household size range, the Pearson $\chi^{2}$ test does not reject the null of equality to the household size distribution measured in the census ($p$-value = $0.948$), meaning that the THBS sample reasonably approximates the distribution of family households in Tsukishima. Online Appendix B.5 illustrates the household size distributions in finer detail. This confirms the validity of the statement in the report that the THBS households can be “regarded as representative of the family form in Tsukishima” (Department of Health, Ministry of the Interior 1923a, p. 145).
The official report of the Tsukishima Survey suggests that roughly nine out of ten factory workers in the skilled-age range were classified as skilled workers. Roughly six of those skilled workers were employed in the middle- to large-scale factories, and three worked in small-scale factories. [^38] Online Appendix A.5 summarizes this point in detail using the complete survey on the primary school students in 4th to 6th grades in Tsukishima. Given this, the THBS was designed to investigate the budgets in skilled workers’ households. [^39] Shortly, Iwasaburo Takano, a survey director, regarded that the skilled workers in the modern industrial sector will constitute the core of the labor force, labor movement, and citizens. The finer details in the background of this survey design are summarized in Online Appendix B.1. The average THBS head’s age is 33, which is indeed in the skilled workers’ representative age range in the machinery factories (Kitazawa 1924). [^40] Specific occupations are not available for most heads because they usually list “factory workers” in the occupation section. However, sub-categories available at a few heads indicate the fact that they were lathe operator and finisher, the typical skilled workers in machinery factories. Online Appendix A.6 summarizes the works in the machinery factories in Tsukishima. Column 1 of Panel C in Table 2 indicates that the average monthly earning of the THBS heads is $53$ yen. In Column 2, I calculated the average monthly earning of male machinery factory workers in the skilled-age range using the manufacturing censuses, which shows similar monthly earnings ($55$ yen). [^41] Shortly, this figure is calculated using the average daily wage of male factory workers aged 30-40 and average annual working days measured in the manufacturing censuses. Online Appendix B.6 provides the finer details of the calculation steps. The average wages for the male factory workers had an inverted-U-shaped distribution with respect to age, taking the greatest figure in the 30s (Online Appendix Figure B.7). This is a similar life-cycle pattern to the factory workers observed in late 19th century England (Horrell and Oxley 2000, p. 42). This provides evidence that the monthly earnings of the THBS heads and skilled workers’ households in Tsukishima are materially similar. To summarize, this evidence indicates that the THBS sample is useful in inferring the mean tendency of those households.
3.3 Aggregation
Measurement Error
Figure 4: Relationships between Average Income and Expenditure
by Different Time-series Frequencies
(a) Daily panel dataset:
income and expenditure
(b) Daily panel dataset:
net income
(c) Semi-monthly panel dataset:
income and expenditure
(d) Semi-monthly panel dataset:
net income
(e) Unadj. semi-monthly panel dataset: income and expenditure
(f) Unadj. semi-monthly panel dataset: net income
Note: Figure 4a4c, and 4e illustrate the time-series plots of the average daily, semi-monthly, and unadjusted semi-monthly income and expenditure, respectively. Figure 4b4d, and 4f illustrate the time-series plots of the average daily, semi-monthly, and unadjusted semi-monthly net income (i.e., income minus expenditure). Figures 4a and 4b shows the daily data between 12th January and 11th November 1919. The (adjusted) semi-monthly series defines the 12-26th for the first and 27-11th for the second half. Figures 4c and 4d show the semi-monthly series calculated using the daily data between 12th January and 11th November 1919. The unadjusted semi-monthly series uses the 15th of each month as a threshold. Figures 4e and 4f show the unadjusted semi-monthly series calculated using the daily data between 15th January and 31st October 1919. Source: Created by the author using the THBS sample.
Figure 4: Relationships between Average Income and Expenditure
by Different Time-series Frequencies
It is necessary to summarize the daily budget information into a more extended time range because households need a set of time to smooth their consumption given a realized amount of income. To explore the reasonable way of aggregation, I first overview the daily time series on the total income and expenditure. Figure 4a shows that the head’s income had been paid twice monthly: the semi-monthly wages were paid on the middle (14–16th) and last day (30th or 31st). [^42] The head’s income was the dominant way in earnings because the breadwinner households had been typical at that time (Panel B in Table 3). Importantly, Figure 4a indicates that the households expensed much of their income just on the payday and spent the rest of the payment little by little until the next payday. In other words, the expenditures had never clearly increased just before the paydays. This means that the households had made their consumption decisions (schedules until the next payday) based on the income they obtained. Notice that calendar (semi-)monthly aggregation can not reflect this household behavior because calendar (semi-)month does not place the payday at the beginning of its cell. [^43] For example, a calendar semi-month cell in January, say the cell from 1st to 14th, starts on the 1st but its payday is set on the 14th. This assumes that the aggregate consumption in this semi-month cell is based on the earnings at the end of its cell. This is clearly an unrealistic assumption. Another source of measurement error is the shift of the payday: the payday for the latter half of a month was sometimes set on the first date of the following month. Although this shift did not occur frequently, the calendar month cell shall contain a substantial noise from the shift as it leads to the cells including more (three paydays) and fewer (one payday) payments than normal. Nelson (1994) suggests that this sort of mismatching in the timing of income and expenditure tends to induce the type II error in the test of consumption smoothing.
Adjustment
In order to deal with the mismatch, I aggregated the daily observations between the 12th and 26th for the first half and those between the 27th and 11th of the next month for the second half. [^44] This means that the number of days included in a semi-month ranges from 13 to 16. Note that this heterogeneity in the number of days over different semi-months is a cross-sectional constant. Thus, these differences are entirely captured using the semi-month fixed effects in the regression. This range is based on the fact that almost all of the paydays in the second half of a month measured in the THBS dataset, are 27th–31st. [^45] Note that I do not use the definition of semi-month starting from the 2nd (to the 16th) in each month because this way of stratification places the payday at the end of each semi-month cell. Online Appendix B.7 provides the evidence that while the definition using 2nd in each month as the threshold provides an improved series in income, the expenditure can not chase the income in the right way because the expenses measured under this definition are based on the income in the previous semi-month cell. Again, I confirmed that my preferred definition using the 12th in each month as the threshold offers a clean series of income and expenditure. Figure 4c shows the time-series plots of the income and expenditure from this adjusted semi-monthly panel dataset, whereas Figure 4e illustrates the unadjusted (calender) semi-monthly series using the 15th of each month as the threshold of the semi-month. Clearly, Figure 4e provides very rough measures of income and expenditure due to the measurement errors from the misadjustment of paydays. The most significant improvement in Figure 4c is that the expenditure can chase the income in each month, reflecting the households’ consumption behaviors based on their income measured in the nearly exact timing. Given the improvements in both income and expenditure measures, the net income series shows a much smoother trend in Figure 4d, which are not available under the unadjusted series (Figure 4f). In Section 4.1, I will demonstrate how these measurement errors in the unadjusted dataset attenuate the estimates.
3.4 Data Description
Variables
The THBS budget book has two primary sections: expenditure and income. I digitized records in each category following the categorization used in the official report. Panel A of Table 3 lists the $11$ subcategories for expenditures: food, housing, utilities, furniture, clothes, education, medical, entertainment, transportation, gifts, and miscellaneous. These are used to test the smoothness of the consumptions in each subcategory. To test the roles of risk-coping strategies, I consider five net income categories: savings, insurance, borrowing, credit purchase, and gifts. Net variables are defined as the income minus expenditure in each category; for example, net savings refers to the difference between the withdrawal amount and deposits to savings. I also consider the sales of miscellaneous goods and labor earnings from the family members except for the head to test the contribution of the asset sales and intrahousehold labor-supply adjustments. Panel B of Table 3 summarizes these variables. [^46] Online Appendix B.9 presents the results for the panel unit root tests. The null hypothesis of unit roots in all the panels is rejected at the conventional level for all these variables.
Alterations in family size may shift the household’s preferences (Jappelli and Pistaferri 2017). [^47] In village economies, landholding and asset stocks such as livestock and grain may also be correlated with the household’s risk-coping behaviors (e.g., Udry 1994, 1995). Since this study analyzes urban factory workers’ households, the roles of these assets can be discarded. Fortunately, the archives of the Tsukishima Survey include a set of sheets documenting the changes in the composition of the THBS households (Kazoku idō hyō). [^48] The Ohara Institute for Social Research, Archives of the Tsukishima Survey (THBS, unreleased). I found that the size of households had been stable over time: one household experienced changes in the number of family members due to the birth of a son during the sample periods. This supports the evidence that the preference shifts had rarely occurred among the THBS households. Although family size is included in all regressions to be conservative, it has little effect on the results. [^49] The family size variable shall be omitted in several regressions due to the near-perfect multi-collinearity in the fixed-effects models given its small within variations. Although the number of family members in different age bins is available from the THBS, those family composition variables are also time-constant and are absorbed by the household fixed effect. Panel C of Table 3 shows the summary statistics on the family size variable.
Table 3: Summary Statistics
Panel A: Variables for testing consumption smoothing (in yen)
Semi-monthly Panel Dataset Adj. Monthly Panel Dataset
Level Log Level Log
Mean Obs. Mean Obs. Mean Obs. Mean Obs.
Total expenditure 34.5 289 3.4 289 61.1 162 4.0 162
    Food 10.9 289 2.3 270 19.5 162 2.8 152
    Housing 2.7 289 1.6 134 4.9 162 1.8 118
    Utilities 1.2 289 -0.1 232 2.2 162 0.5 147
    Furniture 0.4 289 -1.6 177 0.7 162 -1.2 126
    Clothes 2.8 289 0.3 248 5.0 162 1.0 148
    Education 0.7 289 -0.6 201 1.3 162 -0.2 135
    Medical 1.7 289 0.2 269 3.0 162 0.7 152
    Entertainment 1.4 289 -0.1 255 2.5 162 0.5 150
    Transportation 0.4 289 -1.3 195 0.8 162 -0.9 125
    Gift 3.1 289 0.5 243 5.6 162 1.2 148
    Miscellaneous 0.5 289 -1.3 222 0.9 162 -0.9 140
Disposable income 37.6 289 3.5 278 67.3 162 4.1 159
Panel B: Variables for testing risk-coping mechanisms (in yen)
Semi-monthly Panel Dataset Adj. Monthly Panel Dataset
Mean Std. Dev. Obs. Mean Std. Dev. Obs.
Panel B-1: Savings, insurance, borrowing, and gifts
    Net savings -0.01 3.97 289 -0.02 5.52 162
    Net insurance -0.03 6.67 289 -0.05 8.97 162
    Net borrowing 0.31 5.03 289 0.61 7.32 162
    Net credit purchase -2.47 11.33 289 -4.67 19.38 162
    Net gifts -2.04 6.11 289 -3.56 9.20 162
Panel B-2: Labor supply adjustments & sales of miscellaneous assets
    Other family member’s earning 3.74 6.11 289 6.72 14.79 162
    Sales of assets 0.02 0.12 289 0.04 0.15 162
Panel B-3: Income variable
Head’s earning 29.43 18.11 289 52.70 30.30 162
Panel C: Family size (people)
Semi-monthly panel dataset
Mean Std. dev. Min Max Obs.
Number of family members 4.3 1.6 2.0 8.0 289
Panel D: Health shocks
Semi-monthly panel dataset
Mean Std. dev. Min Max Obs.
Number of days with illness (head) 0.1 0.8 0 11 254
Number of days with illness 0.2 1.7 0 16 254
(the other family members)
Notes: Panel A: The summary statistics of the variables used to test the consumption smoothing for the 33-panel units in the semi-monthly and adjusted monthly panel datasets are shown in this panel. Disposable income is the total income minus tax payments. Panel B: This panel shows the summary statistics of the variables used to test the risk-coping strategies for the 33-panel units in the semi-monthly and adjusted monthly panel datasets. Each net income variable is defined as the difference between income and expenses. “Net savings” refers to the difference between the amount withdrawn and the total deposits. “Net insurance” refers to the amount received from insurance minus the expenses paid for insurance. “Net borrowing” refers to the amount of borrowing minus the debt payments. “Net credit purchase” refers to the amount of credit purchases minus credit redeemed. “Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts. “Other family member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles. Panel C: The summary statistics of the family size variable for the 33-panel units in the semi-monthly panel dataset are listed. The statistics for the adjusted monthly panel dataset are not reported because they are materially similar to those for the semi-monthly panels. Panel D: The summary statistics of the number of days with illness for family members are listed. Since both are lagged variables, the initial time period in all the units is omitted. Similarly, a household with two semi-month observations is excluded. Similarly, . Sources: Created by the author from the THBS sample.
Trend
Next, I illustrate the time-series figures of the THBS dataset to overview the trend in income and expenditure. Figure 4c shows a weak increasing trend in the average income, which is consistent with the overall macroeconomic trend in the nominal income during this period. This trend in earnings was, however, offset by the steep inflation under the expansions after the war (Section 2.1). In Figure 4d, the deviation of expenditure from income in January must capture the expenses on the goods for the New Year events. The gaps in June to the first half of July may reflect the payments of bonuses. The humps in March may reflect the preparation for the new fiscal and academic year from April, suggesting a higher proportion of temporary income from withdrawals and loans. These support the evidence that the budget data from the THBS households can capture a large part of the macroeconomic trends.
Idiosyncratic Shocks
[…]
Considering this, I first provide the overall results from the specification using the changes in earnings as the idiosyncratic income shock in Sections 4.1 and 4.2. I then consider the structure behind the adverse income shocks using the heads’ temporary illness to investigate the households’ consumption and risk-coping behaviors in Section 4.3. Panel D of Table 3 shows the summary statistics on the health shock variables.
Risk Preferences
This study investigates consumption smoothing behaviors among the skilled factory workers’ households in the machinery sector in Tsukishima. This empirical setting is preferable because the different attitudes toward risks among households with different occupations and regions may lead to bias in the standard risk-sharing regressions (Schulhofer-Wohl 2011; Mazzocco and Saini 2012). To assess the homogeneity of the risk preference among THBS households, I tested whether the earnings of the heads who worked in a few specific factories or sectors showed different responses to the aggregate shocks in the spirit of Schulhofer-Wohl (2011). Specifically, I consider the heads who worked in large-scale or governmental factories to have potentially different risk preferences because those workers might have lower risk tolerance than those who worked in smaller enterprises. I confirmed that the sensitivities to the aggregate shocks are materially similar in both groups in the statistical sense. [^53] Online Appendix B.10 summarizes the finer details of this analysis. In short, the heads’ earnings are regressed on the aggregate consumption, its interaction term with respect to the indicator variable for the large-scale and government factories, family size variable, and household fixed effects. The results are unchanged if I trim a few semi-months with the smaller number of cross-sectional observations. This result is unchanged if I assume that the heads who worked in the smithing factory or in the non-machinery sectors had different risk preferences. [^54] The skilled workers in the smithing sector might be a business owner (Online Appendix A.6). Similarly, the workers in the non-machinery sectors in Tsukishima, a representative machinery production area, might have had a different, presumably higher, risk-tolerance. This supports the evidence that conditional on the household fixed effects, it is plausible to assume that the THBS households shared similar risk preferences.
4 Empirical Analysis
4.1 Consumption Smoothing
Estimation Strategy
I consider a linear fixed-effects model designed to test the consumption smoothing. [^55] Another empirical specification used to test the consumption smoothing is the first-difference model, including the aggregate consumption to capture the macroeconomic shocks in the economy (Mace 1991). In other words, this alternative model requires the interpretation of the estimated coefficient on the aggregate measure of consumption. Although the consumption of the THBS households tends to capture the macroeconomic trend in the consumption (Section 3.3), those households do not cover entire households in the target economy, Tsukishima. This disturbs the interpretation of the coefficient on aggregate consumption. To be conservative, therefore, I use the two-way fixed-effects model to control for the macroeconomic shocks (Cochrane 1991; Ravallion and Chaudhuri 1997). Online Appendix C.1 provides the conceptual framework and the derivation of my empirical specification. For household $i$ in time $t$, the reduced form equation is characterized as follows:
$\log c_{i,t}=\gamma\log y_{i,t}+\delta x_{i,t}+\mu_{i}+\phi_{t}+u_{i,t},$ (1)
where $c_{i,t}$ is consumption, $y_{i,t}$ is disposable income, $x_{i,t}$ is the family size control, $\mu_{i}$ is the household fixed effect, $\phi_{t}$ is the semi-month fixed effect, and $u_{i,t}$ is a random error term. The household fixed effect captures the time-constant unobservable factors such as permanent income and consumption preference. The time-fixed effect effectively removes the unobservable macroeconomic shocks and trends described in Section 3.4. The estimate of $\gamma$ shows income elasticity, which can range from zero for perfect insurance to one for the absence of insurance. The cluster-robust variance-covariance matrix estimator is used to deal with heteroskedasticity and serial dependency (Arellano 1987).
Results
Table 4: Results of Estimating Income Elasticities
(1) Semi-monthly Panels (2) Adj. Monthly Panels (3) Unadj. Monthly Panels
Disposable income Obs. Disposable income Obs. Disposable income Obs.
Coef. Std. error Coef. Std. error Coef. Std. error
Total consumption 0.358 [0.033]*** 278 0.548 [0.065]*** 159 0.261 [0.054]*** 144
    Food 0.136 [0.030]*** 259 0.436 [0.088]*** 149 0.100 [0.072] 135
    Housing 0.099 [0.415] 134 0.291 [0.243] 118 -0.263 [0.167] 126
    Utilities -0.003 [0.138] 223 0.431 [0.149]*** 144 0.197 [0.333] 133
    Furniture 0.426 [0.223]* 169 0.943 [0.335]*** 123 0.250 [0.339] 114
    Clothes 0.378 [0.190]* 238 0.374 [0.324] 145 0.806 [0.262]*** 135
    Education 0.070 [0.092] 194 0.337 [0.212] 133 0.093 [0.213] 125
    Medical expenses 0.039 [0.066] 258 0.498 [0.196]** 149 -0.128 [0.094] 135
    Entertainment expenses 0.355 [0.109]*** 244 0.484 [0.148]*** 147 -0.049 [0.199] 134
    Transportation 0.278 [0.110]** 187 0.635 [0.190]*** 124 0.362 [0.290] 117
    Gift 0.523 [0.106]*** 234 0.599 [0.115]*** 146 0.622 [0.305]* 133
    Miscellaneous -0.076 [0.147] 214 0.549 [0.184]*** 137 0.246 [0.267] 129
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level. Notes: This table shows the results of equation 1: the regressions of the 11 measures of log-transformed consumption on log-transformed disposable income as well as on the family size control, household-fixed effect, and time-fixed effect. Columns 1, 2, and 3 show the results for the semi-monthly, adjusted monthly, and unadjusted monthly panel datasets, respectively. The estimated coefficients on log-transformed disposable income are listed in the columns named “Coef.”.
Column (1) of Table 4 presents the results from the semi-monthly panel dataset. In the total consumption, the estimated coefficient is $0.358$. This means that a one percentage point decrease in disposable income decreases the total expenditure by $0.36$ percentage point, suggesting that the factory worker households could not thoroughly smooth their consumptions. Among the subcategories, the food, furniture, clothes, entertainment, transportation, and gifts expenditures are not well-insured. The households are likely to cut off (increase) their expenses on these categories when they face adverse (favorable) income shocks. The estimates suggest that the expenses on furniture and clothes are relatively sensitive to the shocks ($0.426$ and $0.378$, respectively), which is consistent with the fact that both categories are luxury categories. Similarly, the estimate for gifts shows a large magnitude ($0.523$), reflecting that the expenses on the others are essentially optional. The estimates for the entertainment and transportation expenses are slightly smaller but statistically significant ($0.355$ and $0.278$). This also makes sense, given that these expenses include family travel and theatergoing. [^56] The workers can enter their factories by walking in Tsukishima. Thus, the costs of commuting to the factory are not usually documented in the budget books. An important finding is that the estimate for food categories is statistically significant but much smaller ($0.136$). Therefore, the households might have had some means to deal with idiosyncratic shocks to smooth their expenses on food. I will return to this point later.
I present the results for the adjusted monthly panel dataset in Column (2) to see whether the income elasticities changed in a longer span. Not surprisingly, the adjusted monthly series yields greater estimates than the semi-monthly series. The estimate for total consumption is now $0.548$, which is roughly $1.5$ times greater than that from the semi-monthly series. Generally, the panels with short-run time bins are more likely to be influenced by unexpected shocks in consumption, which tends to attenuate the income elasticity (Nelson 1994). If this mechanism worked, subcategories that are sensitive (insensitive) to unexpected shocks should be more (less) attenuated. In my result, however, the estimates for the adjusted monthly panels are greater than those for the semi-monthly panels across almost all subcategories. Thus, this suggests that households may face more difficulties in smoothing their consumption in the longer run. The clothes category alone shows a similar but statistically insignificant estimate. This may reflect the cases in which the workers bought their work clothes for the new season. Since work clothes are rather necessary items, this sort of seasonal expense may attenuate the estimate.
A few subcategories turn out to be statistically significantly positive: utilities and medical expenses. The estimate for the utility category is relatively large. Given that the sample period includes the winter season and that the firewood, charcoals, and medicines are relatively expensive, this may reflect the households’ behavior of reducing their expenses on fuels and medicines. By contrast, a few subcategories are not sensitive to the idiosyncratic shocks over different time-series frequencies. The estimates for the housing and education categories are not statistically significant in both semi-monthly and adjusted monthly series. This may reflect that the rent payment and school fees were fixed and paid once per month. [^57] Online Appendix C.2 confirmed this feature in the housing category. In addition, the payments for education shared a very small amount of households’ expenses (Panel A of Table 3).
Overall, I found that the estimated income elasticity is smaller in the semi-monthly than in the adjusted monthly series. In the total consumption, the estimated coefficient is $0.36$ in the semi-monthly panels, while that is $0.55$ in the adjusted monthly panels. One potential explanation behind this difference is that the available sources of temporary income have been limited, and those sources are not useful for compensating for adverse income shocks lasting for at least one month. [^58] The positive income shocks might not be enough to sustain a comfortable lifestyle in the middle run. In other words, while the household could compensate for their income losses for a few weeks, they could not do so for several weeks, presumably because of borrowing constraints and/or a limited amount of liquidity assets. The most noticeable example is the food category, which comprises the largest portion of the expenditure (Panel A of Table 3). The estimate for the food category is $0.14$, which is more than three times smaller than the estimate from the adjusted monthly series ($0.44$). In addition, the estimates for the utilities and medical expenses show a relatively clear tendency. This implies that the households’ expenditures on food, fuel, and medicine are likely to be insured in a few weeks but not over one month. As I will delve into the mechanism behind this behavior in the following subsections, the credit purchases may work to smooth consumptions for these categories in the short run.
Finally, I assess the validity of the aggregation method used in this study. Adjusting the aggregate data to account for the timing of paydays reveals more transparent relationships between income and expenditure than the unadjusted data (Section 3.3). To demonstrate this mechanism, I listed the estimates from the unadjusted monthly panel dataset in Column (3). The estimate for the total consumption is $0.261$, roughly half of the estimate from the adjusted series ($0.548$). In addition, most subcategories show systematically smaller estimates than those from the adjusted data listed in Column (2). These estimates are clearly odd, with several subcategories taking values close to one or large negative values. Those are no longer economically interpretable. [^59] For example, the estimate for housing is now negative. The households paid a fixed rent amount monthly. Given that the rent payday is fixed, for example, the shifts in wage paydays might create cells with less income with a fixed amount of rent. In the case of seasonal consumption categories, the shifts may coincide with the timing of seasonal spending. The THBS households tended to purchase work clothes in March to prepare for the start of the new fiscal year in April. They also spent more on clothing in June, likely to buy summer clothes. The shift of paydays to these particular months may have created a strong but false positive correlation. This means that measurement errors in the unadjusted aggregation lead to strong attenuation bias, which results in false negatives in most of the consumption categories and may cause critical issues in statistical inference for a few subcategories.
4.2 Risk-coping Mechanisms
Preceding results show that the skilled workers’ households in Tsukishima were less likely to smooth their consumption. The estimated elasticities, however, become smaller when I focus on the semi-monthly frequency. Specifically, food expenditure is likely to be insured within at least half a month. The workers might have dealt with their shorter-run shocks relatively well but could not cope with the longer-run shocks. In this section, I test the mechanisms behind their smoothing behavior using several variables on the risk-coping devices.
Estimation Strategy
To test the roles of risk-coping strategies, I employ the empirical specification suggested by Fafchamps and Lund (2003). For household $i$ at time $t$, the specification is characterized as follows:
$r_{i,t}=\kappa\tilde{y}_{i,t}+\eta x_{i,t}+\nu_{i}+\zeta_{t}+\epsilon_{i,t},$ (2)
where $\tilde{y}_{i,t}$ is head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ is the household fixed effect, $\zeta_{t}$ is the time fixed effect, and $\epsilon_{i,t}$ is a random error term. Online Appendix C.1 summarizes the derivation of this specification. I prefer to use the head’s income instead of the disposable income in equation 2 to avoid the potential endogeneity. [^60] For instance, the household’s total income is the outcome of the intra-household responses to the shocks, such as the labor-supply adjustments. I confirmed that the results are largely unchanged if I use the disposable income as the idiosyncratic income shock variable instead of the head’s income (not reported). This is consistent with the fact that the labor-supply adjustments did not occur among the THBS households (Table 5). Despite this, I prefer to use the conservative setting to deal with potential biases and to set a comparable setting to the structural equation used in Section 4.3. See Asdrubali et al. (2020) for a recent study using the head’s income as the shock variable in testing the full risk-sharing. As for $r_{i,t}$, I consider several variables to overview various risk-coping strategies. These include the net income on savings, insurance, borrowing, credit purchases, and gifts. Each net variable is defined as the difference between income and expenses. The estimates of $\kappa$ are therefore expected to be negative since net income becomes negative in good times but positive in unfavorable times. I also consider the earnings of all the family members except for the head and the sales of miscellaneous goods because both intra-household labor supply adjustments and asset sales may play a role in mitigating the idiosyncratic shocks as alternative risk-coping strategies (Udry 1995; Blundell et al. 2016). [^61] Rosenzweig and Stark (1989) show that marriage cum migration in South Indian villages contributed to reducing the variability of household food consumption. Since the THBS focuses on the households living in Tsukishima throughout the sample period, I can discard addressing the ex-ante risk-coping strategy through migration. The estimated coefficients for the sales and labor supply adjustment variables are expected to be negative, reflecting additional earnings when the head’s income is reduced.
Results
Table 5: Results of Testing the Risk-Coping Mechanisms
(1) Semi-monthly Panels (2) Adj. Monthly Panels
Head’s earning Obs. Head’s earning Obs.
Coef. Std. error Coef. Std. error
Panel A: Savings, insurance, borrowing, and gifts (yen)
Net savings -0.043 [0.018]** 289 -0.042 [0.018]* 162
Net insurance 0.001 [0.013] 289 0.015 [0.024] 162
Net borrowing -0.021 [0.027] 289 0.036 [0.036] 162
Net credit purchase -0.128 [0.039]*** 289 -0.017 [0.059] 162
Net gifts -0.052 [0.019]** 289 -0.015 [0.014] 162
Panel B: Labor supply adjustments & sales of miscellaneous assets (yen)
Other member’s earnings 0.001 [0.010] 289 0.038 [0.024] 162
Sales of assets -0.000 [0.000] 289 0.000 [0.000] 162
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level. Notes: This table shows the results of equation 2: the family size control, household fixed effects, and time-fixed effects are included in all the regressions. Columns 1 and 2 show the estimates for the semi-monthly and adjusted monthly panel datasets, respectively. The estimated coefficients on the head’s earnings are listed in the “Coef.” columns. Panel A: Each net income variable is defined as the difference between income and expenses. “Net savings” refers to the difference between the amount withdrawn and the total deposits. “Net insurance” refers to the amount received from insurance minus the expenses paid for insurance. “Net borrowing” refers to the amount of borrowing minus the debt payments. “Net credit purchase” refers to the amount of credit purchases minus credit redeemed. “Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts. Panel B: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles. All dependent variables are in yen.
Table 5 presents the results. Columns (1) and (2) summarize the results for the semi-monthly and adjusted monthly panel datasets, respectively. Panel A shows the results for the net income variables, whereas Panel B shows those for the labor supply adjustments and sales of assets.
The estimated coefficient for the net savings is negative and statistically significant in Column (1). The estimate suggests that a one yen decrease (increase) in the head’s earnings increases (decreases) the net income from savings by $0.043$ yen. The estimate from the monthly series is similar ($0.042$), albeit weakly statistically significant (Column (2)). This is basically consistent with the historical fact that many workers had savings accounts both in postal savings and savings banks at that time (Section 2.3). However, the estimated magnitude is small, suggesting that the withdrawals from savings did not compensate for the earnings losses.
The estimates for the insurance category are close to zero and statistically insignificant in both columns. The budget books indicate that the primary source in the category is the mutual loan associations (mujin). As explained in Section 2.2, these associations’ rules were unsuitable for compensating the idiosyncratic shocks. A suggestive case is that a THBS household won a lottery and got $100$ yen in a monthly cell, which is clearly independent of the worker’s earnings. [^62] The OISR, Archives of the Tsukishima Survey (THBS #5, 17th May 1919). The bidding rule may be used to support a specific member in the association if members have a common belief in the aid. However, most of the incomes from mujin measured in the budget books are less than one yen, which is not the case on average.
For the net borrowing category, the estimated coefficients are close to zero and statistically insignificant in both columns. Section 2.2 shows that pawnshop was the popular lending institution for low-income households. Although THBS households occasionally used pawnshops, they never used them frequently. [^63] The OISR, Archives of the Tsukishima Survey (THBS #2; 19; 22). This makes sense because skilled factory workers can not classified as the low-income working class.
The estimated coefficient for the net income from credit purchases is negative and statistically significant in Column (1). The estimate suggests that a one yen decrease (increase) in the head’s earnings increases (decreases) the net income from credit purchases by $0.13$ yen. This magnitude is larger among the other categories. According to the THBS budget books, the credit purchases were mostly used to buy rice, fish, vegetables, and fuels (firewood and charcoal). Since rice, a staple food, is usually an expensive item, credit purchases might have played an important role among these households. The credit purchases were occasionally used in the call doctor and pharmacy. This is consistent with my baseline result for medical expenses in the semi-monthly series (Column (1) in Table 4). While this provides suggestive evidence that credit purchase is the primary risk-coping strategy, Column (2) shows that the credit was less likely to work in the longer run. This seems to be reasonable because borrowers could make monthly or semi-monthly payments but did not borrow for more than a month (Section 5). Instead, the result for the net savings category suggests that precautionary savings have helped mitigate income declines lasting at least one month due to idiosyncratic shocks. Given the magnitudes, however, it is important to note that it could only partially compensate for the idiosyncratic shocks.
The estimates for the net income from gifts are negative and statistically significant in Column (1). The budget books tell that the gifts included seasonal gifts, celebration money, and return gifts. Among them, the return gifts can be regarded as mutual aid from personal networks, which may be reflected in this result. According to an official report of Tokyo, however, the share of average monthly gifts in total income was approximately $3$% (Social Affairs Division 1925, pp. 58-59). In my adjusted monthly panels, this share is calculated to be roughly $4$%. In this light, while the gifts comprised a few percentage points of the monthly income, they may not have provided sufficient aid for the households facing economic hardship. In fact, the estimate is no longer statistically significant in the adjusted monthly series (Column (2)).
Finally, Panel B tests the roles of the intra-household labor supply adjustment and sales of miscellaneous assets. The estimated coefficients in both columns are close to zero and statistically insignificant. Although sewing was the most popular side job among THBS wives, their earnings were too small to replace the lost income from their breadwinning husbands. [^64] The OISR, Archives of the Tsukishima Survey (THBS #1, 13th January 1919). In addition, I find that the sales of miscellaneous assets such as newspapers, bottles, and clocks did not work as the risk-coping.
To summarize, I find that credit purchases have played a role in mitigating the idiosyncratic shocks. The results also indicate that the savings institutions and personal networks might have been used as risk-coping institutions among the workers. However, these were basically available for short-term, say at most half a month. The estimated magnitudes indicate that the workers could only partially compensate for the temporary losses of income by using these institutions, which is indeed consistent with the findings on the partial risk-sharing in Section 4.1. Overall, the results are aligned with the descriptives from the historical documents on the financial institutions available in Tokyo City (Section 2.3). An interesting finding, which is not explicitly described in the historical sources, is the possibility that the credit purchases in the local society and the informal gifts from personal networks might have provided some aid to the factory workers.
4.3 Event-Study Analysis: Response to the Advers Health Shocks
The foregoing analyses used the within variations in the earnings for estimating the income elasticity, meaning that the identification depends both on the adverse and favorable idiosyncratic shocks. While it is useful to illustrate the overall responses of the households to the shocks, it does not offer a clear view of the pathways behind the risk-coping behaviors against the adverse idiosyncratic shocks. To clarify this point, I next focus on the temporary illness of the heads that cause temporary income losses.
Estimation Strategy
Figure 5: Number of Influenza Deaths in Tokyo between 1918 and 1920
Notes: This figure shows the time-series plots of the monthly death counts from influenza (ryūkōsei kanbō) in Tokyo prefecture between January 1918 and December 1920. The death statistics used in this figure were collected based on the comprehensive registration system. Source: Created by the author using the 1918, 1919, and 1920 editions of the Statistics of Causes of Deaths of the Empire Japan (Statistics Bureau of the Cabinet 1921; 1922; 1923).
Figure 5: Number of Influenza Deaths in Tokyo between 1918 and 1920
I begin my analysis by testing whether the illness of the head impacted the earnings. To do so, I decided the exact timing and duration of illnesses of the family members using the Kazoku idō hyō (Section 3.4). Figure 5 shows the number of monthly death counts from influenza (ryūkōsei kanbō) in Tokyo between 1918 and 1920. There appears to have been a moderate spike in early 1919, which is partially captured in the THBS period. In fact, a few THBS budget books suggest that the family members had caught the flu during this period. Five households experienced the illness during the sample period, and the illness affecting the heads was concentrated between February and May 1919, when flu exposure was most prevalent. [^65] Among four cells with health shocks on the heads, two are observed in February, during the peak of the epidemics (Figure 5), and the others are in March and May. In addition, the average duration for the heads is $5.8$ days (Std. Dev. $=3.6$), which is consistent with the fact that Type A influenza causes fever for less than $6$ days (Uyeki et al. 2022). Therefore, it is plausible to see that the influenza epidemics largely led to illnesses during this period, providing a quasi-experimental setting in analyzing to what extent the adverse idiosyncratic shock driven by illness influences the households’ consumption behaviors.
Earnings in a semi-month are based on the work done during the previous semi-month. This means that the head’s earnings shall mainly be influenced by the number of days suffering illness in the previous semi-month cell. Table 6 presents the results from the regression of the head’s earnings on the number of illness days in the previous semi-monthly cell. While the instantaneous health shock has little impact on the head’s earnings (Column 1), the lagged health shock has a statistically significant effect on his earnings (Column 2). This result is unchanged if I include both variables in the same regression in Column 3, supporting the evidence that the health shocks in the previous cell reduced the earnings of the head. The estimate in Column (2) suggests that a one-day increase in the illness decreases the head’s semi-monthly income by roughly $2.4$ yen. Notably, this is close to the average daily wage rate of skilled factory workers (Online Appendix B.6). Thus, this may reflect the reductions in their salary due to the absence. Finally, Column (4) adds the health shocks of the other family members into the specification of Column (3). The estimated coefficients on these placebo variables are statistically insignificant, supporting the robustness of the result in Column (3). These results confirm that the head who suffered illness lost earnings. [^66] Given that the average number of days suffering the illness is roughly $2.0$, the temporary health shock loses more than $5$ yen in that month, accounting for roughly $10$% of the average monthly income. Although this is a rough estimate and the losses shall depend on the severity of the shocks, my result at least shows that the household income had reacted to the adverse health shock among the factory workers’ households.
Table 6: Results of Testing the Impacts of Adverse Health Shockson the Head’s Earning
Dependent Variable: Semi-monthly Earnings of the Head
(1) (2) (3) (4)
Health shocks on:
  The head [level] $-0.061$ $-0.972$ 1.319*-1.319*- 1.319 *
$[0.679]$ $[0.699]$ $[0.666]$
  The head [lag] 2.371***-2.371***- 2.371 * * * 2.868***-2.868***- 2.868 * * * 2.994***-2.994***- 2.994 * * *
$[0.644]$ $[0.734]$ $[0.765]$
  The others [level] $0.889$
$[1.275]$
  The others [lag] $3.675$
$[3.312]$
Observations 287 254 254 254
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level. Notes: This table shows the results from the regressions of the head’s earnings on the health shock variables using the semi-monthly panel dataset. A household with two semi-month observations is excluded from the sample in all the regressions because the lagged health shock variables are included. The family size variable, household fixed effects, and time-fixed effects are included in all the regressions. “Health shocks on the head” refers to the number of days that the head suffered illness. “Health shocks on the others” refers to the number of days that family members, excluding the head, were ill. “[level]” indicates the variable of the head’s earnings in the same semi-month cell. “[lag]” indicates the lagged variable (i.e., the shocks measured in the previous semi-month cell). Although correlation coefficient between the level and lag of the health shock variables for the other family members is substantially high (Pearson’s coefficient $=0.99$), including eather of the two variables does not change the finding on the neglisible impacts of the health shocks on the other family members.
Following the evidence, I characterize equation 1 as a structural consumption smoothing equation by using the first-stage reduced form equation as follows:
$\log y_{i,t}=\beta h_{i,t-1}+\psi x_{i,t}+\varpi_{i}+\tau_{t}+e_{i,t},$ (3)
where $h_{i,t-1}$ is the number of days of head’s illness in the previous semi-month cell, $\varpi_{i}$ is the household fixed effect, $\tau_{t}$ is the semi-month fixed effect, and $e_{i,t}$ is a random error term. For the risk-sharing mechanisms analysis under equation 2, the log-transformed disposable income in equation 3 ($y_{i,t}$) is replaced with the head’s earnings ($\tilde{y}_{i,t}$). The health shock variable is plausibly exogenous because the illnesses were from the influenza epidemics, as explained above. In addition, there is a correlation between the health shock variable and the head’s earnings at a statistically valid level (Table 6). Both support the evidence that the exogeneity of the instrument and relevance condition basically hold under this just-identified system.
Results: To what extent do illnesses affect consumption?
Table 7: Results of Estimating Income Elasticities:Smoothing Consumption against the Adverse Health Shocks
Semi-monthly Panel Dataset
(1) Reduced-form approach (2) Instrumental variable approach
Head’s illness Obs. Disposable income Obs. First-stage
Coef. Std. error Coef. Std. error $F$-statistics
Total consumption -0.118 [0.015]*** 254 0.543 [0.076]*** 245 17.6
    Food -0.075 [0.017]*** 236 0.355 [0.075]*** 227 18.4
    Housing -0.017 [0.070] 126 0.118 [0.477] 126 7.6
    Utilities -0.051 [0.066] 199 0.201 [0.331] 192 12.4
    Furniture -0.157 [0.106] 156 0.761 [0.801] 150 6.2
    Clothes -0.339 [0.050]*** 219 1.233 [0.222]*** 211 65.4
    Education -0.099 [0.039]** 176 0.373 [0.153]** 170 45.9
    Medical expenses -0.002 [0.028] 236 -0.001 [0.142] 227 18.4
    Entertainment expenses 0.019 [0.046] 224 -0.222 [0.501] 215 13.1
    Transportation -0.100 [0.094] 170 0.436 [0.510] 162 20.5
    Gift -0.325 [0.093]*** 212 1.567 [0.797]* 204 17.8
    Miscellaneous -0.075 [0.027]*** 195 0.343 [0.101]*** 189 17.3
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level. Notes: Column 1 shows the results of equation 1, in which the log-transformed disposable income is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell) as the idiosyncratic shock variable. The dependent variables are the 11 measures of log-transformed consumption, and all the regressions include the family size control, household fixed effects, and time fixed effects. The estimated coefficients on the health shock variable are listed in the sub-column named “Coef.”. Column 2 shows the results of equation 1, in which the log-transformed disposable income is regarded as the endogenous variable. The head’s illness is an instrumental variable (i.e., excluded exogenous regressor) for the first-stage reduced-form equation (3). The estimated coefficients on the disposable income are listed in the sub-column named “Coef.”. The first-stage $F$-statistics on the health shock variable are listed in the final column.
I first consider a reduced-form specification using the head’s illness as the exposure variable in equation 1 in the spirit of Cochrane (1991). Column 1 in Table 7 shows the results. Overall, the coefficients on the exposure variable are negative, supporting the evidence that adverse health shocks on the heads reduce consumption levels. The estimate for the food subcategory is relatively small. The estimates for the clothing and gifts categories show a relatively large magnitude. These are in line with my baseline results in Table 4. The estimate for the education category is now statistically significantly negative. This may reflect the reduction of allowances for children in response to income losses due to health shocks.
Next, I explicitly model the structure behind the relationship between the health shocks and earnings using the reduced-form equation of 3. If the negative impacts of the health shocks on consumption were derived from the income losses, the result from the structural estimation approach is similar to that in Column (1). The estimates from the structural equation are listed in Column (2) of Table 7. The results are materially similar to the reduced-form results in Column (1). The estimates are systematically larger than those reported in Table 4. While this result shall reflect the bias reduction thanks to the instrumental variable estimation, it is also important to remember that this system is designed to reduce the bias by specifying the channel from the adverse (instead of favorable) income shocks. For example, expenses on the allowance for children could be easily cut during unfavorable times but may not respond to favorable shocks. Similarly, the expenses for furniture are likely to be cut, but the households might not buy those frequently, meaning that there are no sufficient cases for reacting to the rare health shock events. [^67] In other words, the furniture may have been purchased only when the households had extra income. This may partly reflected in the smaller value of the first-stage $F$-statistics in the furniture subcategory. [^68] Note that the logarithmic transformation leaves the cells with expenses on the furniture. Thus, the reduced-form equation captures the relationship between the health shocks and earnings in the cells with any expenses on the furniture. While the first-stage $F$-statitics is $35.7$ in the linear-linear model, the second-stage result is robust against the functional form assumption. This may reflect the asymptotic theoretical advantage in the panel data model under nearly weak instruments case (Staiger and Stock 1997; Cai et al. 2012; Baltagi et al. 2012). The same argument is applicable to the housing category. A similar interpretation can be applicable to the transportation category because the family travels did not happen frequently. These cases may be another explanation for the difference in the results from my baseline results in Table 4.
While the estimates for clothes and gifts categories now exceed one, the null of equality to the population mean of one is not rejected for both cases. Although the large estimates may be partly due to the operation under the instrument based on the rare health shock events, the findings are materially similar to my baseline results: both categories are sensitive to the shocks as those are dispensable categories.
Results: How did the households compensate for the losses from the ill?
Table 8: Results of Testing the Risk-Coping Mechanisms:Adverse Health Shock as the Idiosyncratic Shock
Semi-monthly Panels
(1) Reduced-form approach (2) Instrumental variable approach
Head’s illness Obs. Head’s earning Obs.
Coef. Std. error Coef. Std. error
Panel A: Savings, insurance, borrowing, and gifts (yen)
Net savings 0.397 [0.318] 254 -0.167 [0.117] 254
Net insurance -0.168 [0.204] 254 0.071 [0.100] 254
Net borrowing 0.116 [0.261] 254 -0.049 [0.103] 254
Net credit purchase 1.294 [0.674]* 254 -0.546 [0.203]*** 254
Net gifts 0.181 [0.177] 254 -0.077 [0.090] 254
Panel B: Labor supply adjustments & sales of miscellaneous assets (yen)
Other member’s earnings 0.001 [0.081] 254 -0.001 [0.034] 254
Sales of assets 0.002 [0.002] 254 -0.001 [0.001] 254
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level. Notes: Column 1 shows the results of equation 2, in which the head’s earning is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell) as the idiosyncratic shock variable. The family size control, household fixed effects, and time-fixed effects are included in all the regressions. The estimated coefficients on the head’s illness are listed in the subcolumn named “Coef.”. Column 2 shows the results of equation 2, in which the head’s earning is treated as the endogenous variable. The head’s illness is an instrumental variable (i.e., excluded exogenous regressor) in the first-stage reduced-form equation. The first-stage result is identical to the result listed in Column (2) of Table 6 with the first-stage $F$-statistics of $13.6$. The estimated coefficients on the head’s earning are listed in the subcolumn named “Coef.”. Panel A: Each net income variable is defined as the difference between income and expenses. “Net savings” refers to the difference between the amount withdrawn and the total deposits. “Net insurance” refers to the amount received from insurance minus the expenses paid for insurance. “Net borrowing” refers to the amount of borrowing minus the debt payments. “Net credit purchase” refers to the amount of credit purchases minus credit redeemed. “Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts. Panel B: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles. All dependent variables are in yen.
I test the risk-coping strategies for mitigating adverse health shocks in Table 8. Column 1 lists the reduced-form results for the risk-coping strategies in the same manner in Table 5. The estimated coefficients on the net income variables are positive in most cases, indicating the households’ compensating responses against the shocks. The estimate for the credit purchases category is weakly statistically significant, with the largest magnitude. It suggests that the one-day increase in the head’s illness increased the net credit purchases by $1.29$ yen, which accounts for roughly half of their average daily wages ($1.29/2.56$). Column 2 presents the estimates from the structural equation using the head’s earnings as the endogenous variable. A similar but much clearer result is obtained: the estimate for the credit purchase category is roughly four times greater magnitude than that reported in Column (1) of Table 5. The estimate indicates that a one yen decrease in the head’s earnings increases the net income from credit purchases by $0.55$ yen. This magnitude is consistent with the direct effect of the health shocks on the net credit purchase estimated in Column (1) of Table 8.696969This means that $2.56\times 0.55=1.4$ yen, which is close to $1.3$ yen in Column (1). The estimates for the savings and gifts categories are still negative but now statistically insignificant, supporting the evidence that credit purchase is the primary risk-coping strategy for dealing with health shocks. Since it shall take more time (steps) from withdrawing and receiving gifts to use the money at the retailers, it is likely that the households prefer to use credit purchases.
Overall, the event-study analysis using adverse health shocks confirms baseline findings and provides a sharper illustration of consumption smoothing behaviors in the case of adverse idiosyncratic shock events. In short, my results show that the unique available response among the THBS households to the idiosyncratic income shock was to use credit purchases. The available compensation via the credit purchases accounts for roughly half of their temporary income losses, which partially contributes to smoothing their food consumption in the short-run.
5 Credit System in Local Economy
Foregoing results shed new light on the roles of credit institutions among skilled workers’ households during historical industrialization. Since the credit purchases are provided by local retailers, this also suggests the importance of the relationships between the workers and retail shops. To the best of my knowledge, however, there is little historical work on credit purchases in prewar Japan. Hence, the details of this credit institution are not yet clear: for example, how retailers get the credit information of their customers, whether the contracts are explicit or implicit, and whether the lenders enforce penalties for the debtors.
Table 9: Customers, Sales Methods, and Competitiveness in the Retailers in Tokyo City
Selling Item Number of Share of sales Social Class of Sales Method (%) Number of
stores for regular the Regular Customer (%) Cash Credit Monthly competitors
reported customer (%) Upper Middle Other purchase installment reported
White Rice 101 73.1 16 45.7 38.2 26.9 71.5 1.7 8.5
Fish 59 64.6 6.8 57.6 35.6 43.9 56.1 0 5.0
Greengrocer 66 45.4 9.4 40.4 50.2 65.8 34.2 0 6.2
Firewood and charcoal 100 51.1 15.7 50.9 33.4 20.4 78.4 1.3 8.6
Notes: This table summarizes the statistics on the sales in the $326$ retail shops selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935. The number of stores reported the share of regular customers (column 2) are $90$, $56$, $61$, and $93$ in the white rice, fish, greengrocer, and fuel (firewood and charcoal) categories, respectively. The “share of sales for regular customers” is defined as the percentage of sales for regular customers in the annual total sales in 1935. The definition of the social classes (“Upper”; “Middle”; “Other”) is not exactly defined in the report. The share of sales method is based on the percentage of sales using each method in the annual total sales in 1935. The number of competitors indicates the number of peers within a three-cho, approximately $327$ meters, radius. The department stores and retail markets are not included in this survey. Source: Tokyo City Chamber of Commerce and Industry 1937, pp. legend; 20–21; questionnaire.
The information on the credit purchase is considerably personal, as in the case of state-contingent loans in the rural credit market (Udry 1994), and can hardly be documented. Therefore, systematic information on credit purchases in the prewar period is unavailable. Fortunately, however, a survey report on retail shops in 1935 published by the Tokyo City Chamber of Commerce and Industry remained. Table 9 summarizes the information on the sales in those retail shops of rice, fish, fruit and vegetables, and fuels. [^70] This survey covers 939 shops and was conducted on 31st December 1935. I digitized the statistics for all the shops selling the representative items (rice, fish, fruit and vegetables, and fuels) in the credit purchases. The survey instrument includes detailed questions regarding sales and purchases, making it the most comprehensive of the available documents (Tokyo City Chamber of Commerce and Industry 1937, pp. legend, questionnaire). Online Appendix Table C.1 decomposes the statistics listed in Table 9 by the scale of businesses. I confirm that the main findings are materially similar. First, many annual sales are from their regular customers, most of whom are from the non-upper classes. For example, $73$% of annual sales in rice retailers are from their regular customers, and more than $80$% of them are from non-upper classes, including the factory workers. Second, and most importantly, credit purchase is the popular sales method, especially among rice and fuel retailers: more than $70$% of the sales were done with credit in both shops. Although cash was also used frequently among greengrocers, $34$% of the sales were settled using credit. This supports the evidence that credit purchase was a popular payment method, and regular customers settled a large part of it. Although these statistics are from the mid-1930s, this confirms the possibility that the retailers were engaged in credit transactions in the context of a long-term relationship with their regular customers. The final column of this table further provides suggestive evidence behind this. The survey asked retailers whether the sales area was competitive or not, and if yes, they were asked to report the number of competitors within a three-cho (roughly $330$ meters) radius. As shown, rice and fuel stores have 2-3 more competitors than the retailers selling fish, fruit, and vegetables. Given the relatively high utilization of credit in rice and fuel retailers, they may have responded to competition by allowing their customers to trade on credit.
Figure 6: Relationships between Average Credit Purchases and Payments
by Different Time-series Frequencies
(a) Daily panel dataset:
purchase and payment
(b) Daily panel dataset:
net purchase
(c) Semi-monthly panel dataset:
purchase and payment
(d) Semi-monthly panel dataset:
net purchase
Note: Figure 6a, and 6c illustrate the time-series plots of the average daily and semi-monthly credit purchase and payment, respectively. Figure 6b and 6d illustrate the time-series plots of the average daily and semi-monthly credit purchase minus payment. Figures 6a and 6b show the daily data between 12th January and 11th November 1919. Figures 6c and 6d show the semi-monthly series calculated using the daily data between 12th January and 11th November 1919. All figures are based on the 30 THBS households that used any credit purchases. Source: Created by the author using the THBS sample.
Figure 6: Relationships between Average Credit Purchases and Payments
by Different Time-series Frequencies
Next, I look into consumers’ credit purchase behavior using the THBS dataset. Most of the THBS households ($29/33$) used any credit purchases during their sample period, supporting the evidence that this was a common practice among factory workers in Tsukishima. Figure 6a illustrates the average daily credit purchases and payments. Generally, there is a peak per semi-month in payments, which reflects the households’ behavior of making the payments for the credit purchase on or soon after their paydays. The purchases series shows a similar tendency but with lower spikes and larger variance. This indicates that the households purchased goods by a certain amount around their paydays and then used it when they needed unconservable food such as fish, fruit, and vegetables. Figure 6b illustrating the net purchases confirms this behavior: the households used the credit on a daily basis and redeemed it near paydays. Figures 6a and 6d show that the average payment exceeds the average purchase in almost all semi-month cells, meaning that many households were able to repay their loans. This is consistent with the findings from the survey report that credit purchases function based on the long-term relationship between retailers and consumers in the local economy.
I then analyze whether the volatility of earnings restricts credit purchases. Since the credit of a worker is unobservable, I use the volatility of earnings as a proxy of credit risk in the following empirical specification:
$I(\overline{CP}_{i}=0)=\varkappa_{0}+\varkappa_{1}CV_{i}+\bar{\mathbf{x}}_{i}^{\prime}\mathbf{\chi}+v_{i},$ (4)
where $I(\cdot)$ is an indicator function, $\overline{CP}$ is the (unit-)average credit purchases, $CV$ is the coefficient of variation for head’s earnings, $\bar{\mathbf{x}}$ is a vector of the (unit-)average family composition variables, and $v$ is a random error term. All variables are calculated using the semi-monthly series. [^71] For example, the coefficient of variation for household $i$ is defined as $CV_{i}=\{\sum_{t}^{T_{i}}(\tilde{y}_{i,t}-\bar{\tilde{y}}_{i,.})/(T_{i}-1)\}^{\frac{1}{2}}/\bar{\tilde{y}}_{i,.}$, where $T_{i}$ indicates the number of time-series observations, and $\bar{\tilde{y}}_{i,.}$ indicates the time-series average of the head’s income. In this cross-sectional model, I include a set of family composition variables to control for unobservable preferences because the household-fixed effect is unable to be modeled as an error component. If retailers allowed households to use credit purchases regardless of the variability of earnings, the estimate of $\varkappa_{1}$ becomes close to zero. In contrast, if retailers did not allow households with high-income volatility to trade on credit, the estimate should be positive and reasonably large.
Table 10: Volatility of Earnings and Use of Credit Purchases
Dependent Variable: Indicator Variable for
Non-CP Users Non-CP + Low-CP Users
(1) (2) (3) (4) (5) (6)
CV of Head’s Earnings $0.003$ $-0.002$ $-0.027$ $0.019$ $0.008$ $-0.005$
$[0.012]$ $[0.014]$ $[0.050]$ $[0.013]$ $[0.017]$ $[0.049]$
Sample mean of the DV $0.12$ $0.12$ $0.12$ $0.24$ $0.24$ $0.24$
Model LPM LPM Probit LPM LPM Probit
Number of THBS households 33 33 33 33 34 33
Family Size Yes Yes Yes Yes Yes Yes
Children aged 6–12 (%) Yes Yes Yes Yes Yes Yes
Children aged 13–16 (%) Yes Yes Yes Yes Yes Yes
Adults aged 17+ (%) Yes Yes Yes Yes Yes Yes
Head’s age No Yes Yes No Yes Yes
Head’s earnings No Yes Yes No Yes Yes
Zero slope ($p$-value) 0.467 0.718 0.336 0.515 0.519 0.451
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Robust standard errors are reported in brackets. Notes: The dependent variable in Columns (1)-(3) is the indicator variable for the households who did not use any credit purchases during the sample period (non-CP users). The dependent variable in Columns (4)-(6) is the indicator variable for the non-CP users and households less than the first quartile in the sample average of credit purchases. Columns (1)-(2) and (4)-(5) use the linear probability model (LPM), whereas Columns (3) and (6) use the Probit model. “CV of Head’s Earnings” indicates the coefficient of variations in the head’s semi-monthly earnings. The sample mean of the CV is $5.76$ (Std. Dev. = $5.72$). All the regressions include the family composition variables: average family size, share of children aged 6–12, share of children aged 13–16, and share of adults aged 17+ (share of children aged 0–5 is used as a reference group). The average head’s age and average head’s earnings are added to the control variables in Columns 3 and 6. $F$-statistics $p$-values for the null of the zero-slope hypothesis are reported in Columns (1)-(2) and (4)-(5). $\chi^{2}$-statistics $p$-values for the null of the zero-slope hypothesis are listed in Columns (3) and (6). The standard errors for the LPM are based on the HC2 estimator proposed by Horn et al. (1975).
Column (1) of Table 10 presents the result for equation 4. The estimated coefficient is close to zero and statistically insignificant. A potential concern herein may be the correlations between the volatility of earnings and the head’s age and earnings level. Column (2) shows the result from the extended model, including the head’s age and earnings in the controls, which shows the robustness of the baseline estimate in Column (1). Another potential issue is the functional form assumption in the linear probability model. Column (3) then lists an alternative non-linear model, including the full set of control variables used in Column (2). It confirms that the estimate is not statistically significantly different from zero. Next, I changed the threshold for the dependent indicator variable to consider a wider range of households using a relatively smaller amount of credit purchases, say less than the first quartile of the average credit purchases. Columns (4)-(5) summarize the results in the same column layouts as Columns (1)-(3), confirming that the estimates are still close to zero in all the specifications. Finally, the final row in this table shows that the null of the joint zero slope hypothesis is strongly rejected in all the regressions. Overall, there is no apparent trend in the characteristics of households to use credit purchases. This suggests that retailers may not have put weight on the volatility of earnings when trading with factory workers on credit. This is consistent with the retailers’ motivation to secure their local customers discussed earlier, and with the fact that factory workers commonly used credit purchases.
The results so far indicate that retailers have been generous in the earnings volatility risk. This, in turn, implies that retailers may factor risk in the form of interest into their loans or retail prices. Unfortunately, there are no systematic statistics on the interest in credit purchases. Thus, I calculate the risk premiums backward by differencing payments and purchases. This backward calculation technique requires information both on the price and type of retailers. However, budget books usually document price-by-item information but not price-by-retailer information. This is problematic because the same items were often sold in different types of retailers: for example, eggs were sold in greengrocers (yaoya) in many cases but sometimes sold in variety stores (aramonoya). Fortunately, I found that a specific month in a budget book has documented the price-by-retailer information. [^72] The OISR, Archives of the Tsukishima Survey (THBS #4, February 1919). According to this budget book, the reimbursements were made semi-monthly, as seen in Figure 6b. They bought various goods in greengrocers, fish stores, and variety stores between 1st and 14th February 1919 and paid exactly the same amounts on the 15th. In the latter half of February, the household paid the same amounts to greengrocers and variety stores on 28th, but less amount to the fish stores. Since the budget book for March does not have information on the type of retailers, it is difficult to know exactly how the reimbursement shortfall was handled. However, this case suggests that the risk premiums may have been included in the retail prices but not in the form of interest.
To summarize, the available documents show that retailers competed with other retailers in the same sales area. Under this condition, they provided regular consumers with the opportunity to trade on credit without interest for at least one month. This ease of accessibility to credit purchases might enable workers to smooth consumption in response to adverse idiosyncratic shocks in the short-run.
6 Conclusion
This study investigates the consumption smoothing behaviors among working-class households in Tokyo. I digitized the daily longitudinal budget survey on the factory worker households conducted in Tsukishima and implemented a systematic econometric method to obtain comparable estimates for income elasticities. The estimated elasticity for total consumption indicates that the factory worker households suffered certain consumption losses when they faced idiosyncratic income shocks. Despite this, I found evidence that while expenses on dispensable categories are likely to be cut in response to the income losses, expenses on indispensable categories are partially smoothed.
The estimated elasticity for total expenditure is in a similar range to those in rural economies in developing countries, although only a few available references exist. According to Ravallion and Chaudhuri (1997), the elasticity for the rural households in Indian villages between 1975 and 1984 was roughly $0.5$. As for the urban households, the estimated elasticity is $0.4$ for Bangkok between 1975 and 1990, which had a similar GDP per capita to 1920 Japan (Townsend 1995). Since my estimate is from a specific occupation, skilled factory workers, one must be careful when making comparisons with these estimates based on the urban samples, including a variety of occupations. At least, however, the factory workers in the historical metropolis of Japan around WWI had been relatively vulnerable to the idiosyncratic income shocks.
A recent study found an income elasticity of $0.39$ for factory workers’ households in Osaka city circa 1920 (Ogasawara 2024). My estimate from the adjusted monthly panels is slightly larger than this estimate. Although the measurement errors in aggregate consumption and income in the monthly panel dataset may explain the discrepancy in these estimates, if taken literally, this implies that the Tsukishima factory workers were more vulnerable than those in Osaka. This indicates that higher levels of urbanization do not necessarily correspond to a greater availability of risk-mitigation institutions. Therefore, inter-city and cross-national comparisons of income elasticities and analysis of the institutional characteristics behind the consumptions there would be another subject of future work.
The finding on the attenuation bias due to the measurement errors is potentially important in that they may prompt a reinterpretation of existing results. Errors in the timing of payroll can result in inconsistencies in the aggregate data for income and expenditure, albeit the household-level studies often use aggregate data. In this case, modifying the aggregation period to align with wage payments can reduce attenuation bias. Even if the aggregation period is unadjustable, clarifying payroll timing and discussing potential bias is useful.
The results from mechanism analyses provide suggestive evidence that credit purchases played an important role in mitigating adverse idiosyncratic shocks on consumption. Although the institutional details of credit purchases should be investigated further, this finding offers a new historical view of the interdependence between consumers and local retailers. Understanding the local retail system would be important in comprehending regional consumption behavior and, in turn, in understanding the improvement in living standards in prewar Japan.
Appendices
Appendix Appendix A Background Appendix
A.1 Developments of the Cities
Figure A.1: Developments of the Large Cities in Japan
(a) Number of people in the large cities
(b) Share of people in the large cities (%)
Note: This figure illustrates the development of large cities in Japan between 1889 and 1920. Figure A.1a shows the number of people in Tokyo, Osaka, Kyoto, and Nagoya cities. Figure A.1b shows the share of people in Tokyo (diamond) and the cities with more than $50,000$ people (circle) in total population in Japan. Sources: Umemura et al. (1983, pp. 303–305); Umemura et al. (1988, pp. 166; 168).
Figure A.1: Developments of the Large Cities in Japan
Table A.1: City Characteristics:Tokyo, Osaka, Kyoto, and Nagoya
Name of cities
Tokyo Osaka Kyoto Nagoya
Panel A: Population and Workers
Population 2,173,201 1,252,983 591,323 429,997
Male workers 768,633 476,229 189,922 138,275
Panel B: Industrial Structure (share in male workers)
Agriculture 1.0 0.8 2.2 1.8
Fisheries 0.1 0.1 0.0 0.4
Mining 0.4 0.3 0.2 0.2
Manufacturing 44.5 45.6 49.0 45.8
Commerce 32.4 34.0 32.3 29.6
Transport 7.6 10.6 5.5 8.0
Public service and professions 11.3 6.7 9.0 11.3
Housework 0.3 1.8 0.1 0.0
Other industry 2.4 0.1 1.5 2.8
Panel C: Local Infrastructures
Number of electrical lights per people in 1921 0.9 0.8 0.9 0.9
Accessibility to modern water supply (%) 78.9 80.5 46.5 39.4
Panel D: Meteorological Features
January 1918 – December 1923
  Average daily temperature 14.1 15.4 14.1 14.6
  Average daily precipitation 4.7 3.9 4.6 4.8
February – April only
  Average daily temperature 8.0 8.8 7.8 8.3
  Average daily precipitation 4.2 3.4 4.1 4.3
Notes: 1. Panel A: The number of people and male workers measured in the Population Census conducted in October 1920 are listed. 2. Panel B: Occupations are classified based on the industrial classification of the 1920 Population Census. Unemployed people are not included in the calculation. 3. Panel C: The number of electrical lights measured at the end of 1921 is divided by the number of people in the 1920 Population Census. The number of houses accessible to water supply measured at the end of September 1920 is divided by the number of total households measured in the 1920 Population Census (%). 4. Panel D: The average daily temperature and precipitation between 1st January 1918 and 31st December 1923 and between 1st February 1920 and 30th April 1920 are listed. Sources: The number of people in 1920 is from the 1920 Population Census (Statistics Bureau of the Cabinet 1928, pp. 186; 214; 222; 225). The number of male workers and their industrial structures are from the Statistics Bureau of the Cabinet (1929a, pp. 20–21; 26–29). The number of houses accessible to the water supply is from the Water Works Association (1922, pp. 69–70). The total number of households is from the 1920 Population Census (Statistics Bureau of the Cabinet 1928, pp. 186; 214; 222; 225). The number of electrical lights is from the Ministry of Posts and Telecommunications (1922, p. 558) The daily temperature and precipitation are from the Japan Meteorological Agency (database).
Figure A.1a illustrates the developments of the large cities in Japan between 1889 and 1920. Table A.1 summarizes the scales, industrial structures, infrastructures, and meteorological features of the large cities.
A.2 Financial Institutions in the Cities
Several financial institutions were available for working-class households in the cities around 1920.
A representative savings institution is the postal savings (yūbin chokin). Postal savings was initially designed as a small savings institution for low-income households, and the share of postal savings to total savings was roughly $10$% in the 1910s (Tanaka 2018, p. 29). However, the number of postal savings accounts per total population in the early 1910s exceeded $20$% (Tanaka 2018, p. 33). [^73] James and Suto (2011) show that the saving rate in prewar Japan was higher than that of the U.S. As shown in Panel A of Table A.2, roughly $1.4$ million people had accounts in Tokyo in 1920. This covers approximately $67$% of the entire population in Tokyo and the average savings amount per capita was $64.5$ yen. In Kyobashi Ward, the average figure for the postal savings accounts was $44.5$ yen at the end of 1919 (Tokyo City Office 1921, pp. 892–893). This amount is less than the average monthly earnings of the skilled-factory workers (Section 3.2). Unfortunately, the statistics on postal savings by depositors’ occupation are unavailable. [^74] There are some statistics for the entire Japan. According to the Postal Savings Bureau (1924, p. 72), the share of depositors in the manufacturing sector was $4.8$% (the share based on the deposit amount is $5.5$%). However, this implies that, on average, postal savings may not provide sufficient temporary income.
Savings bank (chochiku ginkō) is another famous savings institution for working-class households in the 1910s (Ito and Saito 2019, pp 79–81). Panel A of Table A.2 shows that the number of people having savings accounts was roughly $1.46$ million, accounting for approximately $67.3$% of total people in Tokyo at that time. Although this figure is similar to the postal savings, the workers in the manufacturing industry were less likely to rely on the savings banks than the workers in the other industrial sectors. Table A.3 summarizes the number of depositors and the amounts of deposits in savings banks by occupation in Tokyo City (Panel A) and Kyobashi Ward (Panel B). In Kyobashi, the commerce and miscellaneous sectors shared roughly $84.8$% of all the depositors, whereas the manufacturing sector had only $11.5$%. This is unchanged if I use the amounts of deposits to calculate the shares. Tokyo Institute for Municipal Research stated that this tendency captures the fact that savings banks are mainly designed as financial institutions for low-income households (Tokyo Institute for Municipal Research 1925b, p. 81). Thus, skilled factory worker households, like the THBS households, were less likely to use the savings banks. This is consistent with the fact that Tsukishima did not have any savings bank branches, while there were two postal offices. [^75] There were 123 savings banks in Tokyo City in 1922, and 10 of them were in Kyobashi Ward (Tokyo Institute for Municipal Research 1925, p. 76). However, Tsukishima, a representative manufacturing area, did not have savings banks at that time.
Mutual loan association (mujin) is another type of savings institution. The members of this association pay a fixed amount of money several times, and each member withdraws the total amount of payments each time under specific rules. The lotteries or bidding system decided which members could withdraw money each time. This means that the workers could not use this institution to deal with the acute negative income shocks. [^76] This is the most representative mutual loan system called tsūjyō (i.e., common) mujin. There was a different type called sueoki chokin (or tsumitate kai) that people could receive their stakes with interests after a fixed amount had been accumulated for a certain period (Tokyo Legislative Research Association 1915, pp. 3–4). Clearly, both are also not flexible for compensating the income shocks. Instead, it had been used to make a certain amount of money among small business owners and merchants. A report suggests that the workers in the manufacturing sector shared only $10$% of all the mutual loan associations in Tokyo (Tokyo City Office 1935, p. 25). Moreover, the Tokyo Chamber of Commerce stated that the mutual loan association was used for business and rarely for household consumption (Tokyo Chamber of Commerce 1918, p. 121).
Cooperative societies were also available for worker households. Cooperative societies in Tokyo used credit as their payment method: households could buy goods in advance and pay at the end of the month or twice (middle and the end of) a month. The commodity prices were similar to or slightly lower than those in the retail shops (Central Federation of Industrial Associations 1925, pp. 18–19). However, the cooperatives only covered a few citizens. There were 26 cooperative societies in Tokyo prefecture in 1924, and 23 of them were organized by the citizens and workers (Central Federation of Industrial Associations 1925, p. 57–59). Among these 23 cooperatives, two were large societies with roughly 4000 members, and the others were small societies with less than 100 members. The total number of accounts for only a few percentage points in the workers in the city, given that the number of male workers was more than 768 thousand (Table A.1). In fact, most of the THBS households had not used the cooperatives.
An alternative borrowing institution is the credit purchase in retail shops. The THBS households had frequently used credit purchases in various retail shops for rice, fish, vegetables, alcohol, firewood, charcoal, and medicines. The details of the credit purchase are little documented, and the systematic statistics on the credit purchase are also unavailable. This must be because the information on the transactions is personal. Despite this, an available document shows that credit purchases are used in various retail shops (Section 5). The THBS budget data show a consistent tendency for credit purchases to be widely used among factory worker households.
Pawnshops were the popular lending institution among low-income working-class households. Lenders did not need to screen the borrower’s credit, and the borrowers did not worry about incurring heavy debt as the primary articles were inexpensive clothes (Shibuya et al. 1982). In addition, the interest rates were regulated by the Pawnbroker Regulation Act of 1895, and the average redemption rate was substantially high under the lower interest rates. Among $88$ pawnshops in Kyobashi, the average redemption rate was approximately $94$% ($401,222$ cases/$427,265$ cases) in 1920 (Tokyo City Office 1922b, pp. 888–889). The accessibilities were higher than the other lending institutions. The number of pawnshops in Tokyo was $1,334$ in 1918 and $1,261$ in 1919, respectively. The total number of cases was $8,226,883$ and $7,573,406$ each year. These figures are more than three times greater than the entire population in Tokyo City. This means that the accessibility to a pawnshop was relatively high for the workers. As described in the main text, the number of pawnshops in Tsukishima was 13, whereas that of moneylenders was five.
However, the inexpensive article means that pawnshops were used to deal with very short-run necessities only. Among the pawnshops of Kyobashi, for example, the average amount per case was $7.8$ yen, which was $12$% of the monthly earnings of the THBS households. While the pawnshops were the popular lending institutions, they could not offer a sufficient amount of temporary income for relatively large idiosyncratic shocks. According to the statistics on the pawnshop users in Tokyo prefecture in 1923, the share of users in the agriculture, commerce, and manufacturing sectors was 1.7%, 19.5%, and 15.6%, respectively. In turn, the most representative users were the day laborers and workers classified as “miscellaneous”, accounting for $63.4$% (Tokyo Institute for Municipal Research 1926, p. 25). In fact, the Tokyo City Office stated that the pawnshops were mainly used among low-income (poor) working-class households (Tokyo City Social Affairs Bureau 1921, p. 9). This indicates the evidence that the primary pawnshop users were not the skilled-worker households like the THBS households.
Another lending institution is the money lenders (kinsen kashitsuke gyō). Generally, business owners used the money lenders. The interest rates in the money lenders were very high compared with those in the pawnshops (Shibuya 2000, pp. 184; 248). Tokyo Institute for Municipal Research was concerned about the increasing number of unscrupulous companies arrested by the National Police Agency (Tokyo Institute for Municipal Research 1925b, pp. 106–108). Given this fact, while there were two money lenders in Tsukishima, they might have been less likely used by the factory worker households.
Credit unions (shinyō kumiai) are another institution available for working-class households. In Tokyo City, there were 215 unions in 1923. According to the survey of the 184 unions among them, however, the number of union members in the manufacturing sector was only $5,662$ (Tokyo Institute for Municipal Research 1925c, p. 20). This accounts for, at the most, only a few percentage points of the male workers in the manufacturing sector in Tokyo measured in the population census of 1920 (Online Appendix Table A.1).
Table A.2: Financial Institutions in Tokyo, Osaka, Kyoto, and Nagoya Cities in 1920
Name of cities
Tokyo Osaka Kyoto Nagoya
Panel A: Savings Institutions
Postal saving
   Number of depositors 1,462,604 725,642 305,927 172,183
       Population percentage (%) 67.3 57.9 51.74 40.04
   Total amounts (yen) 94,330,789 34,742,000 12,036,439 9,540,725
       Amount per people (yen) 64.50 47.88 39.34 55.41
Savings bank
   Number of depositors 1,461,389 1,648,090 663,831 482,960
       Population percentage (%) 67.25 131.53 112.26 112.32
   Total amounts (yen) 48,682,260 50,674,000 28,495,407 24,014,781
       Amount per people (yen) 33.31 30.75 42.93 49.72
Panel B: Pawnshop
   Number of shops 1,253 867 441 412†
   Number of units lent 6,042,501 2,027,690 1,251,283 319,609
   Amount of loans (yen) 38,667,993 3,785,186 4,424,357 2,983,250
   Amount per loan (yen) 6.40 1.87 3.54 9.33
[…]
Table A.3: Deposits in the Savings Banks in Tokyo City by Occupations
Occupations
Agriculture Commerce Manufacturing Miscellaneous Total
Pane A: Entire Tokyo City
    Number of depositors 76210 550784 166498 619963 1413455
    Amount of deposits 2002124 19112713 6663091 19809591 47587519
    Number of depositors (%) 5.4 39.0 11.8 43.9 100.0
    Amount of deposits (%) 4.2 40.2 14.0 41.6 100.0
Pane B: Kyobahi Ward
    Number of depositors 3809 34665 11664 51481 101619
    Amount of deposits 130781 1218081 450805 1696086 3495753
    Number of depositors (%) 3.7 34.1 11.5 50.7 100.0
    Amount of deposits (%) 3.7 34.8 12.9 48.5 100.0
Notes: This table summarizes the number of depositors and the amount of deposits in the savings banks by occupation. These statistics were measured on 31st December 1919. Panel A and B list the statistics for Tokyo City and Kyobashi Word, respectively. Source: Tokyo City Office (1922a, pp. 864–865).
Figure A.2: Pawnshops and Money Lenders
Notes: This map shows the location of the pawnshops and moneylenders in Tsukishima. The bullseye indicates the pawnshop (shichiya). The black circle shows the money lenders (kanekashi gyō). The white circle shows the rickshaw station (kuruma yado). Source: Department of Health, Ministry of the Interior 1923c, tenth map. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
Figure A.2: Pawnshops and Money Lenders
A.3 Development of Tsukishima
Figure A.3: Tsukishima: Ishikawajima, Tsukudajima, Shin-tsukudajima, and Lots 1–3
Notes: Tsukishima contains Ishikawajima, Tsukudajima, Shin-tsukudajima, and Tsukishima Lots 1–3. This map does not illustrate the northernmost part of Ishikawa jima (see Figure 3 for this part). “Ferries” shows the Tsukishima to Kyobashi Ward sea routes: Tsukuda, Tsukishima, and Kachidoki routes (right to left). “Aioi bridge” links Shin tsukudajima and Fukagawa ward. Source: Department of Health, Ministry of the Interior 1923c, 15th map. The author modified and adjusted the tone using Adobe Photoshop 24.7.0.
Figure A.3: Tsukishima: Ishikawajima, Tsukudajima, Shin-tsukudajima, and Lots 1–3
Tsukishima (transliterated as “moon island”) is a reclaimed land on a sandbar of the Sumida River. Ishikawajima and Tsukudajima (“jima” means island) at the northern end of Tsukishima are the starting point of its history (Figure A.3 for map). The following description of this section is based on Chūō Ward (1994, pp. 122–124).
Ishikawajima was built in 1626 by a boatman, Hachizaemon Ishikawa, who received a fief from the Shogunate. In 1790, the Ishikawajima Prison was built on Ishikawajima. This was a facility for the reintegration of minor offenders and those without lodgings and was called a Ninsoku Yoseba. After Matthew Perry arrived in Japan in 1853, the Shogunate ordered the Mito Clan to build a shipyard on Ishikawajima. A month later, the groundbreaking ceremony was held for a Western-style wooden sailing ship delivered to the Shogunate in 1856. In 1866, the Ishikawajima Lighthouse was built on the west bank of the Yoseba and served to signal the position of ships entering the Nihonbashi direction. In 1870, the Ninsoku Yoseba became a place of imprisonment (Toba) and was renamed the Ishikawajima Jail Station of the Metropolitan Police Department in 1877. The prison was finally moved to Sugamo in Kitatoshima County in Tokyo prefecture, and the penitentiary on Ishikawajima was closed in 1895.
Adjacent to Ishikawajima is called Tsukudajima, where fishermen from “Tsukuda” in Osaka prefecture settled in 1644 after receiving land from the Shogunate (i.e., the island was named after their hometown, Tsukuda). In 1645, the “Tsukuda ferry” began, and throughout the Edo period, Tsukudajima was known as a fishing island for whitefish for the Shogun’s household. In 1863, the Tsukudajima Battery was built on the south side of Tsukudajima, one of 13 batteries used to defend Edo Bay. This battery later became the starting point for the reclamation of Lot 1 of Tsukishima.
In 1884, the Port of Tokyo Waterway Dredging Project was initiated. This was to secure the passage of ships by sweeping away the sediment in the shipping channel, and the Tsukishima landfill was created using the deposit. The first Tsukishima site was completed in 1891, and the “Tsukishima ferry” began the same year. The second site was created in 1894; the Shin (new) tsukudajima was built in 1896. From the west shore of Shin tsukudajima, the island overlooks the Boso Peninsula in Chiba Prefecture, making it a favorite vacation spot for writers. Cultural figures such as Toson Shimazaki and Kaoru Osanai stayed here from 1907 to the early Taisho period. In 1903, the Aioi Bridge was completed, bringing water and electricity to Tsukishima and improving transportation convenience, which contributed to industry development. [^77] Until then, Suzukiiya had been selling water to supply the entire island using a ferryboat. The “Kachidoki Ferry” began to lead to Lot 2 in 1905. In 1913, Lot 3 was finally completed.
A.4 Housings in Tsukishima
Since the Tsukishima Survey included a housing survey (Table B.1), the housing types in the entire island can be determined. Table A.4 summarizes the results of this survey. The description in this section is based on the Department of Health, Ministry of the Interior (1923a, pp. 69–76).
$88$ ($46$%) of the detached homes are in Tsukudajima. Tsukudajima is reported to be “superior to other areas in terms of housing quality, and the percentage of inferior housing is meager”. Shin-tsukudajima had a low percentage of detached homes. The report states that it was indeed “inhabited by unskilled laborers” and “occupies the most inconvenient housing status of all the islands.” Lot 1 was a center of life in Tsukishima and is a cluster of numerous residences, mainly tenements. Lot 2 had a relatively high percentage of single-family and two-family houses and appears to have had a character as a “workers’ residential area” than Lot 1. These indicate that there might have been some geographic differences in income distribution regarding housing type. Overall, however, of the $989$ dwellings in Tsukishima, $799$ ($81\%$) are row houses, and only $190$ are single-family homes. This confirms that Tsukishima was an industrial area with many working families. The workers lived in similar tenements and with similar standards of living. Figure A.4 is a photo showing these tenement houses.
Table A.4: Housing Types in Tsukishima
Tenement (nagaya)
(A) Detached (B) 2 families 3+ families Total (A)+(B) %
2-story 1-story 2-story 1-story
Tsukuda jima 88 32 26 6 23 175 146 83.4
Shin-tsukuda jima 16 27 10 36 59 148 53 35.8
Lot 1 58 62 74 73 200 467 194 41.5
Lot 2 28 27 43 9 92 199 98 49.2
Total 190 148 153 124 374 989 491 49.6
Note: This table summarizes the housing types in Tsukishima surveyed on 30th June 1919. Source: Department of Health, Ministry of the Interior 1923a, pp. 69–76.
Figure A.4: Photographs of the Tenements in Tsukishima
(a) 2-story tenement
(b) 1-story tenement
Note: Figures A.4a and A.4b show the two- and one-story tenements in Tsukishima, respectively. Source: Department of Health, Ministry of the Interior 1923c, figures 62–63; 66.
Figure A.4: Photographs of the Tenements in Tsukishima
A.5 Machinery Factories in Tsukishima
First, I use the results of the Primary School Children Survey included in the Tsukishima Survey to analyze the occupation (industry) of the household heads in Tsukishima. This is a survey of 4th-6th graders at Tsukishima 1st and 2nd Elementary Schools in July 1919. There were 1,173 children in these grades, and the survey included all 1,100 children who resided in Tsukishima. At that time, there were no other elementary schools besides those mentioned above. Thus, this survey is a complete survey for all children in grades 4-6 in Tsukishima. Although it does not cover households without children in grades 4-6, it is undoubtedly a beneficial survey that provides an overview of the households with children that are the main target of this study.
Panel A of Table A.5 shows the industrial structures. The composition ratios are similar to the census-based statistics reported in Panel A of Table 1. It does not specify whether brothers and sisters in the same household were counted twice. Despite this, this similarity confirms that the statistics obtained from this survey are useful in identifying overall trends in family households in Tsukishima.
Panel B of Table A.5 divides the “social classes” of the household heads working in the manufacturing sector into six categories. In the report, large business owners are classified into the capitalist class; small business owners, free enterprise owners, and directors are regarded as the middle class; workers are defined as the working class (Department of Health, Ministry of the Interior 1923a, p. 50). The table reveals at least three things. First, most heads of households are workers, with only 1% of the capitalist class and less than 20% of the middle class. While the definition of “executive” is unclear, “small business owners” must be small factory owners who could have similar earnings to the skilled worker households (Department of Health, Ministry of the Interior 1923a, p. 64).
Table A.5: Industrial Structure and Social Classes in Tsukishima: Evidence from Tsukishima Primary School Survey
Panel A: Industrial Structure
Name of survey Tsukishima Survey
Survey area Tsukishima
Survey month and year July 1919
Survey subject Heads of all the students in grades 4–6
in Tsukishima 1st and 2nd Primary Schools
# of heads % share
Agriculture 1 0.1
Fisheries 13 1.2
Mining 2 0.2
Manufacturing 623 58.6
Commerce 259 24.4
Transport 72 6.8
Public service and professions 55 5.2
Other industry 38 3.6
Observations 1,100 100
Panel B: Social Class of Heads in Manufacturing Sector
Name of survey Tsukishima Survey
Survey area Tsukishima
Survey month and year July 1919
Survey subject All students in grades 4–6
in Tsukishima 1st and 2nd Primary Schools
Tsukudajima Shin-tsukudajima Lot 1 Lot 2 Total
Large business owner 0 0 5 3 8
Small business owner 1 10 49 13 73
Self-employed 0 0 0 0 0
Executive 3 13 28 7 51
Worker 23 94 265 108 490
Unemployed/Unknown 0 0 0 0 0
Total 27 117 347 131 622
% share of workers 85.2 80.3 76.4 82.4 78.8
Notes: This table summarizes the statistics from the complete survey on the 4th-6th graders in Tsukishima1st and 2nd Primary Schools in July 1919. Panel A summarizes the heads’ industries, and Panel B sorts the “social classes” of the heads in the manufacturing sector, respectively. The report does not specify the definition of each “social class”. Source: Department of Health, Ministry of the Interior 1923a, pp. 386–387.
In November 1920, there were 214 factories in Tsukishima, 168 of them (roughly $80$%) were classified as machinery factories. The official report documented that the wooden pattern factories could be included in the machinery sector because they were incorporated into the casting processes (Department of Health, Ministry of the Interior 1923a, p. 388). The share of machinery factories becomes approximately $90$% if I include those 24 wooden pattern factories.
Panel A of Table A.6 uses the manufacturing survey in the Tsukishima Survey to classify the 168 factories by size and industrial sector. First, more than $90$% of the plants are in the three sectors of machinery and equipment manufacturing, can manufacturing, and blacksmithing, with only $15$ plants in the casting sector. 113 ($67$%) of the machinery factories had workers less than 15. Second, many smithing factories have fewer than 15 workers. Hence, almost all of Tsukishima’s smithing industry was considered to be small businesses. [^78] The report suggests that 43 of the 47 smithing factories were the “town factories” or “workplaces”. The definition of “town factory” is not clear in the report (p. 392). Since they were always counted more than the number of factories with 15 or fewer workers, they are likely to include smaller factories operated as small businesses.
Panel B of Table A.6 classifies the types of the 620 workers based on the results of the primary school survey used in Table A.5. There were 455 skilled and 165 unskilled workers. Notably, these 620 workers include non-industrial workers: 490 industrial workers, 13 commercial, 61 in transportation, 16 in public service free trade, 38 in other gainful employment, and 2 in agriculture and mining in total. Since the workers in the industrial sector should include factory workers, workers in the cottage industry, and handicraft workers, the day laborers are considered to be classified into non-industrial workers (in fact, $620$ workers $-120$ day laborers $=490$). Therefore, roughly 90% of factory workers ($455/490$) can be regarded as skilled workers, $65$% of whom worked in factories, $17$% in cottage industries, and $18$% as craftsmen. Since the workers in cottage and handicraft industries are considered to work in small factories, $65$% of these workers might have worked in the middle- to large-scale factories (with 15+ workers).
One must be careful that this survey only covers the heads with primary school children aged 9–11 (i.e., 4th to 6th grades), suggesting that most of these heads are in their 30s. Therefore, the share of skilled workers to the entire workers should be smaller than the $90$%. At least, however, it is plausible to see that most of the Tsukishima workers in their 30s were skilled workers at that time.
Table A.6: Scale of Machinery Factories and Type of Workers in Tsukishima
Panel A: Scale of Machinery Factories
Name of survey Tsukishima Survey
Survey area Tsukishima
Survey month and year November 1920
Survey subject All machinery factories
Machinery and Can Casting Smithing Total
equipment
# of workers $\geq 15$ 22 15 10 8 55
# of workers $<15$ 44 25 5 39 113
Total 66 40 15 47 168
% shares 39.3 23.8 8.9 28.0 100
Panel B: Type of Workers (All Industries)
Name of survey Tsukishima Survey
Survey area Tsukishima
Survey month and year July 1919
Survey subject All students in grades 4–6
in Tsukishima 1st and 2nd Primary Schools
Skilled workers Unskilled workers
Factory workers 295 Factory workers 42
Cottage industry 76 Cottage industry 3
Handicraft industry 84 Day laborer 120
Total 455 165
Notes: Panel A summarizes the number of machinery factories by the scale and type of the factory. Statistics are from the manufacturing census conducted in the Tsukishima Survey (1st November 1920). The threshold of the scale (i.e., 15 workers) is based on the application of the Factory Act enacted in 1916. Panel B sorted the heads by their skill level and job type. Statistics are obtained from the Complete Survey on the 4th-6th graders in Tsukishima1st and 2nd Primary Schools in July 1919. Source: Panel A: Department of Health, Ministry of the Interior 1923a, pp. 386–387; 410–414. Panel B: Department of Health, Ministry of the Interior 1923a, p. 149.
A.6 Works in the Machinery Factories
Tsukishima Survey includes the factory survey that counts the number of machinery factories in Tsukishima. Panel A of Table A.6 summarizes the number of machinery factories by type of plant. The machinery and equipment manufacturing factories shared 39% of all the machinery factories. Typical jobs for workers in machinery and equipment manufacturing plants are finishers (or fitters) and turners. A finisher is “a skilled worker responsible for finishing, assembling, and commissioning machine products” and is “one of the finest skilled workers because it requires skills in all aspects to correct defects in machine making” (Department of Health, Ministry of the Interior 1923a, p. 393). The turner is the “middle axis of machinists,” to which most semi-skilled and skilled workers belong. Becoming a full-fledged finisher and lathe operator takes at least five years. Generally, apprentices enter the workforce around 14–18, but they do not work in the same factory for many years within the apprenticeship system. After learning the skills for 1-2 years, they often move to another factory to improve their skills (Department of Health, Ministry of the Interior 1923a, p. 394). Wages are paid on a piece-work basis or on a time basis. In Tsukishima, a time contract was adopted, and payment was based on the “residual time” method. Under the “residual time” method, “if the work is completed within the contracted time, the remaining time is considered the bonus.” For example, if 12 hours are contracted, and the work is completed in 10 hours, 2 hours are paid as a bonus (pp. 396-397). Wages were paid by dividing one month into two periods (first and second semester). The average work hours are reported to be roughly 10 hours, with occasional all-night work (p. 402). Overtime hours were reportedly around 2 hours per day (Department of Health, Ministry of the Interior 1923a, p. 402).
Next, I summarize labor in the canning factory, which accounts for 24% of all Tsukishima factories. The canning factory is engaged in shipbuilding, bridges, steel frames for construction, and other works, including fabrication and repair of gas reservoirs (p. 406). Workers in a canning factory are called “boiler makers” and are typically involved in bending steel plates and joining them with rivets. Workers are divided into rivetter, holder-up, plater, angle-ironsmith, and caulking. The work is done in groups, with the skilled angle-ironsmith and caulking. The “bōsin” of the rivetter gives the instructions as the head of the group. Wages are paid on a piece-rate contract basis. Because of the contracted wage, the cannery workers’ income is unstable; thus, they are paid a fixed wage (daily wage amount) during the off-season.
Smithing is the second most common industry, accounting for 28% of the total. Panel A of Table A.6 reveals that most of the smithing industry is relatively small scale. They make boulders, nuts, rivets, lathe stops, jigs, and spanners for can manufacturing and construction. They were subcontractors who manufactured metal parts to supply the larger factories in Tsukishima (Department of Health, Ministry of the Interior 1923a, p. 412). There are three types of jobs: yokoza, sakite, and tatara-fuki (bellows-blower). The yokoza is self-employed skilled worker (oyakata), while the other two are young unskilled laborers. In small operations, wages were based on a percentage based on voluntary business discussions with the proprietor, and working hours were not regular (pp. 411–413).
Appendix Appendix B Data Appendix
B.1 Tsukishima Survey: The First Social Survey in Japan
Kawai (1980) summarizes the history of social research methodology. The methodology of social research in Japan was formed in the 1930s. Social research methodologies had been developed in the 1910s and 1920s (Kawai 1980, pp. 47–48). During this period, “numerous social surveys were conducted by various survey entities, including government agencies, research groups, local administrative bodies, private organizations, survey research institutes, and researchers” (p.48). This is because “around WWI, there were many problems in labor, lifestyle, urban, health and sanitation, public safety, and civic due to the heavy and chemical industrialization, the monopolization of capitalism, and urban expansion.” The Report on Field Survey in Tsukishima, Kyobashi Ward, Tokyo (called Tsukishima Survey) [^79] Takano himself, who was responsible for overseeing the survey, refers to this field survey as the “Tsukishima Survey” (Department of Health, Ministry of the Interior 1923a, p. 6). Considering this, I call the Tsukishima Survey throughout this paper. used in this study is the most pioneering of these surveys of urban communities. Sekiya (1970, p. 43) regards the Tsukishima Survey as a pioneering household survey in Japan.
First, I summarize the investigators, research questions, and subjects of the Tsukishima Survey. The person responsible for the Tsukishima Survey was Iwasaburō Takano (hereafter “Takano”), a professor of social statistics and social policy at the Law School of Tokyo Imperial University. Takano became a professor of economics at Tokyo Imperial University in 1919, and became the director of the OISR in 1920. Takano placed great importance on gathering objective statistics about workers through social surveys (Sekiya 1970, p. 33). He had the experience of conducting the first household survey in Japan in 1916.808080This is called the “Survey of Household Income and Expenditure of 20 Artisans and Workers in Tokyo” and was inspired by Charles Booth (Miyoshi 1980, p. 33). Although the original documents of this survey were preserved at Tokyo Imperial University, those were destroyed in the Great Kantō Earthquake (Takano 1933, pp. 727–728). More precisely, before this household survey, Takano conducted a field survey in a poor area of Tokyo when he was a student at Tokyo Imperial University. The findings were reported in April 1894 as an exercise entitled “East London in Tokyo” (Kawai 1981b, p. 3). Takano has been a member of the seventh subcommittee (Rural Sanitation Subcommittee) of the Ministry of Home Affairs’ Health and Sanitation Investigation Committee since 1916.818181The Health and Sanitation Investigation Committee of the Ministry of Home Affairs was established on June 27, 1916, under the second Shigenobu Okuma Cabinet by issuing Imperial Ordinance No. 172. Initially, 34 members were appointed, and the first meeting was held on July 8. The reason for the establishment of this committee was based on the need for a policy response regarding the establishment and institutionalization of various investigation organizations and the penetration of the capitalist economy during the Meiji period. Another reason is the requirement for a public health response to the rising mortality rate, exemplified by the tuberculosis epidemic (Kawai 1981a, pp. 10–13). In 1917, the Seventh Section evolved into the Section on Urban and Rural Sanitary Conditions (Kawai 1981a, pp. 12–13). On October 22, 1918, Takano submitted a proposal to the subcommittee for a field survey of urban sanitary conditions, which was approved. The survey aimed to determine the housing conditions, household budget status, and health status (child health, production, stillbirths, deaths, and diseases) in the areas where skilled-worker families congregated (Sekiya 1970, p. 6). Initially, the survey area was planned to be Yanagibashi Yokogawa-chō in Honjo Ward. However, at a committee meeting on November 23, the survey site was changed to Tsukishima in Kyobashi Ward, which had more skilled worker households than in Yokogawa-chō (Kawai 1980, pp. 55–59). [^82] Takeda (2015, pp. 75–76) provides more details on how the survey site was changed. Takano was looking for a commissioned researcher to conduct the survey and asked Sakuzō Yoshino, a colleague at Tokyo Imperial University, to recommend a suitable person. Yoshino approached his former student Kotora Tanahashi, who had become a member of the fraternity club, a mutual aid organization for workers, to accept the investigator position. The diary written by Kotora Tanahashi (“Tanahashi Kotora Diary”) is in the possession of the OISR, and Takeda used the diary as the basis for her analysis of the circumstances surrounding the change of survey location. Tanahashi eventually recommended Yoshitsuru Yamana, who was his classmate in high school and Tokyo Imperial University, as a researcher and considered Tsukishima to be the best research area because of the ease of selecting survey households using the fraternity’s network (Takeda 2015, pp. 76–77). On October 26, Tanahashi and Yamana, together with the Kyobashi Branch of the fraternity club, conducted an on-site inspection of Tsukishima. Based on this inspection, Tanahashi and Yamana approached Takano about locating the survey station in Tsukishima. Finally, Takano and Yamana conducted an on-site inspection of Tsukishima on October 30 and decided on Tsukishima as the survey site.
Although the statistical method of sampling was not yet common at that time, Takano selected the survey sites based on his visionary thinking. In fact, “There can be no dispute that we have chosen Tokyo as the target of our metropolitan survey. However, it is impossible to survey the entire Tokyo area. We have no choice but to conduct a partial survey. For this purpose, it would be the best idea to select a representative area in Tokyo and survey that area narrowly but deeply.” (Department of Health, Ministry of the Interior 1923a, pp. 2–3). Takano further explained the need to focus on the working-class, which makes up most of the citizens, rather than to see the political and commercial center of Tokyo: “Most of Tokyo’s citizens are wage earners who subsist on labor. As is usual in large cities, a large number of people are engaged in manual labor in Tokyo. However, the most genuine type of worker is the skilled worker. Therefore, the area where many families of skilled workers congregate can be regarded as a representative area of Tokyo. And it seemed to me that a study of the social and sanitary conditions in that area could explain a large part of the social and sanitary conditions in Tokyo.” (Department of Health, Ministry of the Interior 1923a, pp. 3).
It is necessary to delve into the point where Takano states that “the most genuine type of worker is the skilled workers” and the fact that the main target of the Tsukishima Survey was skilled workers. [^83] In other words, the Tsukishima Survey was not designed for poor households. Therefore, unlike the surveys conducted in the UK by C. Booth and B.S. Rowntree, no standard of living was set in the survey (Sekiya 1970, p. 42). The description by Kawai (1981a, pp. 13–18) is helpful in this regard. As summarized in Section 2, Japan’s heavy and chemical industrialization throughout WWI led to a large concentration of city workers. As a result, the ratio of small-scale conventional factories based on “master employees” as is often observed in the textile industry decreased. On the other hand, the number of factories with a direct “factory owner-employee” employment relationship increased, particularly in the machinery and equipment, metal, and chemical industries. These workers in the modern industrial sector accumulated skills, formed households, and settled in the cities. It can be said that a new urban worker lifestyle was being formed. In short, although the city had a class of workers engaged in traditional industries, a new type of workers in the modern industrial sector, and a lower class, Takano believed that the skilled workers in the modern industrial sector would play a central role in terms of the labor force, labor movement, and urban life.
In addition, Sekiya notes the rise of the labor movement. As noted in Section 2, while wages rose throughout WWI, the price surge caused relative wages to decline. In 1916, Japan’s first labor protection legislation, the Factory Law, was enacted. However, the number of working days and hours did not decrease, and the number of labor disputes steadily increased (Sekiya 1970, pp. 24–26). Given this, Takano believed that workers should organize labor unions to help themselves with labor problems that could not be resolved through labor protection legislation. Sekiya (1970, pp. 26–40) speculates that Takano’s intention in conducting the Tsukishima Survey was to obtain basic data for solving labor problems, specifically to identify the scope of matters to be solved by workers on their own initiative (or covered by labor protection legislation). His insights suggest that this historical background should be addressed as a reason for Takano’s focus on skilled workers.
Kawai (1981b, pp. 15; 17) assessed the 40 households covered by the household survey included in the Tsukishima Survey as “the representativeness of the sample has remained ambiguous.” There is some mention but no clear description of how the sample households were characterized among the worker households in Tsukishima. It is reported that “to find suitable persons to fill in the form, we must ask for the assistance of school principals, police officers, physicians, and laborers, and we have requested a meeting of these persons at the survey office to explain the purpose of the survey, and to solicit applicants for the distribution of the household account book” (Department of Health, Ministry of the Interior 1923a, pp. 8–9). As explained, the target households were not randomly selected because the statistical sampling method was not prevalent. However, it is plausible that Takano designed this household budget survey in the sampling framework to capture the representative households in Tsukishima. As described in Section 3, the THBS households can be regareded as the standard skilled factory worker households in Tsukishima.
Table B.1: Survey Categories in the Field Study of Tsukishima Survey
Periods of the Chapter in the
field studies published report
1. Social mapping Nov. 1918–May 1919 A2
2. Physical examination of children Jan. 1919–July 1919 3
3. Physical examination of workers June 1919 3
4. Nutrition survey of the worker households May 1919 3
5. Survey of tenement houses Autumn, 1919 3
6. Survey of sanitation occupation June 1919 3
7. Sanitary survey in elementary schools Unknown 3
8. Factory labor survey Unknown 4
9. Labor household survey Nov. 1918–Jan.1920 2
10. Elementary school survey July 1919 2
11. Restaurant survey Dec. 1919 & April 1920 2
12. Field survey of vaudeville Dec. 1919 2
13. Survey of stalls and passersby Dec. 1919 2
14. Photographing Oct. 1920–Nov.1920 A2
Notes: The years/months listed in the first column (“Periods of the field studies”) are based on the description in the Department of Health, Ministry of the Interior (1923a, pp. 10–17). Figures listed in the second column (“Chapter in the published report” ) are based on the table of contents in the Department of Health, Ministry of the Interior (1923a, pp. 1–12; 1923b, pp.1–2). “A2” indicates the Appendix 2. Sources: Department of Health, Ministry of the Interior (1923a, 1923b, 1923c).
Next, I overview the survey items and methodology. The investigation was primarily divided into “documentary” and “field” surveys. The former was an analysis of the characteristics of Tsukishima, using published statistics such as the Police Department Statistics Book (Department of Health, Ministry of the Interior 1923a, p. 10). This was a pioneering attempt at the time and had influenced subsequent urban studies (Kawai 1980, pp. 61-62). The latter was a field survey including 14 survey items (Table B.1). It is not possible to describe all the survey items in detail, but following the report, they can be divided into four large categories: 1. survey on workers’ livelihood (chapter 2); 2. survey on sanitary conditions (chapter 3); 3. survey on labor conditions (chapter 4); and 4. social mapping and photography (appendix). The survey was conducted over two years, generally from November 1918 to November 1920, although the timing varied depending on the survey items (Table B.1).
A unique feature of the survey methodology is using young researchers as commissioned researchers. The household budget survey and leisure life survey were conducted by Yasunosuke Gonda, a faculty member of the Doitsugaku kyōkai gakkō (Association for German Sciences), who had researched activity photography and entertainment. Tetsuo Hoshino, a medical doctor, was responsible for the sanitary survey. The labor survey was handled by Yoshitsuru Yamana, a graduate of Tokyo Imperial University who contributed to the organization of labor unions. In addition, Toyotaro Miyoshi, a sociology student at the Tokyo Imperial University, and others became temporary employees. In other words, although the Tsukishima Survey was initially planned as an insurance and sanitation survey by the Ministry of Home Affairs, in reality, it was a long-term survey conducted as a joint study by young researchers in different fields, led by Takano (Kawai 1980, p. 63). At the time, the Health Bureau of the Ministry of the Interior was conducting several sanitary surveys of rural areas. Takano may have designed the urban sanitary survey from this background (Kawai 1981a, pp. 8–9). However, based on the idea that “health and sanitation surveys can at least encompass economic and social surveys,” Takano included a wide range of survey items, not limited to sanitation surveys (Department of Health, Ministry of the Interior 1923a, p. 56). Kawai (1980, p. 55) credits the Tsukishima Survey as a noteworthy pioneering concept: “going beyond a mere insurance and sanitation survey, it attempted to accurately depict the state of urban workers’ lives as they became established in the city during modernization, with skilled workers as its core.”
Figure B.1: Photographs of the Survey Office in Tsukishima
(a) Outside of the office
(b) Inside of the office
Note: Figures B.1a and B.1b show the outside and inside of the survey office in Tsukishima, respectively. Source: Department of Health, Ministry of the Interior 1923a, figures 1 and 2.
Figure B.1: Photographs of the Survey Office in Tsukishima
Takano stated, “To be as complete as possible, it is best to continue the survey over as long a period as possible and to obtain accurate facts by recording the workers’ household income and expenditures over a long period.” For this reason, he believed that “the best method of investigation would be to establish a survey station directly in the survey area and to have a full-time investigator reside there as much as possible, familiarize himself with the area and the people, and conduct the field survey” (Department of Health, Ministry of the Interior 1923a, p. 4). In fact, through the mediation of the Tsukishima police station, a survey office was rented at 9-3 Higashinakadori, Tsukishima, where Yoshitsuru Yamana moved in from November 1918 and contacted worker households (Department of Health, Ministry of the Interior 1923a, p. 7; Takeda 2015, p. 77). The survey office was open from November 1918 to December 1920, including the entire survey period. Figure B.1 shows the photographs of the survey office.
It is also important to note that Takano and his team held weekly meetings at the research center during the survey period. Toyotarō Miyoshi, who was an assistant in the survey, recalls the situation at that time as follows:
After the start of the survey office, all the members gathered once a week with Iwasaburo Takano present to discuss the implementation and progress of the survey and administrative matters. Iwasaburo Takano would always attend each meeting, happily listening to the results of the week’s research, inspecting the materials, providing detailed and thoughtful guidance, and usually ending the study session in a friendly atmosphere. Even now, I can recall the warm face of sensei (i.e., Professor Takano) of those days. It was always a friendly, laboratory-like atmosphere in which the C. Booth and S. Rowntree investigations were occasionally discussed.
Miyoshi (1980, p. 38)
He also states the following about Takano:
When the Tsukishima Survey was launched, the Faculty of Economics (in the Tokyo Imperial University) was becoming independent, and he was swamped both inside and outside the university. However, his enthusiasm for the Tsukishima Survey was extreme, and he was always present on his business trips to Tsukishima. Moreover, he made every effort to carry out his duties with precision.
Miyoshi (1980, p. 40)
Kawai (1981b, p. 28) assessed that there were no interconnections among the survey items mentioned above and that the survey “ended up being several fragmentary surveys rather than a comprehensive and systematic survey of workers’ lives.” This difficulty may be why previous studies have not frequently used the Tsukishima Survey. However, as already explained, Takano did not intend to conduct a large-scale social survey in Tsukishima in the first place. Instead, as he noted, the best method is to investigate a representative area “narrowly but deeply,” the goal was to improve the quality of the investigation rather than the scale of it. In this light, the fact that the researcher lived in the survey area and was constantly organizing and revising the survey results should be appreciated. Even though the Tsukishima Survey is the early social survey in Japan, it is of high quality as a prewar household survey.
Figure B.2: Official Published Report of the Tsukishima Survey
Notes: This photograph shows an example of the official report of the Tsukishima Survey published by the Department of Health, Ministry of the Interior, in 1923. Source: Department of Health, Ministry of the Interior 1923b, title page. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
Figure B.2: Official Published Report of the Tsukishima Survey
Finally, I summarize the format of the published report and the unpublished materials. The results of the Tsukishima Survey were officially summarized in one report published in December 1921 by the Sanitary Bureau of the Ministry of the Interior and in two appendices (Figure B.2 is an example of the front cover). The report was published under the name “Report on Field Survey in Tsukishima, Kyobashi, Tokyo,” and consists of four sections. The first section is a comprehensive review by Takano and outlines the survey. The second section summarizes the living conditions of the workers and was written by Gonda (partly by Yamana). Hoshino wrote the third section and summarized the sanitary conditions of workers. In the fourth section, Yamana summarizes the labor situation in Tsukishima. As for the two appendices, one contains statistical tables related to the above report, and the other organizes social maps and photographs. The correspondence between the survey items and the report is summarized in Table B.1. The Ministry of Home Affairs has already published these materials, which are available at the Gonda Yasunosuke Library of the OISR and the University of Tokyo Library.848484 In the OISR, these are included in the Archives of Gonda Yasunosuke (7-2; 7-3; 7-4). In addition, a second manuscript of the Tsukishima Survey report exists. This is the manuscript of the second series, which was preserved in the home of Gonda Hayao, the second son of Gonda Yasunosuke, and was discovered by Terade Koji. The second series includes a summary of the results of a household survey of elementary school teachers’ households in Tokyo, conducted as a supplementary survey to the Tsukishima Survey (see Terade (1982) for details). The manuscript of this second series was later donated to the OISR and is now available as part of the Gonda Yasunosuke Collection at the Institute (The OISR, Archives of Gonda Yasunosuke (8-12; 8-13; 8-14; 8-15)).
B.2 Original Micro-spreadsheets from the Household Budget Survey in Tsukishima Survey
To analyze the consumption smoothing behavior of THBS households, it is necessary to use the original micro-spreadsheets. The official report does not include household-level information, and a large part of the original forms used to prepare the report have little survived. Fortunately, however, some of the original micro-spreadsheets are kept in the archives of the OISR because Takano was its first director. This study uses the original micro-spreadsheets from the household budget survey (No. 9 in Table B.1), which was used in writing the second section of the official report by Gonda. While they have yet to be publicly available, I received permission from the OISR to use the original forms stored in the archives: the OISR, Archives of the Tsukishima Survey (THBS, unreleased). [^85] I would like to express my sincere thanks to Professor Kazue Enoki, Deputy Director of the OISR, for allowing me access to the documents. I wish to thank Ms. Mika Nakamura, an archivist at the institute, for her careful assistance during the research.
Figure B.3: An Example Page of the Budget Book in the Tsukishima Household Budget Survey
Notes: This photograph depicts an example page from the household expenses account of the THBS. The right page documents the details of costs. The left page includes the details of revenue and receiving and making payments on the goods (actual things, including gifts). Source: The THBS, account #6, 15th February 1919. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
Figure B.3: An Example Page of the Budget Book in the Tsukishima Household Budget Survey
The budget book used for the household budget survey is named the Kinsendeiri hikaechō (money receipts and disbursements account book). This budget book is designed based on the form used in Takano’s first household survey (Online Appendix B.1). [^86] Department of Health, Ministry of the Interior (1923a, p. 152). The budget book is to be filled out daily and includes four sections: income, expenses, receipts from credit purchases, and receipts and disbursements of actual goods. Most of the entries are in the income and expenses sections, indicating that living was basically conducted through wage receipt and consumption in cash. Figure B.3 shows an example of the budget book.
As described, the method used to select households for the survey was not recorded in detail (Section B.1). According to the report, they initially recruited just over 90 households through the factory, police station, and elementary school in Tsukishima. Of these, 50 households completed the book for at least one month, but 10 were deemed incomplete and not included in the report. In addition to the 40 budget books used in preparing the report, therefore, the OISR has a set of other budget books marked as “discarded”. According to my survey, about six incomplete books of the worker households have been determined to be in this “discarded” category. Therefore, based on the remaining conditions of the books, it is plausible to believe that roughly 50 households participated in the survey. Of the 40 households included in the report tally, two households had recorded for more than one year, 11 households had recorded for 6–12 months, and the remaining 27 households had registered for less than six months (Department of Health, Ministry of the Interior 1923a, pp. 152– 153). Initially, the survey period was targeted to be six months, but the average bookkeeping period for the 40 households was four months.
B.3 Trimming
Figure B.4: Structures of the THBS Datasets by Different Time Frequencies
(a) Daily panel dataset
(b) Semi-monthly panel dataset
(c) Adj. monthly panel dataset
Note: Figure B.4aB.4b, and B.4c show the observations in the daily, semi-monthly, and adjusted monthly panel datasets, respectively. For example, a marker in Figure B.4a indicates an observation in a specific year-month-date cell. Strictly, semi-monthly, and monthly datasets include both panel and cross-sectional units because aggregation shall trim fractions in the edges of the survey period. Figure B.4b shows 35-panel units and 5 cross-sectional units (#29; #31-33; #35). Figure B.4c also shows 35-panel units and 5 cross-sectional units (#29; #31-33;#35). Source: Created by the author using the THBS sample.
Figure B.4: Structures of the THBS Datasets by Different Time Frequencies
To check whether the statistics from THBS budget books contained any missing information, I digitized all 40 household budget books. Throughout this digitization process, I found that there are a few misreporting issues in the budget books.
First, there is a household that does not include income information (household #37). Although the official report by Gonda reported the average monthly income for this household (Department of Health, Ministry of the Interior 1923a, p. 183), no documentation would allow me to know how this average income was calculated. Hence, I excluded this household’s budget book. Similarly, a household does not include information on the payments to the credit purchases (household #39). Although this household’s budget book documented the amount purchased in credit purchases, I excluded this household’s budget book because the expenditure series becomes systematically smaller due to the omission of the payment information.
Second, the final pages of the budget books are sometimes omitted. This may be because the vendor who bound the budget books accidentally truncated the last page. All these incomplete sheets can not be included in the analysis.
Third, several households did not complete their budget books in the first and/or last month of the survey period. For example, the budget book of household #1 has statistics from 1st January 1919 to 12th January 1920. Similarly, the book of household #29 contains statistics from 26th January to 23rd February 1919. In semi-monthly and monthly datasets, these censored budget books may not be included in the aggregation. In the latter case, for instance, the budget sheets of 26th January and after 11th February 1919 are not included in a semi-month cell because the semi-month cell is defined from the 27th to 11th of the next month in this case. Thus, household #29 only has a cross-sectional observation and can not be included in my analytical sample.
Figure B.4 summarizes the structure of the 40 THBS households. Figure B.4aB.4b, and B.4c shows the observations in the daily, semi-monthly, and adjusted monthly datasets, respectively. As explained, there are five cross-sectional units (#29; #31-33; #35) in the semi-monthly and adjusted monthly datasets. Therefore, there are 35 units with a panel structure. After excluding a unit without complete income information (#37), I get the 34 THBS households as the analytical sample for my panel data analysis.
Regarding the time dimension, I trim all the observations before 12th January 1919 and after 11th November 1919 because the number of surveyed households is substantially small. For example, Figure B.4b shows that the number of cross-sectional units is three or less in both periods, causing substantial fluctuations in the net income. Consequently, my analytical sample includes 34 households measured between 12th January and 11th November 1919. The average number of semi-months per unit is $8.8$ (Std. Dev. $=5.0$). The characteristics of the 34 THBS households are summarized in detail in Section 3.
B.4 Testing the Potential Influence of the Unbalancedness
Table B.2: Results for the Balancing Tests
DV: Indicator variable for the households
with shorter time-series observations
Threshold:
(1) Median (2) 75 percentile (3) 25percentile
($150$ days) ($184$ days) ($79$ days)
Size -0.243 0.238 -0.135
[0.244] [0.254] [0.294]
Children aged 6–12 (%) 0.017 -0.011 0.026
[0.020] [0.019] [0.022]
Children aged 13–16 (%) 0.029 -0.050 0.057
[0.038] [0.045] [0.041]
Men aged 17+ (%) -0.008 0.008 -0.010
[0.019] [0.018] [0.026]
Intercept 1.149 -0.576 -0.164
[2.107] [2.048] [2.805]
Zero slope ($p$-value) 0.710 0.753 0.160
Maximized Log-likelihood -21.8 -18.5 -15.2
Pseudo $R$-squared 0.046 0.046 0.171
Number of households 33 33 33
The results from Probit models are reported. Robust standard errors are in brackets. Notes: The dependent variable is an indicator variable that takes one if the household has shorter time-series observations. Column 1 uses the median of the number of days per unit as the threshold. The result from the regression using the average number of days per unit ($144$ days) as the threshold is identical to that listed in column 1. Column 2 uses the 75 percentile of the number of days per unit as the threshold. Column 3 uses the 25 percentile of the number of days per unit as the threshold. All family size variables are the time-series average in each unit over the sample period. The proportion of children aged 0–5 years (%) is used as the reference group. Wald $\chi^{2}$ statistics $p$-values for the null of the zero-slope hypothesis are reported in the sixth row. The results from the expanded regressions including the head’s age and earnings are materially similar. The null of the zero-slope hypothesis is not rejected in all the specifications at the conventional level (not reported).
In this subsection, I test whether attrition leads to selection bias in the THBS sample. If the difference in the preference for consumptions predicted the large attritions, there must be statistically significant correlations between the family size variables and attritions. In other words, the units with shorter (longer) time-series observations have different preferences from those with longer (shorter) time-series observations. Table B.2 summarizes the results of the balancing tests (Section 3.2). The dependent variable is an indicator variable that takes one if the household has time-series observations less than the median (column 1), 75 percentile (column 2), and 25 percentile (column 3). In Column 1, all the estimated coefficients on the family size variables are close to zero and statistically insignificant, and the Wald statistics $p$-value suggests the null results. This is unchanged if I use the third and first quantiles in columns (2) and (3), respectively. Therefore, the family characteristics are similar between the households with short-term observations and those with long-term observations. This result supports the evidence that the unbalancedness in the THBS dataset is unlikely to lead to selection bias.
B.5 Comparing Distributions of the Household Size
Figure B.5 illustrates the household size distributions measured in the census and THBS sample. Pearson chi-squared test does not reject the null hypothesis of the equality of distributions with $p$-value of $0.948$. This supports the evidence that the THBS sample has a similar household size distribution to the population of Tsukishima (Section 3.1). As described in Section 3.2, while the THBS sample does not cover households with one person and with a very large number of people, it covers most of the classes in the household sizes with reasonable approximation to the population distribution. [^87] The number of households with one people is $362$, accounting for only $6$% of the total number of households. Similarly, the number of households with nine or more people is $294$, accounting for only $5$% (Section 3.2).
Figure B.5: Distribution of the Household Size:
Population Census v. the THBS sample
Notes: This figure shows the share of the household in each household size category to the total number of households in Tsukishima measured in the 1920 Population Census and the THBS samples. Pearson chi-squared test does not reject the null hypothesis of the equality of distributions with $p$-value of $0.948$. The share in each household size measured in the census is used as the theoretical probability in calculating $\chi^{2}$ test statistic (Pearson 1900). Sources: Created by the author using the Tokyo City Office (1922a, pp. 262–283) and the THBS dataset.
Figure B.5: Distribution of the Household Size:
Population Census v. the THBS sample
B.6 Estimating Average Head’s Monthly Income
Comprehensive estimation effort on the time series figures of average wages in prewar Japan was done in the Long-Term Economic Statistics (hereafter LTES). Specifically, the LTES estimates the average national salary series of “general occupations,” “manufacturing sector,” and “manufacturing occupations (by medium classification).” For the manufacturing occupations category, the LTES reports the average wage per day of factory workers in Tokyo City and Osaka cities from 1917 to 1922 (Ohkawa et al. 1967, p. 255). These values were estimated using the “Annual Statistics of the City of Tokyo” (hereafter, the ASCT) and “Osaka City Statistical Table.” However, the household survey includes the household head’s monthly income but no information on the daily wage. Hence, the “average daily wage per factory worker” estimated in the LTES is not useful for comparing the heads’ average monthly income from the THBS. In this section, I use available wage and labor statistics to estimate the average monthly income of male factory worker households in Tokyo and compare it to the THBS heads’ average monthly earnings.
To clarify the problem, I first estimate average monthly earnings most simply. Specifically, I use the ASCT (volume 17), which provides wage statistics as of the end of 1918, as a case study. For each industrial sector, the ASCT (volume 17) provides the “average daily wage of male factory workers aged 15 and over,” “the number of male factory workers aged 15 and over,” and “average number of working days per year” for four factory categories: factories with ten or fewer workers, and factories using or not using motive power. Using the “number of workers over 15 years old” in each factory category cell ($c$) as weights, I can get the average wage per month for a worker, for example in the machinery sector, as:
[…]
This value is estimated to be $27.8$ yen. This is the most straightforward estimate using daily wages, the number of workers, and the number of working days. Thus, it has a few critical issues that cause a downward bias in the estimate.
Figure B.6: Distribution of the THBS Heads’ Age
Notes: This figure shows the density of the THBS heads’ age calculated from the THBS daily panel dataset. The dashed line (red) indicates the average age of all the heads (33.2 years old). The dashed line (blue) shows the average age of the heads, excluding an elder head aged 67.6 (32.1 years old). The difference between the two average figures is statistically insignificant with $p=0.4087$. Source: Created by the author using the THBS dataset.
Figure B.6: Distribution of the THBS Heads’ Age
The first issue is that the wages of factory workers are aggregated in a “15 years old or older” category in the ASCT (volume 17). Figure B.6 shows the THBS heads’ age distribution. The sample mean is $33.2$ years old, which is unchanged if I exclude a head aged more than $60$ (i.e., $32.1$ years old). This means that I have to refer to the average daily wage for the 30s to 40s for comparison. However, since many workers are in their 20s in the distribution of entire workers, the average wage reported in the ASCT shall be considerably lower than the average wage for factory workers around 30 years old. Thus, I need the wage profiles by age to correct for this downward bias.888888 The ASCT (volume 18), which contains statistics for 1919, does not list such wages by age group but only average wages. As is clear, the same problem arises in this case.
Second, there is no consideration of ancillary pay such as bonuses and various allowances. As explained later, these wages shared a non-negligible proportion of the head’s earnings. Thus, estimation without considering ancillary wages would have a strong downward bias. In the following, I summarize the procedure for correcting these downward biases.
B.6.1 Daily Wage Distribution by Ages
Since the early wage surveys focused on average wages, official statistics in the 1910s do not include average wage statistics by age (bin). By the “Statistical Data Field Survey Law” (Law No. 52) in April 1922, average wages by age group were surveyed for the first time in the “Labor Statistics Field Survey” in October 1924 conducted by the Social Affairs Bureau of the Ministry of Home Affairs (Committee for the History of the Labor Movement 1964, p. 14). The result of this survey in Tokyo was published as the “Labor Statistics Table for Tokyo City and Suburbs” (hereafter, LST) by the Tokyo City Statistics Division in 1926.
Figure B.7: Average Daily Wage by Age Bins
(a) 1918
(b) 1919
Notes: This figure shows the average daily wage distribution by age bins based on the manufacturing census in 1924. Figures B.7a and B.7b show the distributions under 1918 and 1919 prices, respectively. The weighted average daily wage for a 30-39-year-old bin is shown in red. The weighted averages for all and 14+ bins are shown in gray and blue lines, respectively. The fitted fractional polynomial models are shown in the dashed lines. For the 1918 price data, the fitted model is $\hat{y}^{\textit{Daily Wage}}=-10.01+4.15a^{1/2}-0.33a^{2}$ with $R^{2}=0.99$. For the 1919 price data, the fitted model is $\hat{y}^{\textit{Daily Wage}}=-13.18+5.47a^{1/2}-0.43a^{2}$ with $R^{2}=0.99$. Sources: Created by the author using the Tokyo City Statistics Division (1926a, pp. 16; 244-245). Data on the consumer price index for the cities are obtained from Ohkawa et al. (1967, p. 135).
Figure B.7: Average Daily Wage by Age Bins
Based on the LST, Figure B.7 shows the average daily wages of male factory workers in Tokyo City’s machinery and equipment manufacturing industry by age group. In the LST, there are 13 age bins, and the number of workers and average wages are listed for each bin. Figure B.7a shows the distribution of wages adjusted to 1918 prices using the consumer price index (CPI) for urban areas (Ohkawa et al. 1967, p. 135). The weighted average for ages $14$ and older is $2.58$ yen, which is about the same as the weighted average for daily wages of $2.56$ yen. The $2.58$ yen is between the average values for the 20-24 age bin and the 25-29 age bin. Thus, it likely to represent the daily wage of workers around 25. This is consistent with the fact that workers are concentrated in the 20s. However, the weighted average for the 30-39 age bin is $3.02$ yen, which deviates significantly from $2.58$ yen. The distribution of wages at 1919 prices shows a similar result (Figure B.7b). Using weighted averages for earnings comparisons leads to the erroneous conclusion that the sample households are upwardly biased.
To address this issue, I use the following steps to estimate the distribution of wages of the factory workers in Tokyo City in 1918-1919, the observation period of the THBS sample. Since the same estimation method is applied to both years, I use 1918 as an example below.
First, the distribution of daily wages by age group was projected from statistics obtained from the LST. Following Royston and Altman (1994), I consider a fractional polynomial of degree $m$ as follows:
$\phi_{m}(a;\bm{\varphi},\mathbf{p})=\varphi_{0}+\sum_{j=1}^{m}\varphi_{j}a^{(p_{j})},$ (6)
where $a$ ($>0$) denotes class value of age bin, $\bm{\varphi}$ is a vector of coefficients and $\mathbf{p}$ is a vector of powers, satisfying $p_{1}<p_{2}<...<p_{m}$. $a^{(p_{j})}$ indicates the Box-Tidwell transformation, taking $\text{ln}(a)$ if $p_{j}=0$ and $a^{p_{j}}$ if $p_{j}\neq 0$. As suggested in Royston and Altman (1994, p. 433), the models requiring a degree higher than two ($m>2$) are rare in practice. The wage by age bin distribution illustrated in Figure B.7 clearly shows a simple quadratic curve. Thus, I consider a fractional polynomial function with degree two as follows:
$\phi_{2}(a;\varphi_{0},\varphi_{1},\varphi_{2},p_{1},p_{2})=\varphi_{0}+\varphi_{1}a^{(p_{1})}+\varphi_{2}a^{(p_{2})}.$ (7)
To model the function, I consider a set of powers $\mathscr{P}=\{-2,-1,-0.5,0,0.5,1,2,3\}$ in the spirit of Royston and Altman (1994, p. 434).
The maximum likelihood estimation is used to search for the most suitable candidate of the coefficients for a given $m$. Since $m=2$, there are ${}_{8}\mathrm{H}_{2}$ ($=36$) repeated combinations for the candidates of the parameters. By adding the $8$ simple candidates under the nested models (i.e., the models with degree-1 fractional polynomial) to be conservative, I run $44$ regressions in total to find out the model with the highest maximized log-likelihood. The best power vector suggested is $(p_{1},p_{2})=(0.5,1)$. [^89] The deviance, defined as the twice negative (maximized) log-likelihood, is calculated to be $-31.5$. Royston and Altman (1994) proposed the deviance difference, $D(m,\mathbf{p})-D(m,\mathbf{\tilde{p}})$, where $\mathbf{\tilde{p}}$ is the best power vector to compare the different models. This statistic has an asymptotic chi-squared distribution with $m$ degree of freedom. The estimated deviance difference with the simplest linear model (i.e., $m=1$ and $p_{1}=1$) is $54.6$ with $p$-value = $0.000$. Similarly, the difference with an another non-linear model with $m=1$ and $p_{1}=-2$ is $34.2$ with $p$-value = $0.000$. The projection model fitted using the best power vector is as follows:
$\hat{y}^{\textit{Daily Wage}}=-10.01+4.15a^{\frac{1}{2}}-0.33a^{1},\quad R^{2}=0.99.$ (8)
The dashed curve in Figure B.7 is the predicted distribution in equation 8. The fitting suggests that the predicted distribution works properly, especially below the 50-54 bin.
Next, I shift the predicted distribution based on the difference in the average daily wages in the ASCT and LST. The weighted average of the daily wage is $2.56$ yen for the LST, whereas that for the ASTC is $1.05$ yen, meaning that there is roughly $1.5$ yen difference. [^90] For the LST, I use the number of factory workers in each age bin as weight. The ASCT reports the average daily wage for four different categories of factories in each manufacturing sector: the factory scale (10+ or below ten workers) and engines (use or not). I calculated the weighted average of the daily wage using the number of factory workers in each age bin (12–14 and 15+) and in each category as weights. Thus, I shift the functional polynomial function by this amount of difference to yield the proxy distribution for 1918.919191This operation, therefore, assumes that the wage-by-age distributions in 1918 and 1924 are reasonably similar. In the LST (i.e., the statistic on which the projected distribution is based), the weighted average for those aged 14 and older was 3.25 yen. In comparison, the overall weighted average was 3.23 yen, a difference of only 0.02 yen. A similar trend can be observed in the statistics obtained from the ASCT: the weighted average for those aged 15 and older is 1.14, while the overall average is 1.13, a difference of only 0.01 yen. At least, in the light of the first moment, this confirms that the wage distribution in the ASCT is similar to that in the LST. Figure B.8a illustrates this operation: dotted curve is a fractional polynomial shifted from the original fractional polynomial shown as a short dashed curve.
The average daily wage for 12–14-year-old bin is reported to be $0.42$ yen in the ASCT, whereas that for 12–13-year-old bin in the LST is $0.60$ yen. This difference (i.e., $0.18$ yen) is smaller than the difference in the weighted average for all age bins (i.e., $1.5$ yen). This implies that the daily wage by age distribution in the ASCT could be smoother (i.e., flatter) than that of the LST. Thus, the shifted fractional polynomial would provide stronger shrinkage for those bins of younger age. In other words, the shifted curve shall get closer to the actual (unobservable) wage distribution as age increases.
The long dashed curve in Figure B.8a examines this relationship by fitting another fractional polynomial model using the 12-14-year old bin figure of the ASCT and the predicted wages for 25–65 from the shifted fractional polynomial.929292 The fitted model is as follows: $\hat{y}^{\textit{Daily Wage}}=-1.19+0.14a^{1}-0.002a^{2},\quad R^{2}=0.98.$ The deviance is reported to be $-149$. I used the 25–65 age range for the prediction based on the age bins used in the LST. Let this curve predict wage distribution for the true distribution. The average daily wage for the 30–40-year-old bin in the shifted fractional polynomial is $1.55$ yen, which is similar to that for the same-age bin in the predicted wage distribution (i.e., $1.52$ yen). However, the wage gaps for the younger ages show more significant differences. As long as our interest is in the late 20s to the early 40s, therefore, it is plausible to employ the predicted wages for the wage comparison in Section 3.
Figure B.8b illustrates the result of the same analysis from the wage data of 1919 price, reflecting the increasing trend in price between 1918 and 1919. The average daily wage for a 30–40-year-old bin in the shifted fractional polynomial is $2.46$ yen. Again, this is similar to that for the predicted wage distribution (i.e., $2.42$ yen).
Figure B.8: Daily Wage by Ages Distribution:
Shifted Fractional Polynomial
(a) 1918
(b) 1919
Notes: The short dashed curve illustrates the fractional polynomial (FP) in Figure B.7. The dotted curve indicates the FP shifted by the difference in the average daily wage between the ASCT and LST. The ASCT documents the wage statistics for four different categories for each manufacturing sector: the scale of the factory (10+ or below ten workers) and engines (use or not). I calculated the weighted average of the daily wage using the number of factory workers in each age bin and the category as weights. The long-dashed curve shows the FP refitted using the average daily wage for 12-14-year-old bin ($0.42$ yen) measured in the ASCT and the daily wages for 25-65-year-old suggested in the shifted FP. The average daily wage for a 12–14-year-old bin is only available for the ASCT of 1918. Therefore, this figure for the 1919 price is calculated using the ratio between the average daily wage ($1.05$ yen) and that for 12–14 years old bin ($0.42$) in 1918 as $1.80\times(0.42/1.05)$, where $1.80$ is the average daily wage measured in the ASCT of 1919. The red dashed line is the average daily wage for 30–40-year-old in the shifted FP. The blue dashed line is the average daily wage for 30–40-year-old in the refitted FP. Sources: Created by the author using the Tokyo City Office (1921, pp. 726–757); Tokyo City Statistics Division (1926a, pp. 16; 244-245).
Figure B.8: Daily Wage by Ages Distribution:
Shifted Fractional Polynomial
B.6.2 Ancillary Wage
Figure B.9: Log-Relationship between the Number of Workers and Share of Ancillary Wage: Original Data from the Factory Survey by Kitazawa (1924)
Notes: This scatter plot shows the relationship between the log-transformed number of factory workers and the log-transformed share of ancillary wage. The raw data are listed in Table B.3. The dashed line indicates the linear fit using the log-transformed variables: the fitted regression is $\hat{y}=0.747+0.428x$ with $R^{2}=0.82$. The predictions are materially similar if I use the fractional polynomial using the raw data on the number of factory workers and share of ancillary wage in Table B.3. In the case of fitting based on the raw data, the fitted fractional polynomial model is $\hat{y}^{\textit{Daily Wage}}=53.6-399.3a^{-1/2}+(2.84\times 10^{-10})a^{3}$ with $R^{2}=0.87$. I prefer to use linear regression herein because it can provide more precise fitted values for factories with a small number of workers. Source: Kitazawa (1924, pp. 20; 58–62).
Figure B.9: Log-Relationship between the Number of Workers and Share of Ancillary Wage: Original Data from the Factory Survey by Kitazawa (1924)
The daily wage estimated in Section B.6.1 is a fixed salary and thus, does not include additional salaries such as bonuses and allowances (hereafter, wages other than regular salaries are collectively referred to as ancillary wages). However, there are no systematic statistics on these ancillary wages throughout the prewar period (Committee for the History of the Labor Movement 1964, pp. 14–15). Moreover, since the Tsukishima Survey targets skilled worker households, information on the ancillary wages paid to skilled workers is necessary. Fortunately, I found a survey report named “Wage Survey Report” edited by Kitazawa (1924). This report includes the result of a survey on the wages for machine finishers (fitters) and turners at six machine factories in Tokyo from May–October 1921. Thus, this is particularly useful material to know the wage structure of skilled factory workers.
According to Kitazawa (1924), there are two methods of paying wages to workers in the machine industry: one is to use hourly wages to calculate fixed salaries, and the other is to combine hourly wages with piece-rate wages, with the former being the most common (p. 33). Despite this, wages paid in both methods are usually divided into fixed salaries and other ancillary wages, and there is not much difference in essence (p. 39). Fixed wages are salaries added from a fixed hourly wage rate, while ancillary wages include bonuses and overtime pay, night shift pay, and labor premiums (profit sharing). In the case of a combination of hourly and piece-rate pay systems, the accessory wage includes a share of the piece-rate profit.
Panel A in Table B.3 summarizes the survey results. The surveyed factories included 86, 157, 376, 623, 928, and 2,837 workers, ranging from small to large plants. The percentage of skilled workers (finishers and turners) in the surveyed factories ranged from $27$% to $77$%, and the portion of monthly ancillary wages was 12–53%. Figure B.9 is a linear regression fit using log-transformed data on the percentage of ancillary wages and the number of workers. There is a positive correlation between the factory size and ancillary wages. This is consistent with the historical fact that larger factories have more extensive benefits than smaller ones (Odaka 1999).
Next, I estimate the percentage of ancillary wages at the Tsukishima factories. One of the most frequently used statistics on factories in Tokyo is the Handbook of Factories (kōjyō tsūran), which covers all the factories with ten or more workers as of January 1, 1920. This handbook is the most comprehensive source for ascertaining the number of factories and their sizes around 1919. This document does not cover the number of tiny working places with less than nine workers. Importantly, however, the purpose of Panel C in Table 2 is to compare the average monthly wage of the skilled factory workers in Tsukishima and the THBS heads’ average monthly earnings. Since the THBS heads did not work in the tiny workplaces, the censoring at the left tail of the factory size distribution is unlikely to influence this comparison.
Panel B of Table B.3 summarizes the distributions of the number of factories, the total number of workers, and the estimated average share of ancillary wage in Tokyo City (Column 1) and Tsukishima (Column 2). The average ancillary wage at the Tsukishima factories is 12%, which is slightly larger than the value for the entire city of Tokyo (9%). The percentage shares of the number of factory workers by each bin are listed in Columns (1) and (2) of Panel C in Table B.3. As shown, Tsukishima factories tended to have larger factories than those of the entire city, which is reflected in the slightly larger estimate in the average ancillary wage for the Tsukishima factories. In Column (3) of Table B.3, the number of THBS heads whose occupation, factory name, and factory size are available are listed. While the percentage shares of the heads in each bin show a similar distribution to that of the Tsukishima factories, the share of middle (large) sized factories is larger (smaller) than the population. If interpreted literally, this tendency implies that the (unobservable) true average ancillary wage for the THBS heads may be slightly lower than the population mean. [^93] Note again that the purpose of Panel C in Table 2 is to compare the average monthly wage of the skilled factory workers employed in the Tsukishima factories to the THBS heads’ average monthly earnings. Therefore, the purpose herein is to estimate the average ancillary wage for the target population, say the Tsukishima factories but not to calculate the average ancillary wage for the THBS sample itself.
Table B.3: Estimating the Average Share of Monthly Ancillary Wage
Panel A: Original Data on the Share of Ancillary Wage
Document: Report of Factory Survey by Kitazawa (1924)
Survey subject: Six Machinery Factories in Tokyo City
Survey month and year: May to October 1921
Number of Share of Share of Share of
factory skilled fixed ancillary
Factory workers workers (%) wage (%) wage (%)
A 157 44 83 17
B 86 53 88 12
C 376 27 57 43
D 623 77 68 32
E 2837 NA 47 53
F 928 41 60 40
Panel B: Estimating the Average Share of Ancillary Wage
Document: Handbook of Factories (1921 edition)
Survey subject: Machinery Factories with 10+ workers in Tokyo City
Survey month and year: January 1920
Machinery Factories
(1) Tokyo City (2) Tsukishima
Number Number Average share Number Number Average share
of factories of factory of ancillary of factories of factory of ancillary
Factory size workers wage (%) workers wage (%)
10–99 workers 760 18,760 8 39 1,252 9
100–199 workers 52 7,155 17 5 713 18
200+ workers 34 21,764 31 3 3,851 38
Total 846 47,679 9 47 5,816 12
Panel C: Comparisons of the Distributions of Factory Size
Machinery Factories
(1) Tokyo City (2) Tsukishima (3) THBS
Number % share Number % share Number % share
of factory of factory of THBS
Factory size workers workers heads
10–99 workers 18,760 39 1,252 22 6 21
100–199 workers 7,155 15 713 12 7 25
200+ workers 21,764 46 3,851 66 15 54
Total 47,679 100 5,816 100 28 100
Notes: Panel A: This panel shows the size of factories and the share of fixed and ancillary monthly wages in six factories surveyed in Kitazawa (1924). The number of factory workers surveyed in factories A–F are 14, 23, 9, 11, 11, and 16, respectively. The share of these surveyed skilled workers to all skilled workers is 20.3 in Factory A, 50.0 in Factory B, 5.3 in Factory C, 2.3 in Factory D, and 4.2 in Factory F, respectively. This figure for factory E is unavailable because the number of skilled workers is unknown (Kitazawa 1924, p. 21). Panel B: Subcolumns 1–2 and 4–5 summarize the number of factories and total number of workers in Tokyo City and Tsukishima, respectively (Handbook of Factories 1921). Subcolumns 3 and 6 show the average share of ancillary monthly wage predicted in Figure B.9 for Tokyo City and Tsukishima, respectively. Panel C: Columns 1–3 summarize the number of factory workers and its share in percentage points in Tokyo City, Tsukishima, THBS sample, respectively. The statistics in Columns 1 and 2 (same figures used in Panel B) are from the Handbook of Factories (1921). In Column 3, the THBS heads who worked in the machinery factories are considered. There are 28 heads whose factory names and size are available. For a blacksmith included in these heads, I assume the number of workers to be 10, based on the report (Department of Health, Ministry of the Interior 1923a, p. 412). Sources: Panel A: Kitazawa (1924, pp. 20; 58–62). Panel B: The number of factory workers is taken from the Ministry of Agriculture and Commerce (1921, pp. 613-629; 674). The average shares of ancillary wage are calculated using Kitazawa (1924, pp. 20; 58–62). See also Figure B.9. Panel C: Data on the number of factory workers and factories are from the Ministry of Agriculture and Commerce (1921, pp. 613-629; 674). Data on the number of heads are from the THBS sample.
B.6.3 Estimating Monthly Income of Adult Male Factory Workers
Table B.4: Estimating the Monthly Income of Adult Male Factory Workers in Machinery Sector
Weight for the Ancillary Wage
(1) Tsukishima (2) Tokyo City
$\vartheta=0.12$ $\vartheta=0.09$
Panel A: Components in equation 14: $w^{Daily}_{j}$, $l_{j}$, and $n_{m}$
Average daily wage ($w^{Daily}_{j}$)
$w^{Daily}_{1918}$ 1.52 1.52
$w^{Daily}_{1919}$ 2.42 2.42
Average annual working days ($l_{j}$)
$l_{1918}$ 313 313
$l_{1919}$ 322 322
Frequency weights ($n_{m}$)
$n_{1}$ 11 11
$n_{2}$ 24 24
$n_{3}$ 29 29
$n_{4}$ 26 26
$n_{5}$ 22 22
$n_{6}$ 18 18
$n_{7}$ 11 11
$n_{8}$ 10 10
$n_{9}$ 6 6
$n_{10}$ 5 5
[…]
where $m\in\{0,1,2,...,12\}$ corresponds to the year-month cells from December 1918 to December 1919. This way of weighting assumes that the average daily wage of machinery workers had increased during this period. This assumption is plausible given the historical fact that the average wage in the manufacturing industry indeed showed a secular increasing trend around 1920 in Tokyo City (Ohkawa et al. 1967, p. 255).
Secondly, I use the number of observations in each year-month cell as the frequency weight. Let $n_{m}$ be the number of households measured in year-month cell $m$. Then, the average monthly wage can be defined as:
[…]
where $m\in\{\text{December}~{}1918,\text{January}~{}1919,...,\text{December}~{}1919\}$ is the year-month cell and $n_{m}$ is the frequency weight for year-month cell $m$.
Substitute equation 1112, and 13 into equation 10 to yield the final specification:
[…]
Panel A in Table B.4 summarizes all the components in equation 14. In column 1 of Panel B in Table B.4, I calculate the average monthly wage under the share of the ancillary wage for the factories in Tsukishima ($\vartheta$). The estimate is $55$ yen, which is approximately $2$ yen higher than the average monthly earnings in the THBS heads ($53$ yen, see Panel C of Table 2) as I expected in Online Appendix B.6.2. Therefore, the deviation from the population mean ($55$ yen) is not too large to infer the mean household behaviors of the skilled workers in the machinery factories in Tsukishima. Given that the average daily wage of the male factory workers in the skilled age range is $2.5$ yen (Figure B.8), $2$ yen difference is less than one day’s earnings. This is consistent with the findings in my descriptive analyses that show the similarities in the distribution of family structure, average household size, and share of ancillary wage between the THBS households and Tsukishima households (Section 3; Online Appendix B.6.2).
Column 2 uses the share of the ancillary wage for the entire Tokyo City ($\vartheta=0.09$) to calculate the average monthly wage. The estimate is $53.6$ yen, which is slightly smaller than that in Column 1, reflecting the difference in the weights for the ancillary wages. When I see this figure as an unbiased estimate for the entire city, the difference in the average factory size between the entire Tokyo City and Tsukishima may be reflected in this deviation (Panel B in Table B.3).
B.7 Measurement Errors in the Unadjusted Panel Dataset
Figure B.10: Comparisons of Average Semi-monthly Income and Expenditure
from Different Aggregations
(a) Unadjusted Semi-monthly Panel Dataset: Income and expenditure
(b) Unadjusted Semi-monthly Panel Dataset: Net income
(c) Shifted Semi-monthly Panel Dataset: Income and expenditure
(d) Shifted Semi-monthly Panel Dataset: Net income
(e) Adjusted Semi-monthly Panel Dataset: Income and expenditure
(f) Adjusted Semi-monthly Panel Dataset: Net income
Note: Figures B.10aB.10cB.10e show the average semi-monthly income and expenditure calculated from unadjusted, shifted, and adjusted panel datasets, respectively. Unadjusted series defines the days from the 1st to the 15th of each month as the first half, whereas the days from the 16th to the end of each month as the second half. The shifted series defines the days from the 2nd to the 16th of each month as the first half, whereas the days from the 17th to the 1st of the following month are the second half. Adjusted series is the baseline definition in the main text, which uses the days from 12th to 26th in each month as the “first-half” and those from 27th to 11th in the next month as the “second-half”. As explained in Section 3, the raw series contains considerable measurement errors due to the timing of payday. The shifted series can deal with the measurement errors from the timing of payday but can not adjust the difference in the payday timing and expenses. In other words, the shifted series ignores the households’ decision on the consumption schedule in each semi-month. The adjusted series can deal with measurement errors and take the timing of the household expenses into account. Figures B.10bB.10dB.10f illustrate the average semi-monthly income minus expenditure calculated from unadjusted, shifted, and adjusted panel datasets, respectively. Source: Created by the author using the THBS sample.
Figure B.10: Comparisons of Average Semi-monthly Income and Expenditure
from Different Aggregations
Figure B.10 explains how different ways of aggregation influence the measured time-series in income and expenditure. Figures B.10aB.10cB.10e show the average semi-monthly income and expenditure calculated from raw, shifted, and adjusted panel datasets, respectively.
The unadjusted (calender-month) series defines the days from the 1st to the 15th of each month as the first half and the days from the 16th to the end of each month as the second half. This series shall contain substantial measurement errors due to the mismatches in the income and expenditure data explained in Section 3.3. The time-series plots of average income in Figure B.10a have substantially large variances and fluctuate over time. This is because some semi-monthly income is assigned to the wrong semi-month cells. In these wrong cells, the measured incomes shall take unnaturally greater values than usual. Moreover, the time-series plots of average expenditure do not chase those of the income. For example, the average expenditure in the second half of June greatly differs from the average income in the same cell. Consequently, the average net income (i.e., income minus expenditure) in each semi-month is relatively flat and takes a few odd cells with large positive values (Figure B.10b).
The shifted series defines the days from the 2nd to the 16th of each month as the first half and the days from the 17th to the 1st of the following month as the second half. Although the shifted series can partly deal with the measurement errors in income, it does not adjust the difference in the payday timing and expenses. Figure B.10c indeed gets a much smaller variance in the average incomes over time because all the paydays are now fixed to assign the suitable semi-month cells. However, the expenditure series still does not chase the income series well because the payday is placed on the latter edge in each semi-month cell. In other words, the average expenditure in each semi-month reflects a large part of the consumption based on the income of the previous semi-month cell. Since the average incomes take a better series than Figure B.10a, the average net income offers better time-series plots (Figure B.10d). However, it still has a few cells with large negative and large values due to the deviations of expenditure from income.
The adjusted series uses the days from the 12th to 26th of each month as the “first half” and those from the 27th to 11th of the following month as the “second half”. By shifting the entire semi-month cells, this adjusted series can deal with the measurement errors from the payday shifts and fix the miss-matching in the timing of income and expenditure. In other words, each payday is assigned to a suitable semi-month cell, and each semi-month cell reflects the household’s decision on consumption given the latest income. Figure B.10e shows similar time-series plots of average income in Figure B.10c, in which has a corrected semi-monthly average income series only. Notably, the expenditure series are now chasing the income series in Figure B.10e because each semi-month cell’s expenditure is based on the income in the same cell. Accordingly, Figure B.10f now shows a smoother series in the net income. It shows a clear increasing trend in the net income and has a smaller variance over time.
These illustrations support the evidence that the adjusted semi-monthly datasets can offer cleaner within variations in income and expenditure because they can fix the measurement errors and can reflect the households’ consumption behaviors based on their latest incomes measured in the same semi-month cell.
B.8 Distribution of the Idiosyncratic Shocks
Figure B.11: Distribution of the Idiosyncratic Shocks
Notes: This figure shows the distribution of idiosyncratic shocks in the semi-monthly THBS panel dataset. The idiosyncratic shocks are estimated as the residuals from the regression of the first difference in income on the first difference in aggregate income. The red line indicates the mean of the idiosyncratic shocks ($-0.235$). Source: Created by the author using the THBS sample.
Figure B.11: Distribution of the Idiosyncratic Shocks
Figure B.11 illustrates the distribution of the idiosyncratic shocks predicted using the semi-monthly panel dataset. [^95] I used the semi-monthly panel dataset to overview the entire distribution of the idiosyncratic shocks rather than the daily panel datasets. This is because the denser panels include a set of consecutive household-date cells with no income. These cells can not be used to predict the idiosyncratic shocks as the residuals economically meaningfully, and thus it is practically impossible to illustrate the idiosyncratic shocks based on the daily dataset. The idiosyncratic shocks are calculated as the residuals from the regression in the first difference in semi-monthly income on the first difference in aggregate semi-monthly income. Let $n_{t}$ be the number of households in semi-month $t$ in the year 1919. For household $i$ in semi-month $t$ cell, the residual is then defined as follows:
[…]
B.9 Unit Root and Autocorrelations
Table B.5 presents the results of the Fisher-type panel unit root tests for the total expenditure in the semi-monthly (Column 1) and adjusted monthly (Column 2) panel datasets. Both columns show that the null hypothesis of unit roots in all the panels is rejected at the conventional level. The results of the unit root tests for consumption subcategories and variables used to test the risk-coping mechanisms are not reported in this table because those are materially similar results. In Section 4, I used the cluster-robust variance-covariance estimator in the regressions to allow arbitrary autocorrelations (Arellano 1987). Despite this, I confirmed that the null hypothesis of no serial correlation is not rejected at the conventional level in Wooldridge’s autocorrelation test for almost of the consumption and net income variables (Wooldridge 2002).
Table B.5: Results of the Unit Root Tests
Test statistics (1) Semi-Monthly (2) Adjusted Monthly
$P$-statistic $p$-value 0.0000 0.0001
$Z$-statistic $p$-value 0.0000 0.0000
$L^{*}$-statistic $p$-value 0.0000 0.0000
$P_{m}$-statistic $p$-value 0.0000 0.0000
The results of the Fisher-type panel unit root tests based on augmented Dickey-Fuller (ADF) tests are reported in this table. The null hypothesis is that all the panels contain unit roots, whereas the alternative hypothesis is that at least one panel is stationary. In all the specifications, the process under the null hypothesis is assumed to be a random walk with drift. The demeaned data are used to address the effect of cross-sectional dependence. The number of lagged differences in the ADF regression equation reported is set as one because including two or more two lags reduces a large part of the panel units available for computing the test statistics due to the unbalancedness of the panels. I confirmed that including two lags in the semi-monthly panels does not change the results. Including two lags in the adjusted monthly panels and three lags in the semi-monthly panels is impossible. See Choi (2001) for the details of the tests. These tests need both the first-differenced dependent variable and the lagged dependent variable. Therefore, the units with less than four time-series observations are excluded in the regressions. Source: Calculated by the author using the THBS sample.
B.10 Risk Preferences
The potential influences of heterogeneous risk preference on standard risk-sharing regressions are discussed in a few works (Schulhofer-Wohl 2011; Mozzoco and Saini 2012). In my empirical setting, the heterogeneity in the risk preference is assumed to be removed by using the household fixed-effect. When the risk preferences are difference over households, say $\sigma=\sigma_{i}$, equation 26 becomes:
$\log c_{i,t}=\log c_{t}^{a}+\frac{1}{\sigma_{i}}(\log\omega_{i}-\omega^{a})+(\theta_{i,t}-\theta_{t}^{a}).$ (16)
The first differencing removes the second term with the coefficient of relative risk aversion in the right-hand side of this equation, even though it varies over the households.
Schulhofer-Wohl (2011) shows that the heterogeneous risk preference may lead to bias in the estimation of the income elasticity under the standard risk-sharing regressions and proposes a parametric test for that heterogeneity. This test is valid under a particular utility function. Specifically, the bias could be raised under the case that the coefficient of relative risk aversion interacts with the aggregate shock in the first-order condition for consumption because the interaction means that the aggregate shock can not be removed using the time-fixed effect in the risk-sharing regressions. [^96] Schulhofer-Wohl (2011, pp. 928–929) considers the utility maximization problem under the following weighted sum of discounted expected utilities: $\sum_{i}\dot{\alpha}_{i}E\left[\sum_{t}\dot{\beta}^{t}\frac{[c_{i,t}(s_{t})]^{1-\dot{\gamma}_{i}}}{1-\dot{\gamma}_{i}}\right],$ (17) where $\dot{\alpha}_{i}$ is Parate weight and $\dot{\gamma}_{i}$ is the coefficient of relative risk aversion for household $i$. The log-transformed first-order condition becomes: $\text{log}c_{i,t}=\frac{\text{log}\dot{\alpha}_{i}}{\dot{\gamma}_{i}}-\frac{\text{log}\dot{\lambda}_{t}}{\dot{\gamma}_{i}},$ (18) where $\dot{\lambda}_{t}$ indicates the Lagrangian multiplier. This result implies that the effects of the aggregate shocks ($\dot{\lambda}_{t}$) depend on the household’s attitudes toward risk and that the first-differencing does not remove the second term in equation 18. As explained, this setting differs from the one used in this current study. Mozzoco and Saini (2012) consider a more general case for testing the homogeneity of risk preferences. Their test requires a panel dataset with a very long time dimension that can not be applicable to my THBS dataset. [^97] The data used in their study covers the household-level monthly panel dataset surveyed in rural India between 1975 and 1985. The households with fewer than 80 data points are not included in the homogeneity tests (Mozzoco and Saini 2012, p. 453). See also Shrinivas and Fafchamps (2018) and Mozzoco and Saini (2018) for the subsequent discussions.
An alternative way to test potential heterogeneity in the risk preference is to use direct measures on the risk preferences of workers. Schulhofer-Wohl (2011) uses the results of experimental risk-tolerance questions on the workers’ risk preferences in the Health and Retirement Study to test whether the correlations between earnings and aggregate shocks differ across high- and low-risk-tolerance workers. [^98] This question is asking if the respondents would be willing to take a risk to get a new job with a higher income, even if there’s a 50-50 chance of a reduction in income by a third (Schulhofer-Wohl 2011, P. 931). They found that the high-risk-tolerance workers have more procyclical incomes than the low-risk-tolerance workers.
Since the THBS has no experimental questions, it is impossible to apply the same test. Given this limitation that the direct test using experimental risk-tolerance question is unavailable, I compare the correlations between the heads’ income and aggregate consumption over sub-groups to get insights into the potential difference in the heads’ attitudes toward risks. For household $i$ and semi-month $t$, the specification is characterized as follows:
$\log\tilde{y}_{i,t}=\xi_{1}\log c_{t}^{a}+\xi_{2}d_{i}\log c_{t}^{a}+ox_{i,t}+\iota_{i}+\varepsilon_{i,t},$ (19)
where $\tilde{y}_{i,t}$ is head’s earnings, $c_{t}^{a}$ is aggregate consumption, $x_{i,t}$ is the household size, $\iota_{i}$ is the household fixed-effect, and $\varepsilon_{i,t}$ is a random error term. The aggregate consumption is calculated as the cross-sectional average of the total expenditure among the THBS sample, given that the population mean of consumption expenditures among all the skilled factory workers’ households in Tsukishima is unavailable.
$d_{i}$ is an indicator variable that takes one if household $i$ belongs to a specific group that might have had different preferences on their jobs than the other heads in the sample. Firstly, I consider the case that the heads who worked in the large-scale factory had different risk preferences. Although there are no specific anecdotes behind this, the workers in the large-scale factory might have had a lower risk tolerance than those in the small- and medium-scale enterprises. 13 heads who worked in the Ishikawajima shipyard and governmental factory are classified into this group. Secondly, I assume that the heads who worked in the smithing and the non-machinery factories had different risk preferences. As described in Online Appendix A.5, skilled workers in the smithing sector could be small business owners. [^99] Workers in the canning factories might have had a different preference on the risk preferences (Online Appendix A.6). However, my THBS sample does not include the workers in canning factories. The workers in non-machinery factories might also have different job preferences than the machinery factory workers. This group has only three heads: one worked in a smithing factory, and two worked in food and paper-making factories.
Table B.6 presents the results. In Column (1), the estimated coefficient on the interaction term is close to zero and statistically insignificant. Thus, the sensitivities to aggregate consumption are similar across different scales of the factories. Column (2) further shows that the estimated coefficient on the interaction term is not statistically significant under an alternative proxy of risk tolerance. It suggests that the heads worked in the smithing and non-machinery factories show similar responses to the aggregate shock. Since the aggregate consumption is the cross-sectional sample average, the aggregate variable may contain noise in the period with the small number of cross-sectional observations. To deal with this potential issue, I exclude October and November 1919 in Columns (3) and (4). As shown, the estimated coefficients on the interaction terms are largely unchanged in both columns.
This finding is consistent with the fact that these heads have similar characteristics in nature because the THBS focused on skilled factory workers in a representative manufacturing area, Tsukishima. Therefore, it is plausible to see that the THBS households had more or less similar attitudes toward the risks.
Table B.6: Results for Testing Heterogeneous Risk Preferences
DV: ln(head’s earnings)
Analytical Period
(1) Full (2) Full Jan. to Sept. Jan. to Sept.
Aggregate Consumption 0.277* 0.273 0.361** 0.362*
[0.158] [0.198] [0.164] [0.198]
Aggregate Consumption -0.041 -0.061
$\times$ Large Scale Factory [0.453] [0.415]
Aggregate Consumption -0.089 -0.178
$\times$ Smithing/Non-machinery [0.215] [0.215]
Observations 261 261 239 239
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level. Notes: This table shows the results of equation 19 for the semi-monthly panel dataset. ‘Aggregate Consumption’ indicates the cross-sectional sample average of the total expenditures. ‘Large Scale Factory’ indicates an indicator variable that takes one if the head worked in the Ishikawajima Shipyard or governmental factory. ‘Smithing/Non-machinery’ indicates an indicator variable that takes one if the head worked in the smithing, food, or paper-making factories. All the regressions include the family size variable and household fixed-effect. Columns 1–2 show the results for the entire sample period. Columns 2–3 show the results for the period between January and September 1919 to remove year-month cells with the smaller (cross-sectional) number of households (Online Appendix Figure B.4).
Appendix Appendix C Empirical Analysis Appendix
C.1 Conceptual Framework
Consider the economy where household $i=1,...,N$ receives an uncertain income at time $t$ depending on the state of the world $s_{t}\in S_{t}$ as $y_{i,t}(s_{t})$. For infinite time horizon, the state $s_{t}$ realizes with probability $\pi(s_{t})\in[0,1]$. The weighted sum of the expected lifetime utility of $N$ households in the economy is expressed as:
$\sum_{i=1}^{N}\omega_{i}\sum_{t=0}^{\infty}\rho^{t}\sum_{s_{t}\in S_{t}}\pi(s_{t})u(c_{i,t}(s_{t}),\theta_{i,t}(s_{t})),$ (20)
where $\omega_{i}$ is the social planner’s weight representing the reciprocal of the marginal utility of each household ($10<\omega_{i}<$ 1 and $\sum_{i=1}^{N}\omega_{i}=1$), $10<\rho^{t}<$ 1 is the discount factor, $c_{i,t}(s_{t})$ is consumption, and $\theta_{i,t}(s_{t})$ is a preference shock. For all states $s_{t}$, the aggregate resource constraint is given by:
$\sum_{i=1}^{N}c_{i,t}(s_{t})=\sum_{i=1}^{N}e_{i,t}(s_{t}),$ (21)
where $e_{i,t}(s_{t})$ is household $i$’s exogenous endowment at time $t$.
Following Mace (1991), I assume that the aggregate idiosyncratic shocks approach zero as $N$ gets large. Consider the situation that the household’s endowment can be decomposed into the permanent component ($e_{i,t}^{P}$), shock linked to macroeconomic trends ($\varrho_{i,t}^{a}(s_{t})$), and idiosyncratic shock ($\varsigma_{i,t}(s_{t})$) as:
$e_{i,t}(s_{t})=e_{i,t}^{P}+\varrho_{i,t}^{a}(s_{t})+\varsigma_{i,t}(s_{t}).$ (22)
[…]
Subject to equation(21), the social planner maximizes equation(20) by choosing an allocation of consumption across households. The first-order condition at time $t$ is derived as follows:
$\omega_{i}\rho^{t}\pi(s_{t})u_{c}(c_{i,t}(s_{t}),\theta_{i,t}(s_{t}))=\lambda(s_{t}),$ (23)
where $\lambda$ is the Lagrange multiplier for the single lifetime resource constraint. To derive my empirical specification, I assume a constant relative risk aversion preference as:
$u(c_{i,t}(s_{t}),\theta_{i,t}(s_{t}))=\frac{c_{i,t}^{1-\sigma}(s_{t})}{1-\sigma}\exp(\sigma\theta_{i,t}(s_{t})),$ (24)
where $\sigma>0$ is the coefficient of relative risk aversion. Below, I consider a particular history of the state of the world so that the subscript $t$ can represent the dependence on the state (Hayashi et al. 1996). This simplifies the notation in equation 23 as:
$\omega_{i}\rho^{t}\pi_{t}c_{i,t}^{-\sigma}\text{exp}(\sigma\theta_{i,t})=\lambda_{t}.$ (25)
Taking the log of equation (25) and aggregating over households, consumption for household $i$ is expressed as:
$\log c_{i,t}=\log c_{t}^{a}+\frac{1}{\sigma}(\log\omega_{i}-\omega^{a})+(\theta_{i,t}-\theta_{t}^{a}),$ (26)
where
$c_{t}^{a}=\exp\left(\frac{1}{N}\sum_{i=1}^{N}\log c_{i,t}\right),~{}~{}~{}~{}~{}\omega^{a}=\frac{1}{N}\sum_{i=1}^{N}\log\omega_{i},~{}~{}~{}~{}~{}\theta_{t}^{a}=\frac{1}{N}\sum_{i=1}^{N}\theta_{i,t}.$ (27)
The first-difference in equation (26) eliminates the household time-constant effect to yield:
$\log c_{i,t}-\log c_{i,t-1}=\log c_{t}^{a}-\log c_{t-1}^{a}+(\theta_{i,t}-\theta_{i,t-1})+(\theta_{t}^{a}-\theta_{t-1}^{a}).$ (28)
Individual income is a conventional proxy variable for the preference shock (Mace 1991). Using the growth rate of household income $\log y_{i,t}-\log y_{i,t-1}$ as a proxy for idiosyncratic shock and adding a measurement error in consumption $\epsilon_{i,t}$, the empirical specification can be characterized as follows:
$\log c_{i,t}-\log c_{i,t-1}=\alpha_{1}(\log c_{t}^{a}-\log c_{t-1}^{a})+\alpha_{2}(\log y_{i,t}-\log y_{i,t-1})+\epsilon_{i,t}.$ (29)
The aggregate consumption growth in equation(29) is assumed to reflect macroeconomic shocks in Tsukishima in the case of my empirical setting. As discussed in Section 3.4, the consumption of the THBS households indeed captures the overall macroeconomic trend. Despite this, interpreting $\alpha_{1}$ is not easy because the THBS households include the factory workers’ households in Tsukishima alone. To deal with this potential issue, I use a two-way fixed-effects model instead of the first-difference model in the spirit of Cochrane (1991). Another potential issue is that $\epsilon_{i,t}$ may include the household-specific time-varying preference shock, which might be correlated with the idiosyncratic income shock. Family size influences the household’s preference because the marginal utility of consumption increases with family size (Jappelli and Pistaferri 2017). As explained, the family size had been fairly stable during the sample period in the THBS households, meaning that the preference shifts had rarely occurred (Section 3.4). To be conservative, however, I include the number of family members as a control variable for representing potential preference shock.
Considering these modifications, the empirical specification can finally be characterized as follows:
$\log c_{i,t}=\gamma\log y_{i,t}+\delta x_{i,t}+\mu_{i}+\phi_{t}+u_{i,t},$ (30)
where $c_{i,t}$ is consumption, $y_{i,t}$ is disposable income, $x_{i,t}$ is family size, $\mu_{i}$ is the household fixed effect, $\phi_{t}$ is the time fixed effect, and $u_{i,t}$ is a random error term. The estimate of $\gamma$ shall range from zero (i.e., idiosyncratic shocks are perfectly insured) to one (for the absence of insurance).
In order to test the risk-coping mechanisms, I consider the following representation of individual consumption in the spirit of Fufchamps and Lund (2003).
$c_{i,t}=y_{i}^{P}+y_{i,t}^{T}+r_{i,t},$ (31)
where $y_{i}^{P}$ is permanent income and $y_{i,t}^{T}$ is transitory income. $r_{i,t}=\bm{z}^{\prime}_{i,t}\bm{1}$ is a linear combination of the net temporary incomes expressed using a $k\times 1$ net income vector ($\bm{z}$) and a vector of ones. Substituting equation 31 into equation 26 and taking the first-difference removes the time-constant components, including the permanent income. It follows that my empirical specification can be characterized as follows:
$r_{i,t}=\kappa\tilde{y}_{i,t}+\eta x_{i,t}+\nu_{i}+\zeta_{t}+\epsilon_{i,t},$ (32)
where $\tilde{y}_{i,t}$ is head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ indicates the household-fixed effect, $\zeta_{t}$ indicates the time-fixed effect, and $\epsilon_{i,t}$ is a random error term. I herein use the head’s earnings ($\tilde{y}_{i,t}$) as the idiosyncratic shock variable instead of the disposable income.100100100In the analysis, I consider the labor-supply adjustment variable and sales of miscellaneous assets as the additional dependent variables in addition to the net income variables $r_{i,t}$ (Section 4.2). The former variable is the total earnings of the family members except for the head, meaning that the simultaneous issue rises if I use the disposable income in this equation. The regression is separately run for each component in the net income vector ($\bm{z}$).
C.2 Raw Relationship between Income and Expenditures
Figure C.1 summarizes the densities of the log-differences (LD) of the semi-monthly expenditures. Figure C.1a illustrates the LD distribution of the total expenditure, whereas Figures C.1bC.1l summarize the LD distributions of the 11 subcategories of the expenditure. These figures confirm no specific outliers in all these expenditure categories.
Figure C.2 summarizes the scatter plots between the LD of the semi-monthly disposable income and expenditures. Figure C.2a illustrates the relationship between the LD of the semi-monthly disposable income and total expenditure. Figures C.2bC.2l illustrate the correlations between the LD of the semi-monthly disposable income and 11 expenditure subcategories.
Generally, rent was paid once per month. This means that the semi-monthly dataset fails to capture the variations in rent payment because it includes some semi-monthly cells with zero values. In addition, the water bills are included in the rent subcategory but not in the utility subcategory. This means that a semi-month cell takes a minimal value when the payment of rent is not included in that cell. Figure C.1c indeed shows that the LD of rent generates some larger values in the semi-monthly panel dataset due to those water bill payments. These cells taking large LD values substantially increased the slope, as shown in Figure C.2c.
Given this technical issue, the adjusted monthly panel dataset is suitable for estimating the income elasticity for the rent category. In fact, Figure C.4c confirms that the scatter plot from the adjusted monthly variation shows a flatter relationship.
Figure C.1: Distributions of the Log-differences of the Semi-monthly Expenditures
(a) Total expenditure
(b) Food
(c) Housing
(d) Utilities
(e) Furniture
(f) Clothing
(g) Education
(h) Medical
(i) Entertainment
(j) Transportation
(k) Gift
(l) Miscellaneous
Notes: The distribution of the log-difference in semi-monthly total expenditure is shown in Figure C.1a. Figures C.1bC.1l show the distributions of the log-differences in semi-monthly expenditure for the 11 subcategories listed in panel A of Table 3. Sources: Created by the author using the THBS sample.
Figure C.1: Distributions of the Log-differences of the Semi-monthly Expenditures
Figure C.2: Raw Relationship between the Log-differences
in Semi-monthly Income and Expenditure
(a) Total expenditure
(b) Food
(c) Housing
(d) Utilities
(e) Furniture
(f) Clothing
(g) Education
(h) Medical
(i) Entertainment
(j) Transportation
(k) Gift
(l) Miscellaneous
Notes: Figure C.2a illustrates the relationship between the log-differences in semi-monthly income and expenditure. Figures C.2bC.2l illustrate the relationships between the log-differences in semi-monthly disposable income and the 11 expenditure subcategories listed in Panel A in Table 3. The range of the y-axis is fixed across all the figures for comparability. Sources: Created by the author using the THBS sample.
Figure C.2: Raw Relationship between the Log-differences
in Semi-monthly Income and Expenditure
Figure C.3: Distributions of the Log-differences of the Monthly Expenditures
(a) Total expenditure
(b) Food
(c) Housing
(d) Utilities
(e) Furniture
(f) Clothing
(g) Education
(h) Medical
(i) Entertainment
(j) Transportation
(k) Gift
(l) Miscellaneous
Notes: The distribution of the log-difference in adjusted monthly total expenditure is shown in Figure C.3a. Figures C.3bC.3l show the distributions of the log-differences in adjusted monthly expenditure for the 11 subcategories listed in panel A of Table 3. Sources: Created by the author using the THBS sample.
Figure C.3: Distributions of the Log-differences of the Monthly Expenditures
Figure C.4: Raw Relationship between the Log-differences
in Monthly Income and Expenditure
(a) Total expenditure
(b) Food
(c) Housing
(d) Utilities
(e) Furniture
(f) Clothing
(g) Education
(h) Medical
(i) Entertainment
(j) Transportation
(k) Gift
(l) Miscellaneous
Notes: Figure C.4a illustrates the relationship between the log-differences in semi-monthly income and expenditure. Figures C.4bC.4l illustrate the relationships between the log-differences in semi-monthly disposable income and the 11 expenditure subcategories listed in Panel A in Table 3. The range of the y-axis is fixed across all the figures for comparability. Sources: Created by the author using the THBS sample.
Figure C.4: Raw Relationship between the Log-differences
in Monthly Income and Expenditure
C.3 Retail Sales in Tokyo City: A Retail Store Survey of 1935
Table C.1 summarizes the contribution of regular customers to their total sales, social classes of the customers, and sales methods in the retailers selling white rice, fish, greengrocers, and fuels (firewood and charcoal) by scale of business. In the original document, the figures for “Total” in the firewood and charcoal category are missing in “Sales Method (%)”. For these three cells, I filled the averages over the three business-size categories in each sales method: 20.4% for “Cash”, 78.4% for “Credit purchase”, and 1.3% for “Monthly installment”.
Table C.1: Customers, Sales Methods, and Competitiveness in the Retailers in Tokyo City
Selling Item Scale of Number of Share of sales Social Class of Sales Method (%) Number of
Business stores for regular the Regular Customer (%) Cash Credit Monthly competitors
reported customer (%) Upper Middle Other purchase installment reported
White Rice Small 19 69.7 12.1 49.5 38.4 29.5 70.2 0.3 7.3
Medium 55 69.5 17.5 41.6 40.9 26.9 72.6 0.5 8.7
Large 27 81.7 15.9 51.5 32.6 26.7 71.0 2.3 8.8
Total 101 73.1 16 45.7 38.2 26.9 71.5 1.7 8.5
Fish Small 34 64.5 5.9 53.2 40.9 42.7 57.3 0 5.1
Medium 20 62.6 10 58.0 32.0 43.8 56.2 0 5.3
Large 5 72.0 0 86.0 14.0 44.8 55.2 0 3.4
Total 59 64.6 6.8 57.6 35.6 43.9 56.1 0 5.0
Greengrocer Small 46 42.2 8 39.2 52.7 59.7 40.3 0 6.5
Medium 17 57.1 8.8 43.5 47.6 65.0 35.0 0 6.1
Large 3 36.7 33.3 40.0 26.7 79.7 20.3 0 3.7
Total 66 45.4 9.4 40.4 50.2 65.8 34.2 0 6.2
Firewood and charcoal Small 78 68.5 13.5 50.7 35.8 25.0 74.2 0.8 8.4
Medium 17 77.7 23.5 49.1 27.4 19.3 79.1 1.6 8.9
Large 5 33.3 23.0 60.0 17.0 16.8 81.8 1.4 10.6
Total 100 51.1 15.7 50.9 33.4 20.4 78.4 1.3 8.6
Notes: This table summarizes the statistics on the sales in the retail shops selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935. 1. Small, middle, and large-scale businesses include retailers with annual sales of less than 10,000, 10,000–30,000, and 30,000+ yen, respectively. 2. The number of stores reported the share of regular customers (%) are 90, 56, 61, and 93 in the white rice, fish, greengrocers, and firewood and charcoal categories, respectively. 3. The share of regular customers is defined as the percentage of sales for regular customers in the annual total sales in 1935. 4. The definition of the social classes (“Upper”; “Middle”; “Other”) is not clearly defined in the report. Therefore, this must be based on the subjective judgment of the retailers. 5. The share of sales method is based on the percentage of sales using each method in the annual total sales in 1935. 6. The number of competitors indicates the number of peers within a three-cho (approximately 327 meters) radius. Source: Tokyo City Chamber of Commerce and Industry 1937, pp. legend; 20–21; questionnaire.
References used in the Appendices
  • (1)
  • (2) Arellano, Manuel. “Computing standard errors for robust within-groups estimators.” Oxford Bulletin of Economics and Statistics 49, no. 4 (1987): 431–434.
  • (3)
  • (4) Cochrane, John H. “A Simple Test of Consumption Insurance.” Journal of Political Economy 99, no. 5 (1991): 957–976.
  • (5)
  • (6) Choi, In “Unit root tests for panel data”. Journal of International Money and Finance 20, no. 2 (2001): 249–272.
  • (7)
  • (8) James, John A., and Isao Suto. “Early Twentieth-century Japanese Worker Saving: Precautionary Behaviour before a Social Safety Net.” Cliometrica 5, no. 1 (2011): 1–25.
  • (9)
  • (10) Jappelli, Tullio, and Luigi Pistaferri. The economics of consumption: Theory and evidence. New York, NY: Oxford University Press, 2017.
  • (11)
  • (12) Kawai, Takao. “Shakai chōsa houhoushi ni tsuite (On history of social research methods)” [in Japanese]. Hōgaku Kenkyū 53, no. 9 (1980): 41–67.
  • (13)
  • (14) Kawai, Takao. “Tsukishima chōsa saikōsatsu, ichi (Tsukishima Survey reconsidered (1))” [in Japanese]. Hōgaku Kenkyū 54, no. 8 (1981a): 1–26.
  • (15)
  • (16) Kawai, Takao. “Tsukishima chōsa saikōsatsu, ni kan (Tsukishima Survey reconsidered (2, end))” [in Japanese]. Hōgaku Kenkyū 54, no. 9 (1981b): 1–33.
  • (17)
  • (18) Mace, Barbara J. “Full Insurance in the Presence of Aggregate Uncertainty.” Journal of Political Economy 99, no. 5 (1991): 928–956.
  • (19)
  • (20) Mazzocco, Maurizio and Shiv Saini. “Testing Efficient Risk Sharing with Heterogeneous Risk Preferences.” The American Economic Review 102, no. 1 (2012): 428–468.
  • (21)
  • (22) Mazzocco, Maurizio and Shiv Saini. “Testing Efficient Risk Sharing with Heterogeneous Risk Preferences: Reply.” The American Economic Review 108, no. 10 (2018): 3114–3115.
  • (23)
  • (24) Miyoshi, Toyotaro. “Tsukishima chōsa no seiritsu to sonokeika ni tsuite (The establishment and progress of the Tsukishima Survey)” [in Japanese]. Meiseidaigaku Kenkyū Kiyō 16, (1980): 31–48.
  • (25)
  • (26) Takano, Iwasaburo. Honpō shakaitōkei ron (Social Statistics in Japan) [in Japanese]. Tokyo: Kaizō sha.
  • (27)
  • (28) Pearson, Karl. On the criterion that a given system of deviations from the probable in the case of a correlated system of variables is such that it can be reasonably supposed to have arisen from random sampling. Philosophical Magazine 50, no. 302, (1900): 157–175.
  • (29)
  • (30) Schulhofer-Wohl, Sam. “Heterogeneity and Tests of Risk Preference”. Journal of Political Economy 119, no.5, (2011): 925–958.
  • (31)
  • (32) Sekiya, Koichi. Seikatsu koten sōsho, volume 6: Tsukishima chōsa (A classic series of books about life, volume 6: Tsukishima Survey) [in Japanese]. Tokyo: Kōseikan.
  • (33)
  • (34) Shrinivas, Aditya and Marcel Fafchamps. “Testing Efficient Risk Sharing with Heterogeneous Risk Preferences: Comment.” The American Economic Review 108, no. 10 (2018): 3104–3113.
  • (35)
  • (36) Shibuya, Ryuichi, Kameji Suzuki, and Shojiro Ishiyama. Nihon no Shichiya (Pawn Shops in Japan) [in Japanese]. Tokyo: Waseda University Press, 1982.
  • (37)
  • (38) Shibuya, Ryuichi. Kourigashikinyū no tenkai kōzō (Development and Structure of Usuries) [in Japanese]. Tokyo: Nihon Tosho Center, 2000.
  • (39)
  • (40) Takeda, Naoko. “Shitsuteki chōsa data no nijibunseki (Qualitative Secondary Analysis and Reuse of the Original Research Data Set of the Japanese Sociological Classic)” [in Japanese]. Nihonrōdōkenkyūzasshi 57, no. 12 (2015): 70–80.
  • (41)
  • (42) Tanaka, Hikaru. Mouhitotsu no kinyū system (Another Financial System) [in Japanese]. Nagoya: Nagoya University Press, 2018.
  • (43)
  • (44) Terade, Kouji. “Tsukishima hōkokusho dai nishū (A study of “The volume II of the report on the investigation at Tsukishima”)” [in Japanese]. Mitagakkaizasshi 75, no. 6 (1982): 94–108.
  • (45)
  • (46) Odaka, Konosuke. “Japanese-style Labour Relations.” In The Japanese Economic System and Its Historical Origins, edited by Tetsuji Okazaki and Masahiro Okuno-Fujiwara. Oxford: Oxford University Press, 1999.
  • (47)
  • (48) Patrick, Royston and Douglas G. Altman. “Regression using Fractional Polynomials of Continuous Covariates: Parsimonious Parametric Modeling.” Journal of the Royal Statistical Society. Series C (Applied Statistics) 43 no. 3 (1994): 429–467.
  • (49)
  • (50) Wooldridge, Jeffrey M. Econometric Analysis of Cross Section and Panel Data. Cambridge: MIT Press, 2002.
  • (51)
Documents and Database used in the Appendices
  • (1)
  • (2)
  • (3)
  • (4) Central Federation of Industrial Associations. Sangyō kumiai chōsa siryō, volume 9 (Survey report of the industrial association, volume.9). [in Japanese] Tokyo: Central Federation of Industrial Associations, 1925.
  • (5)
  • (6) Committee for the History of the Labor Movement. Nihon rōdōundō shiryō (Historical Documents of the Japanese Labor Movement, volume 10). [in Japanese] Tokyo: Committee for the History of the Labor Movement, 1964.
  • (7)
  • (8) Chūō Ward. Chūō ku enkaku zushū, Tsukishima hen (Chūō Ward History Chart, Tsukishima edition). [in Japanese] Tokyo: Jinbun sha, 1994.
  • (9)
  • (10) Department of Health, Ministry of the Interior. Tokyo-shi Kyobashi-ku Tsukishima ni okeru jicchichōsahōkoku (dai isshū) (Report of the Field Study in Tsukishima in Kyobashi Ward of Tokyo City, Volume 1) [in Japanese]. Tokyo: Department of Health, Ministry of the Interior, 1923a.
  • (11)
  • (12) Department of Health, Ministry of the Interior. Tokyo-shi Kyobashi-ku Tsukishima ni okeru jicchichōsahōkoku (dai isshū, furoku no ichi) (Report of the Field Study in Tsukishima in Kyobashi Ward of Tokyo City, Appendix 1 of Volume 1) [in Japanese]. Tokyo: Department of Health, Ministry of the Interior, 1923b.
  • (13)
  • (14) Department of Health, Ministry of the Interior. Tokyo-shi Kyobashi-ku Tsukishima ni okeru jicchichōsahōkoku (dai isshū, furoku no ni) (Report of the Field Study in Tsukishima in Kyobashi Ward of Tokyo City, Appendix 2 of Volume 1) [in Japanese]. Tokyo: Department of Health, Ministry of the Interior, 1923c.
  • (15)
  • (16) Ito, Osamau, and Saito Nao. Kinyu gyo(Financial Industry) [in Japanese]. Tokyo: Nihon Keieishi Kenkyūjyo, 2019.
  • (17)
  • (18) Japan Meteorological Agency (database) [in Japanese]. https://www.data.jma.go.jp/gmd/risk/obsdl/index.php.
  • (19)
  • (20) Kitazawa, Shinjiro. Tokyo niokeru kikaikogyō no jyukurenshokkō toshiteno shiagekō narabini senbankō no chinginhōkoku (The report on a survey of wages of fitters and turners as skilled-workers in the machine industry in Tokyo). [in Japanese] Tokyo: Ohara Institute of Social Research, 1924.
  • (21)
  • (22) Kyoto City Office.Kyōto shi tōkeisho, dai jyūsan kai (Statistical Table of Kyoto City, volume. 13). [in Japanese] Kyoto: Kyoto City Office, 1922.
  • (23)
  • (24) Kyoto Prefecture.Kyōto fu tōkeisho, taishō kyū nen, dai san hen (Statistical Table of Kyoto Prefecture, 1920 edition, vol. 3). [in Japanese] Kyoto: Kyoto Prefecture, 1922.
  • (25)
  • (26) Ministry of Agriculture and Commerce. Kōjyō Tsūran (Taishō jyūnen jyūgatsu) (Handbook of Factories, November 1921 edition) [in Japanese]. Tokyo: Ministry of Agriculture and Commerce, 1921.
  • (27)
  • (28) Ministry of Posts and Telecommunications. Denki jigyō yōran, dai jyūyon kai (Handbook of Electricity Industry, vol. 14). [in Japanese] Tokyo: Electric Association, 1922.
  • (29)
  • (30) Nagoya City Office.Nagoya shi tōkeisho, dai nijyūni kai (Statistical Table of Kyoto City, volume. 22). [in Japanese] Nagoya: Nagoya City Office, 1922.
  • (31)
  • (32) Ohkawa, Kazushi, Miyohei Shinohara, and Mataji Umemura. Estimates of Long-term Economic Statistics of Japan since 1868, volume.8. [in Japanese] Tokyo: Toyo Keizai Shinposha, 1967.
  • (33)
  • (34) Osaka City Office. Osakashi tōkeisho, Taishō kyūnen (Statistical Table of Osaka City, 1920 Edition). [in Japanese] Osaka: Osaka City Office, 1922.
  • (35)
  • (36) Postal Savings Bureau. Chochikukyoku nenpō, volume 30 (Annual Statistical Report of the Postal Savings Bureau, volume 30). [in Japanese] Postal Savings Bureau, 1924.
  • (37)
  • (38) Statistics Bureau of the Cabinet. Taishō kyūnen kokusei chōsa hōkoku, zenkoku no bu, dai ikkan (Population Census of Japan, Vol. 1, 1920 Edition). [in Japanese] Tokyo: Statistics Bureau of the Cabinet, 1928.
  • (39)
  • (40) Statistics Bureau of the Cabinet. Taoshō kyūnen kokuseichōsahōkoku, zenkoku no bu, dai nikan (Population Census of Japan, Vol. 2, 1920 Edition). [in Japanese] Tokyo: Statistics Bureau of the Cabinet, 1929a.
  • (41)
  • (42) Tokyo Chamber of Commerce.Tokyoshinai shōshōkōgyō kinyūchōsa siryō (Survey report on the finance for small commercial and industrial enterprises in Tokyo City) [in Japanese]. Tokyo Chamber of Commerce, 1918.
  • (43)
  • (44) Tokyo Chamber of Commerce and Industry.Tokyoshinai niokeu kourigyōkeiei narabini kinyūchōsa (Survey on retail business management and finance in Toky City) [in Japanese]. Tokyo Chamber of Commerce and Industry, 1937.
  • (45)
  • (46) Tokyo City Social Affairs Bureau.Tokyo shinai saimin no nyūshichi nikansuru chōsa (Survey on the use of pawnshops among the poor in Tokyo City) [in Japanese]. Tokyo City Social Affairs Bureau, 1921.
  • (47)
  • (48) Tokyo City Office.Tokyo shi tōkei nenpyō, dai jyūnanakai (Annual Statistics of the City of Tokyo, vol.17) [in Japanese]. Tokyo: Tokyo City Office, 1921.
  • (49)
  • (50) Tokyo City Office.Tokyo shi tōkeisho, dai jyūhachi kai (Annual Statistics of the City of Tokio, volume. 18). [in Japanese] Tokyo: Tokyo City Office, 1922b.
  • (51)
  • (52) Tokyo City Office.Tokyo shi tōkeisho, dai jyūkyū kai (Annual Statistics of the City of Tokio, volume. 19). [in Japanese] Tokyo: Tokyo City Office, 1923.
  • (53)
  • (54) Tokyo City Office. Tokyo shi shisei tōkei genpyō, dai ikkan (Tokyo City Municipal Statistics Original Table, vol.1) [in Japanese]. Tokyo: Tokyo City Office, 1922a.
  • (55)
  • (56) Tokyo City Office. Chūshō shōkōgyōsha no kinyūmondai (Chūshō shōkōgyōsha shinkō chōsakai siryō, daini) (Financial issues for small and medium commercial and industrial enterprises (Material of the Committee for the Promotion of Small and Medium Commercial and Industrial Development, volume 2)) [in Japanese]. Tokyo: Tokyo City Office, 1935.
  • (57)
  • (58) Tokyo City Statistics Division. Tokyo shi oyobi kinkō chōson rōdōtōkeigenpyō (dai ikkai, dai nikan) (Labor Statistics Table for Tokyo City and Suburbs, issue 2 of volume 1) [in Japanese]. Tokyo: Tokyo City Office, 1926a.
  • (59)
  • (60) Tokyo City Statistics Division. Shichiyagyō no tōkei chōsa (Statistical Survey of the Pawnshop Business) [in Japanese]. Tokyo: Tokyo City Office, 1926b.
  • (61)
  • (62) Tokyo Institute for Municipal Research. Toshishominkinyū nikansuru chōsa, daiissatsu toshikinyū no gaikyō (Survey of Financial Institutions among Common People, volume. 1, Overview of the Financial Institutions in the Cities) [in Japanese]. Tokyo Institute for Municipal Research, 1925a.
  • (63)
  • (64) Tokyo Institute for Municipal Research. Toshishominkinyū nikansuru chōsa, dainisatsu toshichochiku ginkō (Survey of Financial Institution among Common People, volume. 2, Municipal Savings Banks) [in Japanese]. Tokyo Institute for Municipal Research, 1925b.
  • (65)
  • (66) Tokyo Institute for Municipal Research. Toshishominkinyū nikansuru chōsa, daisansatsu shigaichi shinyōkumiai (Survey of Financial Institutions among Common People, volume. 2, Credit Unions in the cities) [in Japanese]. Tokyo Institute for Municipal Research, 1925c.
  • (67)
  • (68) Tokyo Institute for Municipal Research. Toshishominkinyū nikansuru chōsa, daiyonsatsu kousetsu shichiho (Survey of Financial Institutions among Common People, volume. 2, Pubic pawnshops) [in Japanese]. Tokyo Institute for Municipal Research, 1926.
  • (69)
  • (70) Tokyo Legislative Research Association. Chokin to Mujin (Saving and Mutual Loan Association) [in Japanese]. Tokyo: Tokyo Legislative Research Association, 1915.
  • (71)
  • (72) Umemura, Mataji, Takamatsu, Nobukiyo, and Itoh, Shigeru. Estimates of Long-term Economic Statistics of Japan since 1868, volume.13. [in Japanese] Tokyo: Toyo Keizai Shinposha, 1983.
  • (73)
  • (74) Umemura, Mataji, Akasaka, Keiko, Minami, Ryoshin, Takamatsu, Nobukiyo, Arai, Kurotake, and Itoh, Shigeru. Estimates of Long-term Economic Statistics of Japan since 1868, volume.2. [in Japanese] Tokyo: Toyo Keizai Shinposha, 1988.
  • (75)
  • (76) Water Works Association. Jyōsuidō tōkei oyobi hōkoku (Statistics and Reports of the Water Supply, Vol. 1). [in Japanese] Tokyo: Water Works Association, 1922.
  • (77)
Archival Sources
  • (1)
  • (2) The Ohara Institute for Social Research, Archives of the Tsukishima Survey (THBS, unreleased).
  • (3)
  • (4) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7-2).
  • (5) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7-3).
  • (6) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7-4).
  • (7)
  • (8) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-12).
  • (9) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-13).
  • (10) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-14).
  • (11) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-15).
  • (12)